### v2

# Approximate independence of permutation mixtures

Abstract
We prove bounds on statistical distances between high-dimensional exchangeable mixture distributions (which we call permutation mixtures) and their i.i.d. counterparts. Our results are based on a novel method for controlling $\chi^{2}$ divergences between exchangeable mixtures, which is tighter than the existing methods of moments or cumulants. At a technical level, a key innovation in our proofs is a new Maclaurin-type inequality for elementary symmetric polynomials of variables that sum to zero and an upper bound on permanents of doubly-stochastic positive semidefinite matrices. Our results imply a de Finetti-style theorem (in the language of Diaconis and Freedman, 1987) and general asymptotic results for compound decision problems, generalizing and strengthening a result of Hannan and Robbins (1955).
1 Introduction
Let $n$ be a positive even integer and $\mu$ a positive real number. Consider two probability distributions on $\mathbb{R}^{n}$:
  • Under $\mathbb{P}_{n}$, the observation $X$ is of the form $\vartheta+Z$, where $\vartheta$ is uniformly distributed on the subset of balanced vectors in {±$\{\pm\mu\}^{n}$ (i.e., vectors with exactly $n/2$ entries equal to $-\mu$ and $n/2$ equal to $+\mu$) and $Z\sim{\mathcal{N}}(0,I_{n})$ is independent Gaussian noise;
  • Under $\mathbb{Q}_{n}$, the observation $X$ is of the form $\vartheta+Z$, where $\vartheta$ is uniformly distributed on {±$\{\pm\mu\}^{n}$ and $Z\sim{\mathcal{N}}(0,I_{n})$ is independent Gaussian noise. Note that the entries of $X$ under $\mathbb{Q}_{n}$ are themselves i.i.d., and that the marginal distribution of each coordinate of $X$ under $\mathbb{Q}_{n}$ and $\mathbb{P}_{n}$ is the same.
In other words, let $\nu_{\mathbb{P}}$ denote the distribution of $n$ uniformly random draws from the multiset $\{-\mu,\dots,-\mu,+\mu,\dots,+\mu\}$ without replacement and $\nu_{\mathbb{Q}}$ the corresponding distribution of draws with replacement. Then $\mathbb{P}_{n}=\nu_{\mathbb{P}}\star{\mathcal{N}}(0,I_{n})$ and $\mathbb{Q}_{n}=\nu_{\mathbb{Q}}\star{\mathcal{N}}(0,I_{n})$. For which values of $n$ and $\mu$ are $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ statistically close?
The law of large numbers implies that a random vector $\vartheta$ with i.i.d. {±$\{\pm\mu\}$ entries will be nearly balanced, and, if $\mu\lesssim 1$, the Gaussian noise will make it difficult to detect whether $\vartheta$ is perfectly balanced or only nearly so. It is therefore natural to conjecture that $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ will not be distinguishable from each other as long as $\mu$ is not too large. Surprisingly, however, as we show in Section 2, existing approaches to bounding statistical distances between mixtures fail to establish this fact. The goal of this paper is to develop new strategies for analyzing such high-dimensional mixtures, which achieve tight bounds.
As an example, a corollary of one of our main results shows that for the model described above,
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})=\begin{cases}O(\mu^{4})&\text{if }\mu\leq 1,\\ O(\exp(\mu^{2}))&\text{if }\mu>1.\end{cases}$ (1)
Strikingly, (1) implies that the $\chi^{2}$ divergence between these two high-dimensional models is bounded independent of the dimension $n$—in particular, the sequence $\mathbb{P}_{n}$ is contiguous to $\mathbb{Q}_{n}$ in the sense of Le Cam [LC60], so that asymptotic properties of the mixture $\mathbb{P}_{n}$ can be analyzed under the simpler product measure $\mathbb{Q}_{n}$. The bound in (1) also reveals other features: since $\chi^{2}({\mathcal{N}}(+\mu,1)\|{\mathcal{N}}(-\mu,1))=e^{4\mu^{2}}-1$, the $\chi^{2}$ divergence between the two models is polynomially smaller than the “single-letter” $\chi^{2}$ divergence that would arise from comparing the two univariate marginal distributions that appear in the definition of $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$. As we shall show, up to constants, the bounds in (1) are tight.
Our results hold for much more general models, which we call permutation mixtures. Let ${\mathcal{P}}$ be a collection of probability measures on a fixed probability space. Given $P_{1},\cdots,P_{n}\in{\mathcal{P}}$, we define two mixture models:
  • The distribution $\mathbb{P}_{n}$ of the observation $X=(X_{1},\cdots,X_{n})$ is the “permutation mixture” defined by
    $(X_{1},\cdots,X_{n})\sim\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\otimes_{i=1}^{n}P_{\pi(i)}\right];$ (2)
  • The distribution $\mathbb{Q}_{n}$ of the observation $X=(X_{1},\cdots,X_{n})$ is an i.i.d. product of one-dimensional mixtures defined by
    $(X_{1},\cdots,X_{n})\sim\left(\frac{1}{n}\sum_{i=1}^{n}P_{i}\right)^{\otimes n}.$ (3)
As before, $\mathbb{P}_{n}$ is a measure with exchangeable but not independent coordinates, whereas $\mathbb{Q}_{n}$ is a product measure whose one-dimensional marginals agree with those of $\mathbb{P}_{n}$. Note that the simple example given above corresponds to the setting where ${\mathcal{P}}=\{{\mathcal{N}}(-\mu,1),{\mathcal{N}}(+\mu,1)\}$ and $P_{1},\dots,P_{n}$ are divided evenly between the two possibilities. Our main contribution is a set of new techniques for bounding $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ for general permutation mixtures.
The study of permutation mixtures is motivated by important theoretical questions in a number of areas:
Compound decisions and empirical Bayes. In the compound decision problem [Rob51], the statistician observes independent data $X_{1},\dots,X_{n}$, where $X_{i}\sim P_{\theta_{i}}$ for $i=1,\dots,n$; the goal is to estimate the vector of parameters under a separable loss $L(\theta,\widehat{\theta}):=\sum_{i=1}^{n}\ell(\theta_{i},\widehat{\theta}_{i})$. This setting is the starting point of the theory of empirical Bayes [Efr19], which views the parameters as having been drawn i.i.d. from a prior distribution, which is in turn estimated from the observations. An oracle version of this problem was suggested by Brown and Greenshtein [BG09] (see also [JZ09]), in which the statistician knows the unordered set $\{\theta_{1},\dots,\theta_{n}\}$, but not their correspondence to the observations $(X_{1},\dots,X_{n})$. Equivalently, the observations may be viewed as arising from the permutation mixture $\mathbb{P}_{n}$. As Greenshtein and Ritov [GR09] observe, however, most oracle inequalities in the empirical Bayes literature compare to an oracle that knows the unordered set $\{\theta_{1},\dots,\theta_{n}\}$ but is restricted to the use of “simple” decision rules, of the form $\hat{\theta}=(\Delta(X_{1}),\dots,\Delta(X_{n}))$ for some fixed univariate function $\Delta$. Such rules are in fact optimal when the observations arise from not from $\mathbb{P}_{n}$ but from its independent counterpart $\mathbb{Q}_{n}$. The empirical Bayes literature contains several ad hoc techniques for bounding the difference between these two approaches, going back to [HR55]. Our results can be used to give new, general risk bounds in this setting, see Corollary 1.7.
de Finetti-style theorems. The classical de Finetti theorem [DF29] asserts that an infinite exchangeable sequence of binary random variables is a mixture of i.i.d. Bernoulli sequences. Since the work of Diaconis [Dia77], there has been interest in obtaining quantitative forms of this theorem for finite exchangeable sequences [Sta78, DF80, GK21], and, more broadly, to obtain “de Finetti-style theorems” in general settings [DF87]. Our main results, which establish that the exchangeable measure $\mathbb{P}_{n}$ is close to the product measure $\mathbb{Q}_{n}$, are theorems of this type. As Theorem 1.5 shows, our bounds yield new approximation results for general exchangeable sequences.
Mean-field approximation. Whether a high-dimensional distribution can be accurately approximated by a product measure is a central question in probability theory. These “mean-field” approximations are useful to study the behavior of large random systems with weak dependence and open the door to precise asymptotic computations of, e.g., the free energy for complicated models arising from statistical physics [Par88, JRK19, LMY24]. The present case can be seen as a simple example of this philosophy: is the interaction between coordinates in a permutation mixture weak enough that the mixture distribution $\mathbb{P}_{n}$ is close to a measure with i.i.d. coordinates? Our main theorems answer this question in the affirmative.
Information geometry of high-dimensional mixture models. High-dimensional mixtures—such as the measures $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ defined above—are ubiquitous models in theoretical statistics (see, e.g., [LNS99, CL11, IS12, WY16, JHW18]). The analysis of such measures is routine in the univariate case (or, more generally, for product measures), but obtaining sharp bounds is difficult for high-dimensional mixtures whose coordinates are not independent. The setting we study is one in which we can rigorously compare $\mathbb{P}_{n}$ to a simpler, independent counterpart $\mathbb{Q}_{n}$. This comparison may be directly relevant in analyzing mixtures with permutation invariance. Moreover, as we highlight in Section 1.2, our proofs are based on new techniques for bounding divergences between high-dimensional mixtures, which may be of broader applicability.
Our main results are stated in terms of two information-theoretic quantities: given two probability measures $P$ and $Q$ on the same probability space, we define the $\chi^{2}$ divergence and squared Hellinger distance:
$\chi^{2}(P\|Q)=\int\frac{(\mathrm{d}P-\mathrm{d}Q)^{2}}{\mathrm{d}Q}\,\qquad H^{2}(P,Q)=\int\left(\sqrt{\mathrm{d}P}-\sqrt{\mathrm{d}Q}\right)^{2}.$
Both are fundamental measures of the statistical similarity between $P$ and $Q$; see, e.g., [Tsy09].
We now present our main bounds of the $\chi^{2}$ divergence, $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$, between the permutation mixture $\mathbb{P}_{n}$ in (2) and its i.i.d. counterpart $\mathbb{Q}_{n}$ in (3). Our upper bounds will depend on several quantities of the distribution family ${\mathcal{P}}$, defined as follows.
Definition 1.1.
For a given family ${\mathcal{P}}$ of probability distributions over the same space, define:
  1. 1.
    The $\chi^{2}$ channel capacity, denoted by ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$:
    ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}}):=\sup_{\rho\in\Delta({\mathcal{P}})}I_{\chi^{2}}(P;X)=\sup_{\rho\in\Delta({\mathcal{P}})}\mathbb{E}_{P\sim\rho}\left[\chi^{2}(P\|\mathbb{E}_{P^{\prime}\sim\rho}[P^{\prime}])\right],$
    where $P\sim\rho$ and $X|P\sim P$ in the $\chi^{2}$ mutual information, and $\Delta({\mathcal{P}})$ denotes the class of all prior distributions over ${\mathcal{P}}$;
  2. 2.
    The $\chi^{2}$ diameter, denoted by ${\mathsf{D}}_{\chi^{2}}({\mathcal{P}}):=\sup_{P_{1},P_{2}\in{\mathcal{P}}}\chi^{2}(P_{1}\|P_{2})$;
  3. 3.
    The $H^{2}$ diameter, denoted by ${\mathsf{D}}_{H^{2}}({\mathcal{P}}):=\sup_{P_{1},P_{2}\in{\mathcal{P}}}H^{2}(P_{1},P_{2})$;
  4. 4.
    The maximum $H^{2}$ singularity, denoted by $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$:
    $\mathsf{\Delta}_{H^{2}}({\mathcal{P}}):=\left(1-\frac{{\mathsf{D}}_{H^{2}}({\mathcal{P}})}{2}\right)^{-2}.$
When the $\chi^{2}$ divergence is replaced by the Kullback–Leibler (KL) divergence in Definition 1.1, the $\chi^{2}$ channel capacity coincides with the usual notion of channel capacity in information theory. We have ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})$ by convexity, and $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1$ by [PW24, Eqn. (7.33)]. In addition, $H^{2}(P_{1},P_{2})=2$ if and only if $P_{1}$ and $P_{2}$ are mutually singular, in which case it is easy to see that $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ must grow with $n$ in the worst case (cf. Lemma 5.1). Therefore, the quantity $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ measures the extent of singularity of ${\mathcal{P}}$. Finally, we note that all quantities (whenever finite) depend only on ${\mathcal{P}}$ but not on the dimension $n$ of $\mathbb{P}_{n}$ or $\mathbb{Q}_{n}$. We present our main theorem below, which upper bounds the $\chi^{2}$ divergence using only the above quantities.
Theorem 1.2.
The following upper bounds hold:
  1. 1.
    If ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})<\infty$, then
    $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq 10\sum_{\ell=2}^{n}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{\ell};$ (4)
  2. 2.
    If $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty$ and ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})<\infty$, then
    $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}-1;$ (5)
  3. 3.
    If ${\mathsf{D}}_{\chi^{2}}({\mathcal{P}})<\infty$, then
    $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq(1+{\mathsf{D}}_{\chi^{2}}({\mathcal{P}}))^{1+{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}-1.$ (6)
Theorem 1.2 shows that if ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq 1-\delta$ for some constant $\delta>0$, then the $\chi^{2}$ divergence is upper bounded by a dimension-independent constant $O_{\delta}({\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{2})$. In addition, even if ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\geq 1$ but $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty$, the $\chi^{2}$ divergence is still upper bounded by a large albeit dimension-independent constant. This shows that the contiguity relation $\{\mathbb{P}_{n}\}\triangleleft\{\mathbb{Q}_{n}\}$ is not unique to the Gaussian family but holds for a large class of probability families ${\mathcal{P}}$. Specializing ${\mathcal{P}}$ to specific classes of distributions leads to the following corollary.
Corollary 1.3.
The following results hold for specific families:
  1. 1.
    Gaussian family ${\mathcal{P}}=\{{\mathcal{N}}(\theta,1):|\theta|\leq\mu\}$: there exists an absolute constant $\mu_{c}>0$ such that
    $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})=\begin{cases}O(\mu^{4})&\text{if }\mu\leq\mu_{c},\\ \exp(O(\mu^{3}))-1&\text{if }\mu>\mu_{c}.\end{cases}$
  2. 2.
    Gaussian family with small support ${\mathcal{P}}=\{{\mathcal{N}}(\theta,1):|\theta|\leq\mu,\theta\in\Theta\}$ with $|\Theta|<\infty$: there exists an absolute constant $\mu_{c}>0$ such that
    $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})=\begin{cases}O(\mu^{4})&\text{if }\mu\leq\mu_{c},\\ \exp(O(|\Theta|\mu^{2}))-1&\text{if }\mu>\mu_{c}.\end{cases}$
  3. 3.
    Bernoulli family ${\mathcal{P}}=\{\text{\rm Bern}(p):p\in[\varepsilon,1-\varepsilon]\}$: there exists an absolute constant $\varepsilon_{c}\in(0,1/2)$ such that
    $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})=\begin{cases}O((1-2\varepsilon)^{4})&\text{if }\varepsilon\geq\varepsilon_{c},\\ O\left(\frac{1}{\varepsilon}\right)&\text{if }0<\varepsilon<\varepsilon_{c}.\end{cases}$
  4. 4.
    Poisson family ${\mathcal{P}}=\{\text{\rm Poi}(\lambda):\lambda\in[0,M]\}$: it holds that
    $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})=\begin{cases}O(M^{2})&\text{if }M\leq 1,\\ \exp\left(O(M^{3/2})\right)-1&\text{if }M>1.\end{cases}$
The proof of Corollary 1.3 follows from Theorem 1.2 and the computations of $({\mathsf{C}}_{\chi^{2}}({\mathcal{P}}),\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))$, which we defer to Section C.1. For the example presented at the beginning of the introduction, a weaker upper bound $\exp(O(\mu^{2}))$ of (1) follows from point 2 of Corollary 1.3; we refer to Section 3.1 for the proof of the stronger form (1) (via the inequality ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq 1-\exp(-\mu^{2})$ in (2)). We will also discuss the tightness of these upper bounds in Section 5.1; typically, the quadratic dependence in (4) is tight when ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ is small, and both the base and exponent of (5) are not improvable in general. However, for specific ${\mathcal{P}}$ there could still be gaps, and we refer the discussions to Section 5.1.
Another direct corollary of Theorem 1.2 is the application of the continuity relation $\{\mathbb{P}_{n}\}\triangleleft\{\mathbb{Q}_{n}\}$ to mutual information with permutation structures. Let $(\theta_{1},\dots,\theta_{n})$ be a (deterministic or random) vector, with $P_{\theta_{1}},\dots,P_{\theta_{n}}\in{\mathcal{P}}$ being probability distributions over a space ${\mathcal{X}}$. Let $\vartheta=(\vartheta_{1},\dots,\vartheta_{n})$ be a uniformly random permutation of $(\theta_{1},\dots,\theta_{n})$, i.e. $\vartheta_{i}=\theta_{\pi(i)}$ with $\pi\sim\text{\rm Unif}(S_{n})$, and $X=(X_{1},\dots,X_{n})$ be a random vector with independent coordinates $X_{i}\sim P_{\vartheta_{i}}$ conditioned on $\vartheta$. From the statistical perspective, the random vector $\vartheta$ represents a permutation prior on the model parameter and $I(\vartheta;X)$ denotes the amount of information $X$ provides for estimating $\vartheta$. From the information theoretic perspective, when the vector $\theta$ represents the random message to be communicated, the maximum mutual information $I(\theta;X)$ over all possible distributions of $\theta$ is the channel capacity of the so-called noisy permutation channel [Mak20]. Our contiguity relation implies that the mutual information in both scenarios is close to its counterpart when $\vartheta\sim\left(\frac{1}{n}\sum_{i=1}^{n}\delta_{\theta_{i}}\right)^{\otimes n}$ is i.i.d.
Corollary 1.4.
The following inequalities hold for the mutual information $I(\vartheta;X)$ and $I(\theta;X)$:
  1. 1.
    For deterministic $\theta=(\theta_{1},\dots,\theta_{n})$,
    00 0 $\leq nI(\vartheta_{1};X_{1})-I(\vartheta;X)$
    $\leq\min\left\{10\sum_{\ell=2}^{n}I_{\chi^{2}}(\vartheta_{1};X_{1})^{\ell},I_{\chi^{2}}(\vartheta_{1};X_{1})(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))\right\}.$ (7)
  2. 2.
    For possibly random $\theta=(\theta_{1},\dots,\theta_{n})$,
    $I(\theta;X)\leq\inf_{\varepsilon>0}\left(n\varepsilon^{2}+\log N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)\right)+{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})),$ (8)
    where $\mathrm{conv}({\mathcal{P}})$ denotes the convex hull of ${\mathcal{P}}$, and
    $N_{\mathrm{KL}}({\mathcal{P}},\varepsilon):=\min\left\{m:\inf_{{\mathcal{P}}_{0}\subseteq{\mathcal{P}}:|{\mathcal{P}}_{0}|=m}\sup_{P\in{\mathcal{P}}}\min_{P_{0}\in{\mathcal{P}}_{0}}{\sf KL}(P\|P_{0})\leq\varepsilon^{2}\right\}$
    is the $\varepsilon$-covering number of the distribution class ${\mathcal{P}}$ under the KL divergence.
In words, (1) implies that the difference in the mutual information $I(\vartheta;X)$ under a permutation prior and an i.i.d. prior on $\vartheta$ is typically independent of $n$, i.e., these priors are “equally informative” for the output $X$ in the asymptotic sense. This implication turns out to find statistical applications in problems with permutation priors; we refer to Section 5.2 for such an example in the Gaussian sequence model. The inequality (8) is an upper bound on the capacity of the noisy permutation channel, where the quantity $\inf_{\varepsilon>0}\left(n\varepsilon^{2}+\log N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)\right)$ is an entropic upper bound on the minimax redundancy of $\mathrm{conv}({\mathcal{P}})^{\otimes n}$ [HO97, YB99], which, by the capacity-redundancy theorem [Kem74], is the channel capacity of an i.i.d. channel $X\sim Q_{\theta}^{\otimes n}$ with $Q_{\theta}\in\mathrm{conv}({\mathcal{P}})$. We note that this entropic upper bound is tight when the KL metric entropy $\log N_{\mathrm{KL}}$ is of the same order as the Hellinger metric entropy [PW24, Theorem 32.5], which holds for Gaussian or Gaussian mixture models [Ney22, Mou23, JPW23]. The inequality (8) claims that the capacity of a noisy permutation channel is essentially upper bounded by its counterpart when the conditional distribution $P_{X|\theta}$ is i.i.d., up to an $O_{\mathcal{P}}(1)$ additive factor. This generalizes the capacity upper bound proved in [TP23] for the finite-dimensional case $\dim({\mathcal{P}})=d$, where $\inf_{\varepsilon>0}\left(n\varepsilon^{2}+\log N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)\right)\sim\frac{d}{2}\log n$.
We mention two additional applications of Theorem 1.2 and its related proof techniques. The first application is a “noisy” finite de Finetti theorem concerning the statistical distance between the output of a kernel applied to a finite exchangeable sequence and a proper mixture of i.i.d. sequences. Let $P_{X^{n}}$ be an exchangeable distribution over ${\mathcal{X}}^{n}$, and ${\mathsf{K}}$ be a probability kernel from ${\mathcal{X}}$ to ${\mathcal{Y}}$ represented by a class of conditional distributions ${\mathcal{P}}:=\{{\mathsf{K}}(\cdot|x):x\in{\mathcal{X}}\}$. Passing $X^{n}$ through the kernel ${\mathsf{K}}^{\otimes n}$ leads to an exchangeable distribution $P_{Y^{n}}$, and the following theorem establishes its statistical distance from a mixture of i.i.d. distributions.
Theorem 1.5.
Let $1\leq k\leq n$ and let $P_{Y^{k}}$ and ${\mathcal{P}}$ be defined as above. There exists a mixture of i.i.d. distributions $Q_{Y^{k}}=\mathbb{E}_{\theta}[Q_{\theta}^{\otimes k}]$ such that
$\chi^{2}\left(P_{Y^{k}}\|Q_{Y^{k}}\right)\leq\frac{k(k-1)}{n(n-1)}\cdot\begin{cases}10\sum_{\ell=2}^{n}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{\ell}&\text{if }{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})<\infty,\\ (e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}-1&\text{if }{\mathsf{C}}_{\chi^{2}}({\mathcal{P}}),\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty.\end{cases}$
In the absence of a smoothing kernel, some well-known results due to Stam [Sta78], Diaconis–Freedman [DF80], and several recent works [GK21, JGK24] yield bounds of the form ${\sf KL}(P_{X^{k}}\|Q_{X^{k}})=O(\frac{k^{2}}{n})$, which can be improved to $O(\frac{k^{2}}{n(n+1-k)})$ when $|{\mathcal{X}}|=O(1)$. In particular, $P_{X^{k}}$ and $Q_{X^{k}}$ are, in general, statistically indistinguishable when $k=o(\sqrt{n})$, and that this requirement can be improved to $k=o(n)$ when ${\mathcal{X}}$ is small. [^1] Note that if no restriction is made on the size of ${\mathcal{X}}$, then the birthday paradox shows that the requirement $k=o(\sqrt{n})$ is tight. In comparison, Theorem 1.5 implies that as long as the exchangeable distribution can be written as a convolution with a kernel ${\mathsf{K}}$ satisfying $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})<\infty$ (such as a Gaussian), the mixtures are statistically indistinguishable for any $k=o(n)$, without any restriction on $|{\mathcal{X}}|$. This observation suggests that the addition of noise in the form of the kernel ${\mathsf{K}}$ has a similar qualitative effect as a bound on the size of ${\mathcal{X}}$; we leave investigation of this interesting connection to future work. A similar upper bound on the KL divergence for the noisy setting also appeared in [TP23, Theorem 2], with a linear dependence on both $\frac{k}{n}$ and the cardinality of ${\mathcal{P}}$; we will provide a detailed comparison in Section 1.3.
The second application concerns the statistical distance between two different permutation mixtures. Given $P_{1},\cdots,P_{n}\in{\mathcal{P}}$, let $Q_{1}\in{\mathcal{P}}$ be an arbitrary distribution, and $Q_{i}=P_{i}$ for all $i\geq 2$. The permutation mixtures $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ are then defined as
$(X_{1},\cdots,X_{n})$ $\sim\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\otimes_{i=1}^{n}P_{\pi(i)}\right]\text{ under }\mathbb{P}_{n};$
$(X_{1},\cdots,X_{n})$ $\sim\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\otimes_{i=1}^{n}Q_{\pi(i)}\right]\text{ under }\mathbb{Q}_{n}.$
In other words, the components of $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ only differ in one coordinate. The following theorem establishes an upper bound on the squared total variation (TV) distance.
Theorem 1.6.
If ${\mathsf{D}}_{\chi^{2}}({\mathcal{P}})<\infty$, then
${\sf TV}(\mathbb{P}_{n},\mathbb{Q}_{n})^{2}\leq\frac{1}{4}\int\frac{(\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{Q}_{n})^{2}}{\mathrm{d}\overline{P}^{\otimes n}}\leq\frac{3{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}{n},$
where $\overline{P}:=\frac{1}{n-1}\sum_{i=2}^{n}P_{i}$. In particular, for every test function $f\in L^{2}(\overline{P}^{\otimes n})$, it holds that
$\left(\mathbb{E}_{\mathbb{P}_{n}}[f]-\mathbb{E}_{\mathbb{Q}_{n}}[f]\right)^{2}\leq\mathbb{E}_{\overline{P}^{\otimes n}}\left[f^{2}\right]\cdot\frac{12{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}{n}.$
Theorem 1.6 shows that the squared TV distance between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ is of the order $O_{\mathcal{P}}(\frac{1}{n})$. In addition, the mean difference of any test function $f$ under $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ is of order $O_{\mathcal{P}}(n^{-1/2})$ as long as the second moment of $f$ under the i.i.d. distribution $\overline{P}^{\otimes n}$ is bounded.
This result turns out to have implications in empirical Bayes models. In the “oracle Bayes” setting of empirical Bayes [GR09, Efr19], a distribution class $(P_{\theta})_{\theta\in\Theta}$ and a set of parameters $\theta:=\{\theta_{1},\cdots,\theta_{n}\}$ are given. The true parameter $\vartheta$ is a uniformly random permutation of $\theta$, i.e. $\vartheta_{i}=\theta_{\pi(i)}$ for all $i\in[n]$ with $\pi\sim\text{\rm Unif}(S_{n})$, and the observation $X=(X_{1},\dots,X_{n})$ follows the distribution $\otimes_{i=1}^{n}P_{\vartheta_{i}}$ conditioned on $\vartheta$. To estimate $\vartheta$, there are two natural types of decision rules to compete with, following the language of [GR09]: one is a simple decision rule, taking the form $\widehat{\vartheta}^{\mathrm{S}}=(\Delta(X_{1}),\dots,\Delta(X_{n}))$ for some fixed function $\Delta$, and the other is a permutation invariant decision rule, satisfying
$\widehat{\vartheta}^{\mathrm{PI}}_{\pi(i)}(X_{\pi(1)},\dots,X_{\pi(n)})=\widehat{\vartheta}^{\mathrm{PI}}_{i}(X_{1},\dots,X_{n}),\quad i\in[n],$
for all $\pi\in S_{n}$. Denote by ${\mathcal{D}}^{\mathrm{S}}$ and ${\mathcal{D}}^{\mathrm{PI}}$ the classes of all simple and permutation invariant decision rules, respectively. Given a separable loss function $L(\vartheta,\widehat{\vartheta})=\sum_{i=1}^{n}\ell(\vartheta_{i},\widehat{\vartheta}_{i})$, we define the optimal risks among all decision rules in ${\mathcal{D}}^{\mathrm{S}}$ and ${\mathcal{D}}^{\mathrm{PI}}$:
$r^{\mathrm{S}}(\theta)=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{S}}}\mathbb{E}\left[L(\vartheta,\widehat{\vartheta})\right],\qquad r^{\mathrm{PI}}(\theta)=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{PI}}}\mathbb{E}\left[L(\vartheta,\widehat{\vartheta})\right].$
As ${\mathcal{D}}^{\mathrm{S}}\subseteq{\mathcal{D}}^{\mathrm{PI}}$, we always have $r^{\mathrm{S}}(\theta)\geq r^{\mathrm{PI}}(\theta)$. Theorem 1.6 implies that the risk difference $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)$ is small (i.e. of the order $o(n)$) in general:
Corollary 1.7.
Let $P_{\theta_{1}},\cdots,P_{\theta_{n}}\in{\mathcal{P}}$ with ${\mathsf{D}}_{\chi^{2}}({\mathcal{P}})<\infty$, and $0\leq\ell(\cdot,\cdot)\leq M$. Then
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)\leq M\sqrt{6n{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}.$
Corollary 1.7 shows that for a large class of models and losses, asymptotically these two types of decision rules achieve similar mean risks $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=O_{{\mathcal{P}}}(\sqrt{n})$, strengthening a “folklore” result $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=o(n)$ in [HR55]. For the quadratic loss $L(\vartheta,\widehat{\vartheta})=\|\vartheta-\widehat{\vartheta}\|_{2}^{2}$, the risk difference can be improved to $O(1)$ via a different strategy, as shown for a class of models in [GR09]; we will discuss this improvement in Section 5.3.
1.1 Notation
Throughout the paper all logarithms are in base $e$. For a positive integer $n$, let $[n]:=\{1,\ldots,n\}$, and $S_{n}$ be the permutation group over $[n]$. For a vector $x$ in $\mathbb{R}^{n}$ or $\mathbb{C}^{n}$, let $x^{\top}$ and $x^{\mathsf{H}}$ be its real and conjugate transposes, respectively, and $e_{\ell}(x):=\sum_{S\subseteq[n]:|S|=\ell}\prod_{i\in S}x_{i}$ be the elementary symmetric polynomial of degree $\ell$ in $x$. For a square matrix $A=(a_{ij})_{i,j\in[n]}\in\mathbb{R}^{n\times n}$, let $\mathrm{Tr}(A)=\sum_{i=1}^{n}a_{ii}$ be its trace, and $\mathrm{Perm}(A)=\sum_{\pi\in S_{n}}\prod_{i=1}^{n}a_{i\pi(i)}$ be its permanent. Let ${\mathcal{N}}(\mu,\sigma^{2})$ denote the normal distribution with mean $\mu$ and variance $\sigma^{2}$, and ${\mathcal{C}}{\mathcal{N}}(0,1)$ denote the complex normal distribution with real and imaginary parts being independent ${\mathcal{N}}(0,\frac{1}{2})$ random variables. For probability measures $\mu$ and $\nu$, let $L^{2}(\mu)$ be the space of all functions $f$ with $\mathbb{E}_{\mu}\left[f^{2}\right]<\infty$, $\mu\otimes\nu$ be the product measure, and $\mu\star\nu$ be the convolution defined as $\mu\star\nu(A)=\int\mu(A-x)\mathrm{d}\nu(x)$. For probability measures $P$ and $Q$ on the same probability space, let
${\sf TV}(P,Q)=\frac{1}{2}\int|\mathrm{d}P-\mathrm{d}Q|,\quad H^{2}(P,Q)=\int\left(\sqrt{\mathrm{d}P}-\sqrt{\mathrm{d}Q}\right)^{2},\quad\mathsf{LC}(P,Q)=\frac{1}{2}\int\frac{(\mathrm{d}P-\mathrm{d}Q)^{2}}{\mathrm{d}P+\mathrm{d}Q}$
be the total variation (TV), squared Hellinger, and Le Cam distances, respectively, and
${\sf KL}(P\|Q)=\int\mathrm{d}P\log\frac{\mathrm{d}P}{\mathrm{d}Q},\qquad\chi^{2}(P\|Q)=\int\frac{(\mathrm{d}P-\mathrm{d}Q)^{2}}{\mathrm{d}Q}$
be the Kullback–Leibler (KL) and $\chi^{2}$ divergences, respectively. A collection of inequalities between the above distances/divergences can be found in [PW24, Chapter 7.6].
We shall use the following standard asymptotic notations. For non-negative sequences $\{a_{n}\}$ and $\{b_{n}\}$, let $a_{n}=O(b_{n})$ denote $\limsup_{n\to\infty}a_{n}/b_{n}<\infty$, and $a_{n}=o(b_{n})$ denote $\limsup_{n\to\infty}a_{n}/b_{n}=0$. In addition, we write $a_{n}=\Omega(b_{n})$ for $b_{n}=O(a_{n})$, $a_{n}=\omega(b_{n})$ for $b_{n}=o(a_{n})$, and $a_{n}=\Theta(b_{n})$ for both $a_{n}=O(b_{n})$ and $b_{n}=O(a_{n})$. We will also use the notations $O_{\theta},o_{\theta}$, etc. to denote that the hidden factor depends on some external parameter $\theta$.
1.2 Proof techniques
The core of our approach is to express the $\chi^{2}$ divergence between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ in terms of matrix permanents. Let us denote the marginal law of the coordinates of $X$ under $\mathbb{Q}_{n}$ by $\overline{P}$. We may then write
$\frac{\mathrm{d}\mathbb{P}_{n}}{\mathrm{d}\mathbb{Q}_{n}}(x_{1},\dots,x_{n})=\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{n}\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(x_{i})\right]=\frac{1}{n!}\mathrm{Perm}(M(x)),$ (9)
where $M:\mathbb{R}^{n}\to\mathbb{R}^{n\times n}$ is given by
$M(x)_{ij}=\frac{\mathrm{d}P_{j}}{\mathrm{d}\overline{P}}(x_{i}).$
Obtaining bounds on matrix permanents is a classical subject [Min84, Chapter 6]. This task is made challenging by the fact that no polynomial time algorithms to compute or approximate matrix permanents are known, which rules out the existence of general purpose bounds on $\mathrm{Perm}(M(x))$. Our strategy is to exploit enough special structure of the matrix $M(x)$ to allow us to compute accurate upper bounds.
We present two approaches to analyzing (9). The first, simpler, approach involves expanding $\mathrm{Perm}(M(x))$ around the $n\times n$ all-ones matrix $J$. This expansion, which we call the “doubly centered” expansion, has the appealing property that the summands are orthogonal in $L^{2}(\mathbb{Q}_{n})$ and involve matrices whose rows sum to zero. We then apply a key inequality, Lemma 3.1, which shows that permanents of matrices with centered rows are small. This inequality is obtained via a new bound on elementary symmetric polynomials: for any vector $x\in\mathbb{R}^{n}$, if $\sum_{i=1}^{n}x_{i}=0$, then
$|e_{\ell}(x)|\leq C\sqrt{\binom{n}{\ell}}\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}\right)^{\ell/2}.$ (10)
This bound may be contrasted with Maclaurin’s inequality [Mac30], a version of which implies the bound $|e_{\ell}(x)|\leq\binom{n}{\ell}\left(\frac{1}{n}\sum_{i=1}^{n}x_{i}^{2}\right)^{\ell/2}$ for all $x\in\mathbb{R}^{n}$, which is easily seen to be tight when $x_{1}=\dots=x_{n}=1$. The crucial difference between this classical bound and our new inequality is the square-root improvement in the leading coefficient from $\binom{n}{\ell}$ to $\sqrt{\binom{n}{\ell}}$, which may be viewed as a consequence of the cancellations induced by the condition $\sum_{i=1}^{n}x_{i}=0$. Our new Maclaurin-type inequality, combined with the doubly centered expansion, gives rise to the first bound in Theorem 1.2.
Our second approach combines the above observations with additional spectral information. Since $(X_{1},\dots,X_{n})$ are independent under $\mathbb{Q}_{n}$, we have that for any $\pi,\pi^{\prime}\in S_{n}$,
$\mathbb{E}_{X\sim\mathbb{Q}_{n}}\left[\prod_{i=1}^{n}M_{i,\pi(i)}(X_{i})\prod_{i=1}^{n}M_{i,\pi^{\prime}(i)}(X_{i})\right]$ $=\prod_{i=1}^{n}\mathbb{E}_{X_{i}\sim\overline{P}}\left[M_{i,\pi(i)}(X_{i})M_{i,\pi^{\prime}(i)}(X_{i})\right]$
$=\prod_{i=1}^{n}\int\frac{\mathrm{d}P_{\pi(i)}\mathrm{d}P_{\pi^{\prime}(i)}}{\mathrm{d}\overline{P}}=:n^{n}\prod_{i=1}^{n}A_{\pi(i),\pi^{\prime}(i)},$
where we write $A\in\mathbb{R}^{n\times n}$ for the matrix given by $A_{ij}=\frac{1}{n}\int\frac{\mathrm{d}P_{i}\mathrm{d}P_{j}}{\mathrm{d}\overline{P}}$. We obtain
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})+1$ $=\frac{1}{(n!)^{2}}\mathbb{E}_{X\sim\mathbb{Q}_{n}}\mathrm{Perm}(M(X))^{2}=\frac{n^{n}}{n!}\mathrm{Perm}(A)\,.$
The matrix $A$ is doubly stochastic, therefore putting the question of bounding $\mathrm{Perm}(A)$ in the setting of the celebrated van der Waerden conjecture [vdW26], proved by Egorychev [Ego81] and Falikman [Fal81], which states that $\mathrm{Perm}(A)\geq\frac{n!}{n^{n}}$ for all $n\times n$ doubly stochastic matrices $A$. We prove a new upper bound on the permanent of such matrices: if $A$ is a doubly stochastic matrix with eigenvalues $1=\lambda_{1}>\lambda_{2}\geq\dots\geq\lambda_{n}\geq 0$, then
$\mathrm{Perm}(A)\leq\frac{n!}{n^{n}}\prod_{i=2}^{n}\frac{1}{1-\lambda_{i}}.$
The second and third bounds in Theorem 1.2 are obtained by relating the spectrum of $A$ to the information theoretic quantities in Definition 1.1.
Both applications in Theorem 1.5 and 1.6 rely on a combination of the above two approaches, by first expanding $M(x)$ around the all-ones matrix $J$ and then relating each term in the expansion to a counterpart appearing in the expression of $\mathrm{Perm}(A)$, through the identities in Appendix A. In particular, a complex version of the inequality (10) for $x\in\mathbb{C}^{n}$ turns out to be crucial in the proof of Theorem 1.6, a technique which could be of independent interest.
1.3 Related work
Comparisons between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ first implicitly arose in the consideration of compound decision problems. An early result in this direction is due to Hannan [Han53] (see [HR55]). In our notation, his results reads as follows: suppose that ${\mathcal{P}}=\{P_{-},P_{+}\}$ consists of two elements, which are not orthogonal to each other. Given natural numbers $n$ and $k$ with $k\leq n$, denote by $\mathbb{P}^{(k)}_{n}$ the permutation mixture arising from taking $P_{1}=\dots=P_{k}=P_{+}$ and $P_{k+1}=\dots=P_{n}=P_{-}$. If $k/n$ is bounded away from 00 and $1$ as $n,k\to\infty$, then
${\sf TV}(\mathbb{P}_{n}^{(k)},\mathbb{P}_{n}^{(k+1)})=o(1)\,.$
In [HR55], this result is used to establish asymptotic equivalence between the performance of permutation invariant and simple decision rules. Subsequent empirical Bayes literature largely focused on simple decision rules and avoided explicit consideration of $\mathbb{P}_{n}$, but the importance of comparing $\mathbb{P}_{n}$ to $\mathbb{Q}_{n}$ to obtain genuine oracle inequalities was emphasized by Greenshtein and Ritov [GR09, GR19], who obtained risk bounds for the two models under the square loss. The general power of permutation-invariant decision rules was recently investigated in [Wei21].
The fact that exchangeable distributions are “essentially” (mixtures of) product measures is an important heuristic which has its roots in the work of de Finetti [DF29] and Hewitt–Savage [HS55]. Diaconis and Friedman observed that versions of this claim hold in a number of different models [DF87], and precise quantitative versions have been established for finite exchangeable sequences under different assumptions [Sta78, DF80, GK21, JGK24]. Recently, motivated by applications in conformal prediction (see, e.g., [TBCR19, BCRT23a]), similar results have been obtained for “weighted” exchangeable sequences as well [BCRT23b].
Bounding the $\chi^{2}$ divergence for mixture distributions is a common task in high-dimensional statistics [LNS99, Bar02, CL11, IS12, WY16, CCT17, BMV+18, JHW18, PWBM18, BW19, BNWR20], and obtaining bounds for “low-degree” versions of the $\chi^{2}$ divergence has become an important technique in establishing statistical-computational gaps (see, e.g., [KWB19]). In most of these applications, $\mathbb{P}_{n}$ is a mixture distribution but $\mathbb{Q}_{n}$ is particularly simple (for instance, its coordinates are i.i.d. ${\mathcal{N}}(0,1)$ or $\text{\rm Bern}(1/2)$); in this case, expanding the likelihood ratio $\frac{\mathrm{d}\mathbb{P}_{n}}{\mathrm{d}\mathbb{Q}_{n}}$ in a basis of $L^{2}(\mathbb{Q}_{n})$-orthogonal polynomials yields explicit bounds in terms of the moments of $\mathbb{P}_{n}$. Few general techniques exist to obtain sharp results outside of this simple setting, though Schramm and Wein [SW22] (see also [RSWY23]) showed that a related expansion based on cumulants yields a (potentially loose) bound when $\mathbb{Q}_{n}$ is a mixture distribution. In a more general context, Kunisky [Kun24] has shown that employing the Hoeffding decomposition [Hoe48] can sometimes obtain comparable results for more complicated choices of $\mathbb{Q}_{n}$. We employ a similar technique in Section 3.
Bounds for permutation mixtures were first obtained by Ding [Din22] in a PhD thesis supervised by the second author. He established contiguity in the setting of the example presented in (1), by showing
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})=\begin{cases}O(\mu^{2})&\text{if }\mu\leq 1,\\ O(\exp(4\mu^{2}))&\text{if }\mu>1.\end{cases}$
He also developed analogous bounds for a Bernoulli version of the problem. A dimension-independent upper bound on the KL divergence ${\sf KL}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ was also observed and obtained in [TP23, Theorem 2] for a finite class ${\mathcal{P}}$, used as a crucial technical step towards establishing the capacity upper bound of the noisy permutation channel in [Mak20]. Their proof relies on an anti-concentration property for the histogram of independent discrete observations with non-identical distributions, leading to an upper bound linear in $|{\mathcal{P}}|$. They also prove a bound analogous to the one presented in Theorem 1.5 for $k$-dimensional marginals with $k<n$ via a convexity argument [TP23, Appendix A]; however, their bound scales as $O(\frac{k}{n})$ rather than $O(\frac{k^{2}}{n^{2}})$.
1.4 Organization
The rest of this paper is organized as follows. In Section 2, we review some existing methods of bounding the $\chi^{2}$ divergence between mixture distributions, and show their failure to obtain the tight bound (1) for the toy Gaussian model at the beginning of the introduction. Section 3 and 4 provide the details of the doubly centered expansion and the matrix permanent approach, respectively, where a key inequality in the analysis is a new upper bound on the elementary symmetric polynomial for centered real or complex vectors in Section 3.3. Section 5 provides more discussions and applications of our results, and we defer further discussions on the relationship between the two approaches (Appendix A), proofs of the new inequality (Appendix B) and other main results (Appendix C) to the appendix.
2 Failure of existing approaches
Before we present our proof of Theorem 1.2, we review several existing approaches to upper bound the statistical distance between mixture distributions. Unfortunately, we will show that all of them fail to yield the bound (1). For these approaches, occasionally we will show the failure under other statistical distances such as the squared TV distance or the KL divergence, both of which are no larger than the $\chi^{2}$ divergence.
2.1 Reduction to two simple distributions
The simplest strategy to deal with two mixture distributions is a reduction to “simple” distributions via coupling and convexity. Taking the KL divergence for an example and returning to the setting of (1), it holds that
${\sf KL}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ $={\sf KL}(\mathbb{E}_{\vartheta\sim\nu_{\mathbb{P}}}[{\mathcal{N}}(\vartheta,I_{n})]\|\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}[{\mathcal{N}}(\vartheta^{\prime},I_{n})])$
$\leq\min_{\rho\in\Pi(\nu_{\mathbb{P}},\nu_{\mathbb{Q}})}\mathbb{E}_{(\vartheta,\vartheta^{\prime})\sim\rho}\left[{\sf KL}({\mathcal{N}}(\vartheta,I_{n})\|{\mathcal{N}}(\vartheta^{\prime},I_{n}))\right]=\frac{W_{2}^{2}(\nu_{\mathbb{P}},\nu_{\mathbb{Q}})}{2},$
where the inequality is due to the joint convexity of the KL divergence, $\Pi(\nu_{\mathbb{P}},\nu_{\mathbb{Q}})$ denotes all possible couplings with marginals $\nu_{\mathbb{P}}$ and $\nu_{\mathbb{Q}}$, and $W_{2}$ is the Wasserstein-$2$ distance. For the last term, note that $\theta\mapsto|\sum_{i=1}^{n}\theta_{i}|$ is $1$-Lipschitz under $\|\cdot\|_{2}$, so we can lower bound it as
$W_{2}(\nu_{\mathbb{P}},\nu_{\mathbb{Q}})\geq W_{1}(\nu_{\mathbb{P}},\nu_{\mathbb{Q}})\geq\mathbb{E}_{\vartheta\sim\nu_{\mathbb{Q}}}\left|\sum_{i=1}^{n}\vartheta_{i}\right|-\mathbb{E}_{\vartheta\sim\nu_{\mathbb{P}}}\left|\sum_{i=1}^{n}\vartheta_{i}\right|=\mu\mathbb{E}\sqrt{\left|2\text{\rm B}\left(n,\frac{1}{2}\right)-n\right|}=\Omega\left(\mu n^{\frac{1}{4}}\right),$
where $\text{\rm B}(\cdot,\cdot)$ denotes the Binomial distribution, and the last step is due to CLT. Consequently, this coupling approach can at best provide an upper bound of $O(\sqrt{n}\mu^{2})$, which not only grows with $n$ but also exhibits a loose dependence on $\mu$ (i.e. $O(\mu^{2})$ instead of $O(\mu^{4})$ for small $\mu$).
2.2 Reduction to one simple distribution
Instead of reducing to the $\chi^{2}$ divergence between two simple distributions, a more careful coupling approach could reduce to the $\chi^{2}$ divergence between a mixture distribution and a simple distribution, where the standard second moment computation [IS12] could then be applied. Still taking the KL divergence for an example, such a general strategy takes the form
${\sf KL}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ $\leq\min_{\{\nu_{\theta^{\prime}}\}_{\theta^{\prime}\in\{\pm\mu\}^{n}}}\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}\left[{\sf KL}\left(\mathbb{E}_{\vartheta\sim\nu_{\vartheta^{\prime}}}\left[{\mathcal{N}}(\vartheta,I_{n})\right]\|{\mathcal{N}}(\vartheta^{\prime},I_{n})\right)\right],$
where the minimization is over all possible families of distributions $\{\nu_{\theta^{\prime}}\}_{\theta^{\prime}\in\{\pm\mu\}^{n}}$ such that $\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}[\nu_{\vartheta^{\prime}}]=\nu_{\mathbb{P}}$. This is the strategy developed in [Din22], where it is employed with a judicious choice of the family $\{\nu_{\theta^{\prime}}\}_{\theta^{\prime}\in\{\pm\mu\}^{n}}$ to show that $\mathbb{P}_{n}$ is contiguous to $\mathbb{Q}_{n}$.
However, even though it succeeds in proving a dimension-free upper bound of ${\sf KL}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$, this strategy cannot yield the correct dependence on $\mu$. Indeed, we have the following lower bound:
$\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}\left[{\sf KL}\left(\mathbb{E}_{\vartheta\sim\nu_{\vartheta^{\prime}}}\left[{\mathcal{N}}(\vartheta,I_{n})\right]\|{\mathcal{N}}(\vartheta^{\prime},I_{n})\right)\right]$
$\overset{\rm(a)}{\geq}\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}\left[W_{2}^{2}\left(\mathbb{E}_{\vartheta\sim\nu_{\vartheta^{\prime}}}\left[{\mathcal{N}}(\vartheta,I_{n})\right],{\mathcal{N}}(\vartheta^{\prime},I_{n})\right)\right]$
$\overset{\rm(b)}{\geq}\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}\left[\left\|\mathbb{E}_{\vartheta\sim\nu_{\vartheta^{\prime}}}[\vartheta]-\vartheta^{\prime}\right\|_{2}^{2}\right]\overset{\rm(c)}{\geq}\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}\left[\frac{1}{n}\left(\sum_{i=1}^{n}\vartheta_{i}^{\prime}\right)^{2}\right]=\mu^{2},$
where (a) is the transportation cost inequality under the Gaussian measure ${\mathcal{N}}(\theta^{\prime},I_{n})$ [Led01, Theorem 6.6], (b) uses the inequality $W_{2}^{2}(P,Q)\geq\|\mathbb{E}_{P}[X]-\mathbb{E}_{Q}[X]\|_{2}^{2}$ due to convexity, and (c) notes that the vector $\mathbb{E}_{\vartheta\sim\nu_{\vartheta^{\prime}}}[\vartheta]$ always lies on the hyperplane $\{\theta\in\mathbb{R}^{n}:\sum_{i=1}^{n}\theta_{i}=0\}$ and uses the orthogonal projection onto this hyperplane. Therefore, this approach can at best lead to an upper bound $O(\mu^{2})$, still larger than the correct dependence $O(\mu^{4})$ for small $\mu$. On a high level, this is because the original mixtures $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ have the same mean, but after applying the convexity, the means of $\mathbb{E}_{\theta\sim\nu_{\theta^{\prime}}}\left[{\mathcal{N}}(\theta,I_{n})\right]$ and ${\mathcal{N}}(\theta^{\prime},I_{n})$ no longer match.
2.3 Method of moments
The previous failures illustrate the importance of directly comparing the mixtures $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ rather reducing one or both measures to simple distributions. The method of moments is a powerful approach for performing such a direct comparison. Expanding the likelihood ratio in the orthogonal basis of Hermite polynomials shows that (see, e.g. [HP15, Lemma 2.2] and [WY20, Lemma 9])
${\sf TV}(\nu_{\mathbb{P}}\star{\mathcal{N}}(0,I_{n}),\nu_{\mathbb{Q}}\star{\mathcal{N}}(0,I_{n}))^{2}\leq\sum_{\alpha\geq 0}\frac{(m_{\alpha}(\nu_{\mathbb{P}})-m_{\alpha}(\nu_{\mathbb{Q}}))^{2}}{\alpha!},$ (11)
where $\alpha=(\alpha_{1},\cdots,\alpha_{n})\in\mathbb{N}^{n}$ is a multi-index, $\alpha!\triangleq\prod_{i=1}^{n}(\alpha_{i}!)$, and $m_{\alpha}(\nu)=\mathbb{E}_{X\sim\nu}\left[\prod_{i=1}^{n}X_{i}^{\alpha_{i}}\right]$ denotes the $\alpha$-th joint moment of $X\sim\nu$. Since the moment difference in (11) vanishes when $|\alpha|=\sum_{i=1}^{n}\alpha_{i}=1$, the method of moments indeed suggests an $O(\mu^{4})$ dependence on $\mu$.
However, the bound in (11) is not dimension free and tends to infinity as $n\to\infty$. To see this, fix some $\ell\in\mathbb{N}$ and consider the multi-index $\alpha=(1,1,2,\cdots,2,0,\cdots,0)$ with $2$ appearing $\ell$ times. By simple algebra, it is easy to check that
$m_{\alpha}(\nu_{\mathbb{P}})=\mu^{2\ell}\mathbb{E}_{\vartheta\sim\nu_{\mathbb{P}}}[\vartheta_{1}\vartheta_{2}]=-\frac{\mu^{2\ell+2}}{n-1},\qquad m_{\alpha}(\nu_{\mathbb{Q}})=0.$
Therefore, the total contribution of all permutations of $\alpha$ to the RHS of (11) is
$\frac{\mu^{4\ell+4}}{2^{\ell}(n-1)^{2}}\cdot\binom{n}{2,\ell,n-\ell-2}=\Theta_{\ell}\left(\mu^{4\ell+4}n^{\ell}\right),$
which grows polynomially with $n$ for any constant $\ell\geq 1$. On a high level, this shows that although the method of moments works well in one dimension, the sum of squared moment differences might become unbounded in high dimensions due to a large number of cross terms.
2.4 Method of cumulants
To address a similar problem arising in the recovery problem under a low-degree framework, a recent work [SW22] established an upper bound on the $\chi^{2}$ divergence based on cumulants. Specifically, it was shown in [SW22, Theorem 2.2] and [RSWY23, Proposition 2.1] that
$\chi^{2}(\nu_{\mathbb{P}}\star{\mathcal{N}}(0,I_{n})\|\nu_{\mathbb{Q}}\star{\mathcal{N}}(0,I_{n}))\leq\sum_{\alpha\geq 0}\frac{\kappa_{\alpha}^{2}}{\alpha!},$ (12)
where $\kappa_{\alpha}$ denotes the joint cumulant
$\kappa_{\alpha}=\kappa_{\vartheta\sim\nu_{\mathbb{Q}}}\left(\frac{\mathrm{d}\nu_{\mathbb{P}}}{\mathrm{d}\nu_{\mathbb{Q}}}(\vartheta),\vartheta_{1},\ldots,\vartheta_{1},\vartheta_{2},\ldots,\vartheta_{2},\ldots,\vartheta_{n}\right),$
where $\vartheta_{i}$ appears $\alpha_{i}$ times in the joint cumulant. We refer to [SW22] for an in-depth discussion of cumulants; we shall only need the following recursive formula to compute $\kappa_{\alpha}$:
$\kappa_{\alpha}=\mathbb{E}_{\vartheta\sim\nu_{\mathbb{P}}}\left[\prod_{i=1}^{n}\vartheta_{i}^{\alpha_{i}}\right]-\sum_{0\leq\beta\lneq\alpha}\kappa_{\beta}\prod_{i=1}^{n}\left[\binom{\alpha_{i}}{\beta_{i}}\mathbb{E}_{\vartheta_{i}\sim\text{\rm Unif}(\{\pm\mu\})}\left[\vartheta_{i}^{\alpha_{i}-\beta_{i}}\right]\right].$ (13)
Although the upper bound (12) enjoys several advantages over the moment-based approach such as a better behavior for product distributions, it still fails in our problem in a delicate way. Consider a multi-index $\alpha=(1,\ell,0,\ldots,0)$ and let $a_{\ell}=\kappa_{(1,\ell,0,\ldots,0)}$. Based on (13) and some algebra, it is easy to show $a_{2\ell}=0$ for all $\ell\in\mathbb{N}$, and that the modified sequence $b_{2\ell+1}=(-1)^{\ell+1}(n-1)a_{2\ell+1}/\mu^{2\ell+2}$ satisfies
$b_{2\ell+1}-\binom{2\ell+1}{2}b_{2\ell-1}+\binom{2\ell+1}{4}b_{2\ell-3}-\cdots=(-1)^{\ell},\qquad b_{1}=1.$
The sequence $\{b_{n}\}$ is well known to be the number of alternating permutations, [^2] Entry A000182 in The On-Line Encyclopedia of Integer Sequences, https://oeis.org/A000182 with the asymptotic growth [FS09, Page 5]
$b_{2\ell+1}\sim 2\left(\frac{2}{\pi}\right)^{2\ell+1}\cdot(2\ell+1)!\qquad\text{as }\ell\to\infty.$
We obtain that
$\kappa_{(1,2\ell+1,0,\ldots,0)}=a_{2\ell+1}\sim(-1)^{\ell+1}\frac{\pi}{n-1}\left(\frac{2\mu}{\pi}\right)^{2\ell+2}\cdot(2\ell+1)!\qquad\text{as}\ell\to\infty.$
As the growth of $(2\ell+1)!$ is much faster than exponential, summing along this subsequence in (12) gives a diverging result, indicating the failure of this approach.
3 Upper bound via the doubly centered expansion
In this section we prove the first part (4) of Theorem 1.2, by developing an orthogonal expansion of the likelihood ratio in terms of a set of “doubly centered” functions. In Section 3.1 we first provide a direct proof of the previous toy example and discuss the intuitions, and then present the proof for the general case in Section 3.2. One key step is a new upper bound of elementary symmetric polynomials for a sequence summing into zero, which we present in Section 3.3.
3.1 A warm-up example
Similar to Section 2, this section focuses on the toy Gaussian model in (1) as a warm-up example. The proof relies critically on the following representation of the Gaussian likelihood ratio:
$\frac{\varphi(x-\theta)}{\varphi(x)}=\exp\left(\theta x-\frac{\theta^{2}}{2}\right)=\cosh(\mu x)\exp\left(-\frac{\mu^{2}}{2}\right)+\sinh(\mu x)\frac{\theta}{\mu}\exp\left(-\frac{\mu^{2}}{2}\right),\quad\theta\in\{\pm\mu\},$ (14)
where $\varphi(x)$ is the density function of ${\mathcal{N}}(0,1)$. In addition, the marginal distribution under $\mathbb{Q}_{n}$ is
$\varphi_{0}(x):=\frac{\varphi(x-\mu)+\varphi(x+\mu)}{2}=\varphi(x)\cosh(\mu x)\exp\left(-\frac{\mu^{2}}{2}\right),$ (15)
corresponding to the first term on the RHS of (14). Based on (14) and (15), we can express the likelihood ratio between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ as
$\frac{\mathrm{d}\mathbb{P}_{n}}{\mathrm{d}\mathbb{Q}_{n}}(X)$ $=\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{n}\frac{\varphi(X_{i}-\theta_{\pi(i)})}{\varphi_{0}(X_{i})}\right]$
=(14)${=}\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{n}\left(1+\tanh(\mu X_{i})\frac{\theta_{\pi(i)}}{\mu}\right)\right]$
$=\sum_{S\subseteq[n]}\left(\prod_{i\in S}\tanh(\mu X_{i})\right)\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i\in S}\frac{\theta_{\pi(i)}}{\mu}\right].$ (16)
Consequently, the likelihood ratio in (3.1) is decomposed into a sum over all subsets $S\subseteq[n]$, and the effects of $X$ and $\pi$ are decoupled in each summand. Some observations are in order:
  1. 1.
    The functions $\{\prod_{i\in S}\tanh(\mu X_{i}):S\subseteq[n]\}$ are orthogonal under $\mathbb{Q}_{n}$. To see this, let $S\neq T$ be two different subsets of $[n]$, so that there exists some $i_{0}\in S\Delta T$. As
[…]
    we have by the product structure of $\mathbb{Q}_{n}$ that
    $\mathbb{E}_{\mathbb{Q}_{n}}\left[\prod_{i\in S}\tanh(\mu X_{i})\prod_{i\in T}\tanh(\mu X_{i})\right]$
    $=\mathbb{E}_{\mathbb{Q}_{n}}\left[\tanh(\mu X_{i_{0}})\right]\cdot\mathbb{E}_{\mathbb{Q}_{n}}\left[\prod_{i\in S\backslash\{i_{0}\}}\tanh(\mu X_{i})\prod_{i\in T\backslash\{i_{0}\}}\tanh(\mu X_{i})\right]=0.$
  2. 2.
    The second moment of $\prod_{i\in S}\tanh(\mu X_{i})$ under $\mathbb{Q}_{n}$ admits a simple expression. In fact, by the i.i.d. structure of $\mathbb{Q}_{n}$, we have
    $\mathbb{E}_{\mathbb{Q}_{n}}\left[\prod_{i\in S}\tanh^{2}(\mu X_{i})\right]=\left(\exp\left(-\frac{\mu^{2}}{2}\right)\int_{\mathbb{R}}\frac{\sinh^{2}(\mu x)}{\cosh(\mu x)}\varphi(x)\mathrm{d}x\right)^{|S|}=:f(\mu)^{|S|}.$
    We can also derive an upper bound of $f(\mu)$:
    $f(\mu)$ $=\exp\left(-\frac{\mu^{2}}{2}\right)\int_{\mathbb{R}}\left(\cosh(\mu x)-\frac{1}{\cosh(\mu x)}\right)\varphi(x)\mathrm{d}x$
    $=1-\exp\left(-\frac{\mu^{2}}{2}\right)\int_{\mathbb{R}}\frac{\varphi(x)}{\cosh(\mu x)}\mathrm{d}x$
    $\leq 1-\exp\left(-\frac{\mu^{2}}{2}\right)\frac{1}{\int_{\mathbb{R}}\cosh(\mu x)\varphi(x)\mathrm{d}x}=1-\exp\left(-\mu^{2}\right).$ (17)
  3. 3.
    The expectation $\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i\in S}\frac{\theta_{\pi(i)}}{\mu}\right]$ depends only on $|S|$, and can be explicitly computed in the toy example. In fact, the generating function $\prod_{i=1}^{n}\left(1+z\theta_{i}/\mu\right)=(1-z^{2})^{n/2}$ gives
    $\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i\in S}\frac{\theta_{\pi(i)}}{\mu}\right]=(-1)^{\ell/2}\frac{\binom{n/2}{\ell/2}}{\binom{n}{\ell}}{\mathbf{1}_{\left\{{\ell\text{ is even}}\right\}}}=:g_{\ell},$ (18)
    where $\ell=|S|$.
    In particular, since the mean vector is balanced, $g_{1}=\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\frac{\theta_{\pi(1)}}{\mu}\right]=0$, and we have the simple inequality:
    $g_{\ell}^{2}\leq\frac{1}{\binom{n}{\ell}}.$
Based on the above observations, we are in a position to compute the second moment of (3.1):
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ $=\mathbb{E}_{\mathbb{Q}_{n}}\left[\left(\frac{\mathrm{d}\mathbb{P}_{n}}{\mathrm{d}\mathbb{Q}_{n}}\right)^{2}\right]-1=\sum_{S\subseteq[n]}f(\mu)^{|S|}g_{|S|}^{2}-1\overset{\rm(a)}{=}\sum_{S\subseteq[n]:|S|\geq 2}f(\mu)^{|S|}g_{|S|}^{2}$
$\leq\sum_{\ell=2}^{n}\binom{n}{\ell}\cdot\frac{f(\mu)^{\ell}}{\binom{n}{\ell}}\leq\frac{f(\mu)^{2}}{1-f(\mu)}=\begin{cases}O(\mu^{4})&\text{if }\mu\leq 1,\\ O(\exp(\mu^{2}))&\text{if }\mu>1.\end{cases}$
where in (a), the contribution of $S=\varnothing$ cancels with $-1$, and $g_{1}=0$ is used for $|S|=1$. This proves (1).
We make some comments on the intuition behind this approach and the challenges moving forward. The methods of moments and cumulants discussed in Section 2 rely on expanding the ratio $\varphi(x-\theta)/\varphi(x)$ in the basis of Hermite polynomials, which are orthogonal under $L^{2}(\varphi)$; this expansion facilitates taking expectations with respect to the standard Gaussian measure—as is done implicitly in deriving (11) and (12)—but computing the $\chi^{2}$ divergence between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ by comparing both measures to the standard Gaussian is inherently loose.
By contrast, the above approach works directly with the ratio $\varphi(x-\theta)/\varphi_{0}(x)=1+\Psi(x,\theta)$, where $\Psi(x,\theta):=\tanh(\mu x)\frac{\theta}{\mu}$. This representation has two benefits with respect to an expansion in Hermite polynomials. First, the product structure of $\mathbb{Q}_{n}$ guarantees that for any $\theta_{i}\in\{\pm\mu\}$, the functions $\{\Pi_{i\in S}\Psi(X_{i},\theta_{i}):S\subseteq[n]\}$ are automatically orthogonal in $L^{2}(\mathbb{Q}_{n})$, so no change of measure is required. Second, and more subtly, for any $x\in\mathbb{R}^{d}$, the function $\theta\mapsto\Psi(x,\theta)$ is automatically mean-zero when $\theta\sim\mathrm{Unif}(\{\pm\mu\})$. This property guarantees that the first-order term in (3.1) vanishes, which is crucial to obtaining the correct dependence on $\mu$, and indirectly ensures that the coefficients $g_{\ell}$ are typically small. On the other hand, as Section 2 makes clear, large values of the moments $m_{\alpha}$ or cumulants $\kappa_{\alpha}$ lead to the failure of the moment and cumulant methods. Further comparisons between the above two approaches are discussed in Section 5.4.
Several challenges arise when attempting to implement this idea in general. There will typically not be a simple expression for the analogue of $\Psi(x,\theta)$ in the general case, and the coefficients no longer admit an explicit formula such as (18). We address these challenges in Section 3.2.
3.2 Proof of the general case
This section is devoted to the proof of the first part (4) of Theorem 1.2. Motivated by (14) and (15) in the warm-up example, we write
$\overline{P}(\mathrm{d}x)$ $=\frac{1}{n}\sum_{i=1}^{n}P_{i}(\mathrm{d}x),$ (19)
$P_{i}(\mathrm{d}x)$ $=\overline{P}(\mathrm{d}x)+\Psi_{i}(\mathrm{d}x).$ (20)
It is clear that $\Psi_{i}\ll\overline{P}$, so the derivative $\frac{\mathrm{d}\Psi_{i}}{\mathrm{d}\overline{P}}$ exists. In addition, $\{\Psi_{1},\cdots,\Psi_{n}\}$ have the following crucial “doubly centered” property:
$\int\Psi_{i}(\mathrm{d}x)=\int(P_{i}(\mathrm{d}x)-\overline{P}(\mathrm{d}x))=0,\qquad\text{for all }i;$ (21)
$\sum_{i=1}^{n}\Psi_{i}(\mathrm{d}x)=\sum_{i=1}^{n}(P_{i}(\mathrm{d}x)-\overline{P}(\mathrm{d}x))=0,\qquad\text{for }\overline{P}\text{-a.e. }x.$ (22)
As we shall see, the first of these equations guarantees that products functions in $\{\Psi_{i}\}_{i\in[n]}$ are orthogonal in $L^{2}(\mathbb{Q}_{n})$, whereas the second guarantees that $\mathbb{E}_{I\sim\mathrm{Unif}[n]}\Psi_{I}(\mathrm{d}x)=0$. Based on (19) and (20), we express the likelihood ratio as
$\frac{\mathrm{d}\mathbb{P}_{n}}{\mathrm{d}\mathbb{Q}_{n}}(X)$ $=\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{n}\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\right]=\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{n}\left(1+\frac{\mathrm{d}\Psi_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\right)\right]$
$=\sum_{S\subseteq[n]}\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i\in S}\frac{\mathrm{d}\Psi_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\right].$
Similar to the warm-up example, the condition (21) ensures that the above summands are orthogonal under $\mathbb{Q}_{n}=\overline{P}^{\otimes n}$. Therefore,
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})=\mathbb{E}_{\mathbb{Q}_{n}}\left[\left(\frac{\mathrm{d}\mathbb{P}_{n}}{\mathrm{d}\mathbb{Q}_{n}}\right)^{2}\right]-1=\sum_{S\subseteq[n],S\neq\varnothing}\mathbb{E}_{\mathbb{Q}_{n}}\left[\left(\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i\in S}\frac{\mathrm{d}\Psi_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\right]\right)^{2}\right].$ (23)
To proceed, we fix any set $S$ with $|S|=\ell$; by symmetry we assume that $S=[\ell]$. For a fixed vector $(X_{1},\cdots,X_{\ell})$, construct a matrix $A=(a_{ij})\in\mathbb{R}^{\ell\times n}$, with $a_{ij}=\frac{\mathrm{d}\Psi_{j}}{\mathrm{d}\overline{P}}(X_{i})$. It is clear that
$\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i\in S}\frac{\mathrm{d}\Psi_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\right]=\frac{1}{\ell!\binom{n}{\ell}}\sum_{T\subseteq[n],|T|=\ell}\mathrm{Perm}(A_{T}),$
where $A_{T}\in\mathbb{R}^{\ell\times\ell}$ is the submatrix of $A$ by taking the columns of $A$ with indices in $T$. In addition, thanks to (22), the matrix $A$ has all row sums zero. The key to upper bounding (23) is the following technical lemma.
Lemma 3.1.
Let $A=(a_{ij})\in\mathbb{R}^{\ell\times n}$ be a real matrix with $1\leq\ell\leq n$ and all row sums being zero. Then the following inequality holds:
$\left|\frac{1}{\ell!}\sum_{T\subseteq[n],|T|=\ell}\mathrm{Perm}(A_{T})\right|\leq\sqrt{10\binom{n}{\ell}}\cdot\prod_{i=1}^{\ell}\left(\frac{1}{n}\sum_{j=1}^{n}a_{ij}^{2}\right)^{\frac{1}{2}}.$
In the special case $\ell=n$, Lemma 3.1 coincides with the Hadamard-type inequality for permanents [CLL06], even without the condition that all row sums are zero. However, for $\ell<n$, this condition becomes crucial: without this condition, if $A$ is the all-one matrix, the LHS of Lemma 3.1 would be a much larger quantity $\binom{n}{\ell}$. The proof of Lemma 3.1 turns out to be involved and is the central theme of Section 3.3.
Applying Lemma 3.1 to (23), for a given subset $S$ with $|S|=\ell\geq 2$ we get
$\mathbb{E}_{\mathbb{Q}_{n}}\left[\left(\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i\in S}\frac{\mathrm{d}\Psi_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\right]\right)^{2}\right]\leq\frac{10}{\binom{n}{\ell}}\mathbb{E}_{\mathbb{Q}_{n}}\left[\prod_{i=1}^{\ell}\left(\frac{1}{n}\sum_{j=1}^{n}\frac{\mathrm{d}\Psi_{j}}{\mathrm{d}\overline{P}}(X_{i})^{2}\right)\right]$
$=\frac{10}{\binom{n}{\ell}}\left(\mathbb{E}_{Z\sim\overline{P}}\left[\frac{1}{n}\sum_{j=1}^{n}\frac{\mathrm{d}\Psi_{j}}{\mathrm{d}\overline{P}}(Z)^{2}\right]\right)^{\ell}=\frac{10}{\binom{n}{\ell}}\left(\frac{1}{n}\sum_{j=1}^{n}\chi^{2}(P_{j}\|\overline{P})\right)^{\ell}\leq\frac{10}{\binom{n}{\ell}}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{\ell}.$
In addition, for $\ell=1$ the LHS of Lemma 3.1 is zero since all row sums are zero. Therefore, (23) gives that
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq\sum_{S\subseteq[n]:|S|\geq 2}\frac{10}{\binom{n}{|S|}}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{|S|}=\sum_{\ell=2}^{n}\binom{n}{\ell}\cdot\frac{10}{\binom{n}{\ell}}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{\ell}=10\sum_{\ell=2}^{n}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{\ell},$
which is the first part (4) of Theorem 1.2.
3.3 A key inequality
This section is devoted to the proof of Lemma 3.1, which consists of several steps. First we show that it suffices to prove Lemma 3.1 with identical rows. Let $r_{1},\cdots,r_{\ell}\in{\mathcal{H}}$ be the rows of $A$, where =${\mathcal{H}}=\{x\in\mathbb{R}^{n}:\sum_{i=1}^{n}x_{i}=0\}$ is a Hilbert space equipped with the Euclidean inner product. Writing the LHS of Lemma 3.1 as a function of $(r_{1},\dots,r_{\ell})$, it is clear that
$L(r_{1},\dots,r_{\ell}):=\frac{1}{\ell!}\sum_{T\subseteq[n],|T|=\ell}\mathrm{Perm}(A_{T})$
is multilinear in $(r_{1},\ldots,r_{\ell})$. We invoke the following deep result due to S. Banach [Ban38].
Lemma 3.2.
Let $L(x_{1},\cdots,x_{n})$ be a symmetric multilinear form from a Hilbert space $({\mathcal{H}},|\cdot|)$ to either $\mathbb{R}$ or $\mathbb{C}$. Suppose that
$\sup\left\{|L(x,x,\ldots,x)|:|x|\leq 1\right\}\leq M.$
Then it also holds that
$\sup\left\{|L(x_{1},x_{2},\ldots,x_{n})|:|x_{1}|\leq 1,\ldots,|x_{n}|\leq 1\right\}\leq M.$
To apply Lemma 3.2, we note that the multilinear form $L$ in Lemma 3.1 is clearly symmetric, and Lemma 3.1 precisely asks for an upper bound on the operator norm of $L$. By Lemma 3.2, it suffices to establish Lemma 3.1 when all rows of $A$ are identical, denoted by a vector $x\in{\mathcal{H}}$; also, by scaling we may assume that $\sum_{i=1}^{n}x_{i}^{2}=n$. In this case, the target quantity in Lemma 3.1 becomes $e_{\ell}(x):=\sum_{S\subseteq[n]:|S|=\ell}\prod_{i\in S}x_{i}$, the elementary symmetric polynomial of $x=(x_{1},\cdots,x_{n})$. The central inequality of this section is summarized in the following theorem.
Theorem 3.3.
For $0\leq\ell\leq n$, the following upper bound holds:
  1. 1.
    If $x\in\mathbb{C}^{n}$ with $\sum_{i=1}^{n}x_{i}=0$ and $\sum_{i=1}^{n}|x_{i}|^{2}=n$, then (define $10^{0}:=$ 1)
    $|e_{\ell}(x)|^{2}\leq\frac{n^{n}}{\ell^{\ell}(n-\ell)^{n-\ell}}<3\sqrt{\ell+1}\cdot\binom{n}{\ell}.$ (24)
  2. 2.
    If in addition the condition $x\in\mathbb{R}^{n}$ holds, then an improved upper bound is available:
    $|e_{\ell}(x)|\leq\sqrt{10\binom{n}{\ell}}.$ (25)
It is clear that the inequality (25) in the real case proves Lemma 3.1 after the reduction through Lemma 3.2. The complex inequality (24) will also be useful in Section 4. Again, we remark that the condition $\sum_{i=1}^{n}x_{i}=0$ is crucial for Theorem 3.3, for $x=(1,\ldots,1)$ would lead to $e_{\ell}(x)=\binom{n}{\ell}$. Theorem 3.3 gives a bound on $|e_{\ell}(x)|$ when $e_{1}(x)$ and $e_{2}(x)$ are known. It therefore adds to a recent line of work [GY14, MRT19, DHH20, Tao23] proving bounds on elementary symmetric polynomials given the values of two consecutive lower-order elementary symmetric polynomials. For real $x\in\mathbb{R}^{n}$, this prior work can be used to deduce the bound
$|e_{\ell}(x)|^{2}\leq C\cdot A^{\ell}\binom{n}{\ell},$
for some constant $A>1$. The best known constant is obtained via the differentiation trick in [Tao23], giving
$|e_{\ell}(x)|^{2}\leq\binom{n}{\ell}^{2}\left(\frac{\ell-1}{n-1}\right)^{\ell},$
corresponding to a constant $A=e$ by Stirling’s approximation. This idea could also be applied to the complex case, where using the Schoenberg conjecture/Malamud–Pereira theorem [Mal05, Per03] leads to the same upper bound.
However, the application to our problem requires to have the best possible constant $A=1$ (note that the example in (18) implies that $A<1$ is impossible). For example, the quantity $f(\mu)$ in (2) could be arbitrarily close to $1$, and only $A=1$ makes the geometric sum $\sum_{\ell}A^{\ell}f(\mu)^{\ell}$ converge; we will also see similar scenarios in the application of Lemma 4.5 in Section 4. As shown in Theorem 3.3, this constant turns out to be achievable (possibly at some tolerable price of $\mathrm{poly}(\ell)$) using the saddle point method.
Proof of Theorem 3.3 (First Part).
It is clear that $e_{\ell}(x)$ is the coefficient of $z^{\ell}$ in $\prod_{k=1}^{n}(1+x_{k}z)$. By Cauchy’s formula, for any $r>0$ we have
$|e_{\ell}(x)|=\left|\frac{1}{2\pi\mathrm{i}}\oint_{|z|=r}\frac{\prod_{k=1}^{n}(1+x_{k}z)}{z^{\ell}}\frac{\mathrm{d}z}{z}\right|\leq\max_{|z|=r}\left|\frac{\prod_{k=1}^{n}(1+x_{k}z)}{z^{\ell}}\right|.$
By AM-GM and the assumptions on $x$,
$\prod_{k=1}^{n}|1+x_{k}z|^{2}$ $=\prod_{k=1}^{n}\left(1+|x_{k}|^{2}|z|^{2}+2\Re(x_{k}z)\right)$
$\leq\left(\frac{1}{n}\sum_{k=1}^{n}\left(1+|x_{k}|^{2}|z|^{2}+2\Re(x_{k}z)\right)\right)^{n}=\left(1+|z|^{2}\right)^{n}.$
A combination of the previous steps gives
$|e_{\ell}(x)|^{2}\leq\inf_{r>0}\frac{(1+r^{2})^{n}}{r^{2\ell}}=\frac{n^{n}}{\ell^{\ell}(n-\ell)^{n-\ell}},$
by choosing $r^{2}=\frac{\ell}{n-\ell}$ for $11\leq\ell\leq n-$ 1, $r\to 0$ for $\ell=0$, and $r\to\infty$ for $\ell=n$, establishing the first inequality of (24). The second step of (24) simply follows from Stirling’s approximation. ∎
The proof of the second part (25) of Theorem 3.3 is more involved. First, relying critically on a property of real-rooted polynomials (which fails for the complex case), we argue that the coordinates of the real maximizer $x$ of $|e_{\ell}(x)|$ can only take two values. The evaluation of $|e_{\ell}(x)|$ in the resulting simplified scenario is still highly challenging, and we apply a more careful saddle point analysis to arrive at (25). We defer the details to Appendix B.
4 Upper bound via matrix permanents
The bounds presented in Section 3 are based on the identity
$\frac{\mathrm{d}\mathbb{P}_{n}}{\mathrm{d}\mathbb{Q}_{n}}(X)=\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{n}\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\right]=\sum_{S\subseteq[n]}\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i\in S}\frac{\mathrm{d}\Psi_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\right].$
As mentioned in Section 1.2, this expression may be viewed as an expansion of the matrix permanent in (9) around the all-ones matrix. The benefit of this approach is that each of the terms in this expansion are orthogonal in $L^{2}(\mathbb{Q}_{n})$. However, our proof relies crucially on Lemma 3.2, which is used to crudely upper bound each permanent appearing in the above expansion by the permanent of a matrix with identical rows. This bound therefore ignores any additional structure in the likelihood ratio, and fails to give tight results when ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})>1$.
In this section, we derive refined bounds for ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})>1$ by treating the $\chi^{2}$ divergence between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ directly. This gives proofs of the other upper bounds (5) and (6) in Theorem 1.2. We also show how to use similar ideas to obtain Theorem 1.6. In Section 4.1 we express the $\chi^{2}$ divergence between $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ as a proper matrix permanent, and upper bound it in two ways, described in Section 4.2 and Section 4.3 respectively, both of which rely on the use of complex random variables.
4.1 Divergence as a matrix permanent
We first recall the expression given in Section 1.2 for $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ as a matrix permanent.
Lemma 4.1.
Given $P_{1},\cdots,P_{n}\in{\mathcal{P}}$, define a matrix $A\in\mathbb{R}^{n\times n}$ as
$A_{ij}=\frac{1}{n}\int\frac{\mathrm{d}P_{i}\mathrm{d}P_{j}}{\mathrm{d}\overline{P}},\qquad\text{with }\overline{P}:=\frac{1}{n}\sum_{i=1}^{n}P_{i}.$ (26)
Then under the setting of Theorem 1.2, the following identity holds:
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})=\frac{n^{n}}{n!}\mathrm{Perm}(A)-1.$
The following lemma summarizes some properties of the matrix $A$ in (26).
Lemma 4.2.
The matrix $A$ in (26) satisfies the following properties:
  1. 1.
    it is doubly stochastic and PSD;
  2. 2.
    its trace satisfies $\mathrm{Tr}(A)\leq 1+{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$;
  3. 3.
    its eigenvalues satisfy $1=\lambda_{1}>\lambda_{2}\geq\ldots\geq\lambda_{n}\geq 0$, with leading eigenvector $v_{1}={\bf 1}$ and a spectral gap
    $1-\lambda_{2}\geq\frac{1}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}.$
By Lemma 4.1 and Lemma 4.2, the computation of the $\chi^{2}$ divergence $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ reduces to upper bounding the permanent of $A$. Moreover, in light of the van der Waerden conjecture/Egorychev–Falikman theorem [vdW26, Ego81, Fal81], which states that $\mathrm{Perm}(A)\geq\frac{n!}{n^{n}}$ for all $n\times n$ doubly stochastic matrices $A$, showing that $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})=O(1)$ amounts to showing that $\mathrm{Perm}(A)$ is nearly as small as is possible for a doubly-stochastic matrix. Motivated by the eigenstructure of $A$ established in Lemma 4.2, we write $A=UDU^{\top}$ for an orthogonal matrix $U\in\mathbb{R}^{n\times n}$ and a diagonal matrix $D=\mathrm{diag}(\lambda_{1},\cdots,\lambda_{n})$ consisting of the eigenvalues. By a generalized Cauchy–Binet formula for permanents [MM65, Theorem 2.1], it holds that
$\mathrm{Perm}(A)=\sum_{\begin{subarray}{c}(\ell_{1},\ldots,\ell_{n})\in\mathbb{N}^{n}\\ \sum_{i=1}^{n}\ell_{i}=n\end{subarray}}\frac{\mathrm{Perm}(U_{\ell_{1},\ldots,\ell_{n}})^{2}}{\ell_{1}!\cdots\ell_{n}!}\lambda_{1}^{\ell_{1}}\cdots\lambda_{n}^{\ell_{n}},$
where $U_{\ell_{1},\ldots,\ell_{n}}$ denotes the $n\times n$ matrix with the first column of $U$ appearing $\ell_{1}$ times, second column appearing $\ell_{2}$ times, and so on. By taking $\lambda_{1}=1$ into account, we obtain the equality
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})+1=\frac{n^{n}}{n!}\sum_{\ell=0}^{n}\sum_{\begin{subarray}{c}(\ell_{2},\ldots,\ell_{n})\in\mathbb{N}^{n-1}\\ \sum_{i=2}^{n}\ell_{i}=\ell\end{subarray}}\frac{\mathrm{Perm}(U_{n-\ell,\ell_{2},\ldots,\ell_{n}})^{2}}{(n-\ell)!\ell_{2}!\cdots\ell_{n}!}\lambda_{2}^{\ell_{2}}\cdots\lambda_{n}^{\ell_{n}}=:\sum_{\ell=0}^{n}S_{\ell},$ (27)
which is a sum of homogeneous polynomials $S_{\ell}$ in $(\lambda_{2},\ldots,\lambda_{n})$ of total degree $\ell=0,1,\ldots,n$.
To relate the permanent view (27) with the doubly centered expansion in Section 3, we refer to Appendix A for some identities between $S_{\ell}$ and related quantities. In particular, we note that the permanent $\mathrm{Perm}(U_{n-\ell,\ell,0,\ldots,0})$ is exactly a multiple of $e_{\ell}(u_{2})$, the elementary symmetric polynomial of second column vector $u_{2}$ of $U$. Therefore, Theorem 3.3 could be used to upper bound this quantity. In addition, Lemma 3.2 reduces the general quantity $\mathrm{Perm}(U_{n-\ell,\ell_{2},\ldots,\ell_{n}})$ to the case $\mathrm{Perm}(U_{n-\ell,\ell,0,\ldots,0})$ with identical columns, so the arguments in Section 3 leads to an upper bound of (27) in terms of $\sum_{i=2}^{n}\lambda_{i}\leq{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ by Lemma 4.2. However, this approach does not capture additional structures of $(\lambda_{2},\dots,\lambda_{n})$ such as the spectral gap. To this end, it might be a natural idea to establish pointwise upper bounds on $\mathrm{Perm}(U_{n-\ell,\ell_{2},\ldots,\ell_{n}})$. We have not succeeded in this approach and leave it as an open direction; instead, we choose to upper bound the individual sum $S_{\ell}$ or the entire sum $\sum_{\ell=0}^{n}S_{\ell}$. These form the topics of the subsequent sections.
4.2 Proof of Theorem 1.2 via bounding the entire sum
In this section, we upper bound the entire sum $\sum_{\ell=0}^{n}S_{\ell}$ in (27) and use it to prove Theorem 1.2. The central result of this section is the following upper bound:
Lemma 4.3.
Using the notation of (27), for any $\lambda_{2},\dots,\lambda_{n}\geq 0$ it holds that
$\sum_{\ell=0}^{n}S_{\ell}\leq\sum_{\ell=0}^{n}\sum_{\begin{subarray}{c}(\ell_{2},\ldots,\ell_{n})\in\mathbb{N}^{n-1}\\ \sum_{i=2}^{n}\ell_{i}=\ell\end{subarray}}\lambda_{2}^{\ell_{2}}\cdots\lambda_{n}^{\ell_{n}}.$
Comparing this expression with (27), Lemma 4.3 shows that replacing all coefficients in (27) by $1$ gives an upper bound. The proof of Lemma 4.3 relies on the use of complex normal random variables. Recall that $z\sim{\mathcal{C}}{\mathcal{N}}(0,1)$ denotes $z=x+\mathrm{i}y$ with independent $x,y\sim{\mathcal{N}}(0,\frac{1}{2})$, and $z\sim{\mathcal{C}}{\mathcal{N}}(0,\Sigma)$ denotes a vector $z=x+\mathrm{i}y$ with independent $x,y\sim{\mathcal{N}}(0,\frac{\Sigma}{2})$ for real PSD matrices $\Sigma$. It is known that the moments of $z\sim{\mathcal{C}}{\mathcal{N}}(0,1)$ are
$\mathbb{E}[z^{m}\bar{z}^{n}]=\begin{cases}n!&\text{if }m=n,\\ 0&\text{if }m\neq n.\end{cases}$ (28)
The following identity between matrix permanents and complex normal random vectors is observed in [AGGS17, Lemma 2] and earlier in [Gur03].
Lemma 4.4.
For $P\in\mathbb{R}^{m\times n}$ and $z\sim{\mathcal{C}}{\mathcal{N}}(0,I_{n})$, the following identity holds:
$\mathrm{Perm}(PP^{\top})=\mathbb{E}\left[\prod_{i=1}^{m}|(Pz)_{i}|^{2}\right].$
Lemma 4.4 follows from the classical Isserlis’ theorem [Iss18] (or Wick’s formula) for products of joint normal random variables, and we include a proof in Section C.7 for completeness. Now we present the proof of Lemma 4.3.
Proof of Lemma 4.3.
Let $A=UDU^{\top}$ be the matrix in Lemma 4.1, and $P=UD^{1/2}$. Then
$\frac{n!}{n^{n}}\sum_{\ell=0}^{n}S_{\ell}=\mathrm{Perm}(A)=\mathrm{Perm}(PP^{\top})=\mathbb{E}\left[\prod_{i=1}^{n}|(Pz)_{i}|^{2}\right]$
by Lemma 4.4. Since
$\sum_{i=1}^{n}|(Pz)_{i}|^{2}=\|Pz\|_{2}^{2}=z^{\mathsf{H}}P^{\top}Pz=z^{\mathsf{H}}Dz=\sum_{i=1}^{n}\lambda_{i}|z_{i}|^{2},$
the AM-GM inequality then gives that
$\mathbb{E}\left[\prod_{i=1}^{n}|(Pz)_{i}|^{2}\right]$ $\leq\mathbb{E}\left[\left(\frac{1}{n}\sum_{i=1}^{n}|(Pz)_{i}|^{2}\right)^{n}\right]=\mathbb{E}\left[\left(\frac{1}{n}\sum_{i=1}^{n}\lambda_{i}|z_{i}|^{2}\right)^{n}\right]$
=$\overset{\rm(a)}{=}\frac{1}{n^{n}}\sum_{\begin{subarray}{c}(\ell_{1},\ldots,\ell_{n})\in\mathbb{N}^{n}\\ \sum_{i=1}^{n}\ell_{i}=n\end{subarray}}\binom{n}{\ell_{1},\ldots,\ell_{n}}\lambda_{1}^{\ell_{1}}\cdots\lambda_{n}^{\ell_{n}}\mathbb{E}\left[|z_{1}|^{2\ell_{1}}\cdots|z_{n}|^{2\ell_{n}}\right]$
=$\overset{\rm(b)}{=}\frac{n!}{n^{n}}\sum_{\begin{subarray}{c}(\ell_{1},\ldots,\ell_{n})\in\mathbb{N}^{n}\\ \sum_{i=1}^{n}\ell_{i}=n\end{subarray}}\lambda_{1}^{\ell_{1}}\cdots\lambda_{n}^{\ell_{n}},$
where (a) is the multinomial theorem, and (b) follows from (28). The proof is complete. ∎
Next we prove the second upper bound (5) of Theorem 1.2 from Lemma 4.3. Based on (27) and Lemma 4.3, we have
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})+1\leq\sum_{\ell=0}^{n}\sum_{\begin{subarray}{c}(\ell_{2},\ldots,\ell_{n})\in\mathbb{N}^{n-1}\\ \sum_{i=2}^{n}\ell_{i}=\ell\end{subarray}}\lambda_{2}^{\ell_{2}}\cdots\lambda_{n}^{\ell_{n}}\leq\sum_{(\ell_{2},\ldots,\ell_{n})\in\mathbb{N}^{n-1}}\lambda_{2}^{\ell_{2}}\cdots\lambda_{n}^{\ell_{n}}=\prod_{i=2}^{n}\frac{1}{1-\lambda_{i}}.$
Thanks to the spectral gap bound in Lemma 4.2, we have $\lambda_{i}\in\left[0,1-\frac{1}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}\right]$, so that
$\log\frac{1}{1-\lambda_{i}}\leq\lambda_{i}\cdot\frac{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})-1}\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$
by the convexity of $\lambda\mapsto\log\frac{1}{1-\lambda}$ on $[0,1)$. Summing over $i$ gives the upper bound
$\log\left(\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})+1\right)$ $\leq\sum_{i=2}^{n}\lambda_{i}\cdot\frac{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})-1}\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$
$\overset{\rm(c)}{\leq}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\cdot\frac{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})-1}\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\overset{\rm(d)}{\leq}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\left(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\right),$
where (c) follows from Lemma 4.2, (d) uses $\log x\leq x-1$ for all $x>0$. For the inequality (6), we have
$\log\left(\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})+1\right)$ $\leq{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\cdot\frac{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})-1}\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$
$\overset{\rm(e)}{\leq}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\cdot\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1}{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}\log({\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1)\overset{\rm(f)}{\leq}({\mathsf{C}}_{\chi^{2}}({\mathcal{P}})+1)\log({\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1),$
where (e) is due to $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\leq 1+{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})$ and
$\frac{\mathrm{d}}{\mathrm{d}x}\left(\frac{x\log x}{x-1}\right)=\frac{x-1-\log x}{(x-1)^{2}}\geq 0,$
and (f) uses ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})$. Therefore the upper bounds (5) and (6) are established.
4.3 Proof of Theorem 1.6 via bounding the individual sum
In this section, we upper bound the individual sum $S_{\ell}$ in (27) and use it to prove Theorem 1.6. The central result of this section is the following upper bound:
Lemma 4.5.
Using the notation of (27), for any $\lambda_{2},\cdots,\lambda_{n}\geq 0$ and $\ell=0,\ldots,n$ it holds that
$S_{\ell}\leq 3\sqrt{\ell+1}\cdot\sum_{\begin{subarray}{c}(\ell_{2},\ldots,\ell_{n})\in\mathbb{N}^{n-1}\\ \sum_{i=2}^{n}\ell_{i}=\ell\end{subarray}}\lambda_{2}^{\ell_{2}}\cdots\lambda_{n}^{\ell_{n}}.$
Compared with Lemma 4.3, the result of Lemma 4.5 has an additional $O(\sqrt{\ell})$ factor, due to the use of Theorem 3.3 in the complex case. However, such an individual bound is crucial to deal with weighted sums $\sum_{\ell=0}^{n}w_{\ell}S_{\ell}$ in applications such as Theorem 1.6. Again, the proof of Lemma 4.5 relies on the use of complex normal random variables. Recalling the notation $A=UDU^{\top}$ from Section 4.1, let $\widetilde{U}\in\mathbb{R}^{n\times(n-1)}$ be the matrix $U$ with its first column (which is $\frac{\bf 1}{\sqrt{n}}$ by Lemma 4.2) removed, and $\widetilde{D}=\mathrm{diag}(\lambda_{2},\ldots,\lambda_{n})$ with the first eigenvalue $\lambda_{1}=1$ removed. The following lemma establishes an alternative expression of $S_{\ell}$.
Lemma 4.6.
Let $\widetilde{P}=\widetilde{U}\widetilde{D}^{1/2}\in\mathbb{R}^{n\times(n-1)}$ and $z\sim{\mathcal{C}}{\mathcal{N}}(0,I_{n-1})$. Then for $\ell=0,\cdots,n$,
$S_{\ell}=n^{\ell}\frac{(n-\ell)!}{n!}\cdot\mathbb{E}\left[|e_{\ell}(\widetilde{P}z)|^{2}\right],$
where as usual $e_{\ell}$ denotes the elementary symmetric polynomial of order $\ell$.
The proof of Lemma 4.6 again follows from the Isserlis’ theorem and several permutation identities in Lemma A.1, and is deferred to Section C.8. Now we present the proof of Lemma 4.5.
Proof of Lemma 4.5.
Since the matrix $U$ is orthogonal with the first column proportional to ${\bf 1}$, all columns of $\widetilde{U}$ are orthogonal to ${\bf 1}$. Consequently, the vector $\widetilde{P}z=\widetilde{U}\widetilde{D}^{1/2}z$ sums to zero, so (24) in Theorem 3.3 gives
$|e_{\ell}(\widetilde{P}z)|^{2}\leq 3\sqrt{\ell+1}\binom{n}{\ell}\left(\frac{1}{n}\sum_{i=1}^{n}\left|(\widetilde{P}z)_{i}\right|^{2}\right)^{\ell}.$
Since
$\sum_{i=1}^{n}\left|(\widetilde{P}z)_{i}\right|^{2}=z^{\mathsf{H}}\widetilde{P}^{\top}\widetilde{P}z=z^{\mathsf{H}}\widetilde{D}z=\sum_{i=2}^{n}\lambda_{i}|z_{i-1}|^{2},$
by Lemma 4.6 and (28) we get
$S_{\ell}$ $\leq n^{\ell}\frac{(n-\ell)!}{n!}\cdot 3\sqrt{\ell+1}\binom{n}{\ell}\mathbb{E}\left[\left(\frac{1}{n}\sum_{i=2}^{n}\lambda_{i}|z_{i-1}|^{2}\right)^{\ell}\right]$
$=\frac{3\sqrt{\ell+1}}{\ell!}\sum_{\begin{subarray}{c}(\ell_{2},\ldots,\ell_{n})\in\mathbb{N}^{n-1}\\ \sum_{i=2}^{n}\ell_{i}=\ell\end{subarray}}\binom{\ell}{\ell_{2},\ldots,\ell_{n}}\lambda_{2}^{\ell_{2}}\cdots\lambda_{n}^{\ell_{n}}\mathbb{E}\left[|z_{1}|^{2\ell_{2}}\cdots|z_{n-1}|^{2\ell_{n}}\right]$
$=3\sqrt{\ell+1}\sum_{\begin{subarray}{c}(\ell_{2},\ldots,\ell_{n})\in\mathbb{N}^{n-1}\\ \sum_{i=2}^{n}\ell_{i}=\ell\end{subarray}}\lambda_{2}^{\ell_{2}}\cdots\lambda_{n}^{\ell_{n}},$
which is the claimed upper bound. ∎
Next we use the individual upper bounds in Lemma 4.5 to prove Theorem 1.6. Similar to (26) and (27), let the matrix $A\in\mathbb{R}^{(n-1)\times(n-1)}$ be constructed based on $P_{2},\ldots,P_{n}\in{\mathcal{P}}$, and $S_{\ell}$ be the $-\ell$ homegeneous polynomial in its eigenvalues $(\lambda_{2},\ldots,\lambda_{n-1})$. Applying a similar likelihood ratio computation to Lemma 4.1 leads to the following upper bound.
Lemma 4.7.
Under the above notations, the following inequality holds:
${\sf TV}(\mathbb{P}_{n},\mathbb{Q}_{n})^{2}\leq\frac{1}{4}\int\frac{(\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{Q}_{n})^{2}}{\mathrm{d}\overline{P}^{\otimes n}}\leq\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}\sum_{\ell=1}^{n}\ell S_{\ell-1}.$
The proof of Lemma 4.7 is postponed to the appendix, and we proceed to prove Theorem 1.6. By Lemma 4.5, we have
$\sum_{\ell=1}^{n}\ell S_{\ell-1}$ $\leq 3\sum_{\ell=1}^{n}\ell^{3/2}\sum_{\begin{subarray}{c}(\ell_{2},\dots,\ell_{n-1})\in\mathbb{N}^{n-2}\\ \sum_{i=2}^{n-1}\ell_{i}=\ell-1\end{subarray}}\lambda_{2}^{\ell_{2}}\cdots\lambda_{n-1}^{\ell_{n-1}}$
$\overset{\rm(a)}{\leq}3\sum_{\ell=1}^{n}\sum_{\begin{subarray}{c}(\ell_{2},\dots,\ell_{n-1})\in\mathbb{N}^{n-2}\\ \sum_{i=2}^{n-1}\ell_{i}=\ell-1\end{subarray}}(\ell_{2}+1)^{3/2}\lambda_{2}^{\ell_{2}}\cdots(\ell_{n-1}+1)^{3/2}\lambda_{n-1}^{\ell_{n-1}}$
$\leq 3\prod_{i=2}^{n-1}\left(\sum_{\ell=0}^{\infty}(\ell+1)^{3/2}\lambda_{i}^{\ell}\right)\overset{\rm(b)}{\leq}3\prod_{i=2}^{n-1}\left(\frac{1}{2}\sum_{\ell=0}^{\infty}(\ell+2)(\ell+1)\lambda_{i}^{\ell}\right)\overset{\rm(c)}{=}3\prod_{i=2}^{n-1}\frac{1}{(1-\lambda_{i})^{3}},$
where (a) is due to $\prod_{i=2}^{n-1}(1+\ell_{i})\geq 1+\sum_{i=2}^{n-1}\ell_{i}=\ell$, (b) uses $22\sqrt{\ell+1}\leq\ell+$ 2 for all $\ell\geq 0$, and (c) follows from the identity
$\sum_{\ell=0}^{\infty}(\ell+2)(\ell+1)x^{\ell}=\frac{\mathrm{d}^{2}}{\mathrm{d}x^{2}}\sum_{\ell=0}^{\infty}x^{\ell+2}=\frac{2}{(1-x)^{3}},\qquad\text{for }|x|<1.$
By the same argument used to deal with the final product in the proof of Lemma 4.3, Lemma 4.7 leads to
${\sf TV}(\mathbb{P}_{n},\mathbb{Q}_{n})^{2}\leq\frac{3{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})},$
which establishes the first claim of Theorem 1.6. The second claim of Theorem 1.6 simply follows from the Cauchy–Schwarz inequality.
5 Discussion
5.1 Tightness of upper bounds
In this section we discuss the tightness of the upper bounds in Theorem 1.2. Recall that the upper bound (4) is quadratic in ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ when ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ is small, and the upper bound (5) or (6) is a power function of $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ when ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ is large. The next lemma shows that this is essentially the right behavior of $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ for every family ${\mathcal{P}}$.
Lemma 5.1.
In the setting of Theorem 1.2, the following lower bound holds:
$\sup_{n}\sup_{P_{1},\ldots,P_{n}\in{\mathcal{P}}}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\geq\sup_{n}\sup_{P_{1},\ldots,P_{n}\in{\mathcal{P}}}\frac{1}{\sqrt{1-\lambda_{2}(A)^{2}}}-1,$
where $A=A(P_{1},\dots,P_{n})$ is defined in (26), and $\lambda_{2}(A)$ denotes the second largest eigenvalue of $A$. In particular, it holds that
$\sup_{n}\sup_{P_{1},\ldots,P_{n}\in{\mathcal{P}}}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\geq\max\left\{\frac{{\mathsf{D}}_{H^{2}}({\mathcal{P}})^{2}}{8},\frac{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})^{1/4}}{\sqrt{2}}-1\right\}.$
Combined with Theorem 1.2, Lemma 5.1 shows that when the capacity of the family ${\mathcal{P}}$ is small, we have ${\mathsf{D}}_{H^{2}}({\mathcal{P}})^{2}\lesssim\sup_{n,P_{1},\dots,P_{n}}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\lesssim{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})^{2}$, which is often tight as ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})\approx{\mathsf{D}}_{H^{2}}({\mathcal{P}})$ in view of the fact that most $f$-divergences are locally $\chi^{2}$-like [PW24, Chapter 7.10]. At the other extreme, when the family ${\mathcal{P}}$ is rich, we have $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})^{\Omega(1)}-1\leq\sup_{n,P_{1},\dots,P_{n}}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}-1$. Therefore, up to a gap in the exponent, the maximum $H^{2}$ singularity $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ plays a central role when it is large. The proof of Lemma 5.1 computes the variance of a linear test function $\frac{1}{n}\sum_{i=1}^{n}f(X_{i})$ under $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$.
The next natural question is on the tightness of the exponent ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ in the upper bound (5). Using the permanent representation of the $\chi^{2}$ divergence in Lemma 4.1, we can equivalently ask whether the following bound in Section 4 is tight:
$\mathrm{Perm}(A)\leq\frac{n!}{n^{n}}\left(\frac{1}{\Delta}\right)^{1+{\mathsf{C}}},$ (29)
where $A\in\mathbb{R}^{n\times n}$ is PSD and doubly stochastic, with $\mathrm{Tr}(A)\leq 1+{\mathsf{C}}$ and $1-\lambda_{2}(A)\geq\Delta$. Our next lemma shows that the permanent upper bound (29) is essentially tight if we only make use of the trace and spectral gap of $A$.
Lemma 5.2.
There exist absolute constants $r,r^{\prime},{\mathsf{C}}_{0}>0$ and $\Delta_{0}\in(0,1)$ such that for any ${\mathsf{C}}\geq{\mathsf{C}}_{0}$ and $\Delta\leq\Delta_{0}$, one may find some $n\in\mathbb{N}$ and $A\in\mathbb{R}^{n\times n}$ satisfying: 1) $A$ is PSD and doubly stochastic; 2) $\mathrm{Tr}(A)\leq 1+{\mathsf{C}}$; 3) $1-\lambda_{2}(A)\geq\Delta$; and 4)
$\mathrm{Perm}(A)\geq\frac{n!}{n^{n}}\left(\frac{1}{\Delta}\right)^{r{\mathsf{C}}}.$
Similarly, one may also find a family ${\mathcal{P}}$ of distributions with ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{C}}$, $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\leq\Delta^{-1}$, and
$\sup_{n}\sup_{P_{1},\dots,P_{n}\in{\mathcal{P}}}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\geq\left(\frac{1}{\Delta}\right)^{r^{\prime}{\mathsf{C}}}-1.$
By Lemma 5.2, the exponent ${\mathsf{C}}$ in both the permanent upper bound (29) and the $\chi^{2}$ divergence upper bound (5) in Theorem 1.2 is tight in the worst case, up to multiplicative constants. However, as opposed to Lemma 5.1, the statement of Lemma 5.2 is not pointwise in ${\mathcal{P}}$; in other words, this does not mean that the upper bounds of Theorem 1.2 are always tight for a specific class ${\mathcal{P}}$. For example, for the Gaussian family (cf. point 1 of Corollary 1.3) with $\mu>1$, Theorem 1.2 and Lemma 5.1 only imply that $\exp(\Omega(\mu^{2}))\leq\sup_{n,P_{1},\dots,P_{n}}\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq\exp(O(\mu^{3}))$, still exhibiting a gap on the exponent. An intuitive reason is that, for specific ${\mathcal{P}}$, further properties of the matrix $A$ (besides the trace and spectral gap) could potentially be exploited to lead to better results.
Although we leave the exploitation of further properties of $A$ as an open question, we mention one possible strategy here. Note that the upper bound in Section 4.2 claims that
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})+1\leq\prod_{i=2}^{n}\frac{1}{1-\lambda_{i}}=\left(n\sum_{T}\prod_{(i,j)\in T}A_{ij}\right)^{-1},$
where we sum over the edge set $T$ of all $n^{n-2}$ spanning trees of the complete graph over $[n]$, and the equality is due to the weighted matrix tree theorem [Kir47] applied to the graph Laplacian $I-A$. We conjecture that lower bounding this spanning tree polynomial of $A$ could lead to an improved upper bound for $\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})$ in the Gaussian family.
5.2 Application to a Gaussian sequence model
The contiguity relation $\{\mathbb{P}_{n}\}\triangleleft\{\mathbb{Q}_{n}\}$ in Theorem 1.2, or more specifically the mutual information relation (1) in Corollary 1.4, can be applied in statistical problems. We review a classical example of a Gaussian sequence model as follows. Let the observation vector be $X\sim{\mathcal{N}}(\theta,I_{n})$, with an unknown mean vector $\theta\in\mathbb{R}^{n}$ lying in the following $\ell^{p}$ ball $\Theta_{p}(R)$:
$\Theta_{p}(R):=\left\{\theta\in\mathbb{R}^{n}:\frac{1}{n}\sum_{i=1}^{n}|\theta_{i}|^{p}\leq R^{p}\right\}.$
Here $p>0$ is a given norm parameter, and $R=R_{n}$ is a given radius; we will assume that $R=o(1)$ to promote sparsity in this problem. Our target is to understand the minimax $\ell^{q}$ risk $r_{p,q}^{\star}(n,R)$ for the mean estimation:
$r_{p,q}^{\star}(n,R):=\inf_{\widehat{\theta}}\sup_{\theta\in\Theta_{p}(R)}\mathbb{E}_{\theta}\left[\frac{1}{n}\sum_{i=1}^{n}|\widehat{\theta}_{i}-\theta_{i}|^{q}\right].$
It is a classical result [DJ94] that for $q\geq p\vee 1$, the minimax risk is $r_{p,q}^{\star}(n,R)=(1+o(1))\mu^{q-p}R^{p}$ if $R=o(1)$ and $\frac{nR^{p}}{\mu^{p}}=\omega(1)$, where $\mu=\sqrt{2\log(1/R^{p})}$ is the threshold used in the soft-thresholding estimator. The quantity $s:=\frac{nR^{p}}{\mu^{p}}$ denotes the average sparsity level of $\theta$, and the extra condition $s=\omega(1)$ is used in the arguments of the minimax lower bound: when lower bounding the Bayes risk under the i.i.d. two-point prior [^3] In [DJ94], a three-point prior $(1-\alpha)\delta_{0}+\frac{\alpha}{2}(\delta_{-\mu}+\delta_{\mu})$ was used. However, their argument gives the same lower bound under the two-point prior. $(1-\alpha)\delta_{0}+\alpha\delta_{\mu}$ with $\alpha=(1+o(1))\frac{s}{n}$, the condition $s=\omega(1)$ ensures that the i.i.d. prior concentrates on $\Theta_{p}(R)$ with high probability. This condition was later removed in [Zha12], by lower bounding the Bayes risk under a permutation prior $\theta=(v_{\pi(1)},\dots,v_{\pi(n)})$ with $\pi\sim\text{\rm Unif}(S_{n})$ when $s=O(1)$, with a vector $v=(\mu,\dots,\mu,0,\dots,0)$ up to a rounding error under possibly non-integer $s$. The major benefit of the permutation prior is that $\theta\in\Theta_{p}(R)$ almost surely. However, based on the technical tools in [Zha12, Prop. 2.3], the current lower bound for permutation priors only works for a low sparsity level $s=O(\sqrt{n})$.
It is then a natural question to ask if the minimax risk $r_{p,q}^{\star}(n,R)=(1+o(1))\mu^{q-p}R^{p}$ can also be established using the above permutation mixture for all $s=\omega(1)$. It turns out that a combination of Corollary 1.4 and the Gaussian I-MMSE formula [GSV05] directly provides an affirmative answer for the squared loss $q=2$.
Lemma 5.3.
Let $q=2\geq p$ and $R=o(1)$. For $\mu=\sqrt{2\log(1/R^{p})}$ and $s=\frac{nR^{p}}{\mu^{p}}=\omega(1)$, the Bayes $\ell^{2}$ risk under the permutation prior $\theta=(v_{\pi(1)},\dots,v_{\pi(n)})$ with $\pi\sim\text{\rm Unif}(S_{n})$ and
$v=(\sqrt{1-\varepsilon}\mu,\dots,\sqrt{1-\varepsilon}\mu,0,\dots,0),\quad\|v\|_{0}=\lfloor s\rfloor$
is lower bounded by $(1-o(1))(1-c(\varepsilon))\mu^{2-p}R^{p}$, with $c(\varepsilon)\to 0$ as $\varepsilon\to 0^{+}$.
Choosing $\varepsilon\to 0^{+}$, Lemma 5.3 recovers the lower bound of $r_{p,2}^{\star}(n,R)$ via a permutation prior.
5.3 Empirical Bayes with quadratic losses
In this section we present an alternative strategy developed in [GR09] for empirical Bayes models under the quadratic loss. We begin with a technical result. Let $P_{0},\dots,P_{n}\in{\mathcal{P}}$ be $n+1$ probability distributions, and for $i\in\{0,\dots,n\}$, define the permutation mixture $\mathbb{P}_{-i}$ as
$\mathbb{P}_{-i}:=\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{\{0,\dots,n\}\backslash\{i\}})}\left[\otimes_{j\neq i}P_{\pi(j)}\right],$
where $S_{A}$ denotes the permutation group over the finite set $A$. In other words, $\mathbb{P}_{-i}$ is the permutation mixture of $\{P_{0},\dots,P_{n}\}\backslash\{P_{i}\}$. The idea of [GR09] leads to the following inequality (though originally stated in a more complicated way):
Lemma 5.4.
For ¯$\overline{\mathbb{P}}=\frac{1}{n+1}\sum_{i=0}^{n}\mathbb{P}_{-i}$ and every $i\in\{0,\dots,n\}$ it holds that
$\chi^{2}(\overline{\mathbb{P}}\|\mathbb{P}_{-i})\leq\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n+1}.$
In particular, for every $i,j\in\{0,\dots,n\}$ we have
$H^{2}(\mathbb{P}_{-i},\mathbb{P}_{-j})\leq\frac{4{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n+1}.$
Unlike Theorem 1.6, the proof of Lemma 5.4 relies on a clever coupling between $\overline{\mathbb{P}}$ and $\mathbb{P}_{-i}$ to reduce $\chi^{2}(\overline{\mathbb{P}}\|\mathbb{P}_{-i})$ to the $\chi^{2}$ divergence between a mixture and a simple distribution (the strategy in Section 2.2). Since ${\sf TV}(P,Q)^{2}\leq H^{2}(P,Q)$, Lemma 5.4 indeed provides a better dependence on ${\mathcal{P}}$ in the $O_{{\mathcal{P}}}(\frac{1}{n})$ upper bound of ${\sf TV}(\mathbb{P}_{n},\mathbb{Q}_{n})^{2}$ in Theorem 1.6. However, the results of Theorem 1.6 and Lemma 5.4 are incomparable in general, because of the product measure $\overline{P}^{\otimes n}$ used in Theorem 1.6 which could be easier to handle in the computation of $\mathbb{E}_{\overline{P}^{\otimes n}}[f^{2}]$.
The upper bound on the Hellinger distance in Lemma 5.4 helps to study empirical Bayes models under the quadratic loss. In the compound setting of empirical Bayes above Corollary 1.7, the risk difference $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)$ could enjoy a smaller upper bound under a quadratic loss $L(\vartheta,\widehat{\vartheta})=\|\vartheta-\widehat{\vartheta}\|_{2}^{2}$. In this case, it is shown in [GR09] that the optimal simple and permutation invariant decision rules are
$\widehat{\vartheta}^{\mathrm{S}}=\left(\mathbb{E}\left[\vartheta_{1}|X_{1}\right],\dots,\mathbb{E}\left[\vartheta_{n}|X_{n}\right]\right),\quad\text{and}\quad\widehat{\vartheta}^{\mathrm{PI}}=\left(\mathbb{E}\left[\vartheta_{1}|X\right],\dots,\mathbb{E}\left[\vartheta_{n}|X\right]\right),$
respectively. In addition, the risk difference admits a simple form
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=n\mathbb{E}\left[\left(\mathbb{E}\left[\vartheta_{1}|X_{1}\right]-\mathbb{E}\left[\vartheta_{1}|X\right]\right)^{2}\right].$
For a class of models ${\mathcal{P}}$ (such as the Gaussian location model), it is shown in [GR09] that $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=O_{{\mathcal{P}}}(1)$, improving over the $O_{{\mathcal{P}}}(\sqrt{n})$ upper bound in Corollary 1.7 for general losses. The following lemma shows that this result could be extended to more general classes ${\mathcal{P}}$ using Lemma 5.4.
Lemma 5.5.
Let $P_{\theta_{1}},\dots,P_{\theta_{n}}\in{\mathcal{P}}$ with ${\mathsf{D}}_{\chi^{2}}({\mathcal{P}})<\infty$, and $|\theta_{i}|\leq M$ for all $i\in[n]$. Then
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=n\mathbb{E}\left[\left(\mathbb{E}\left[\vartheta_{1}|X_{1}\right]-\mathbb{E}\left[\vartheta_{1}|X\right]\right)^{2}\right]\leq 4M^{2}{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(1+{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})).$
5.4 General basis expansions
As mentioned in Section 3, in the Gaussian case where $P_{i}={\mathcal{N}}(\theta_{i},1)$, the methods of moments and cumulants described in Section 2 rely on the expansion $\varphi(x-\theta)/\varphi(x)=1+\sum_{k\geq 1}h_{k}(x)\frac{\theta^{k}}{k!}$, where $\varphi$ is the density of ${\mathcal{N}}(0,1)$ and $h_{k}$ is the $k$th Hermite polynomial. By contrast, the doubly centered expansion is based on the representation $\varphi(x-\theta)/\varphi_{0}(x)=1+\Psi(x,\theta)$, where $\phi_{0}(x)=\frac{1}{n}\sum_{i=1}^{n}\varphi(x-\theta_{i})$ denotes the marginal density of each coordinate under $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$. This representation possesses the important property that the functions $(x_{1},\dots,x_{n})\mapsto\prod_{i\in S}\Psi(x_{i},\theta_{i})$ are orthogonal with respect to $L^{2}(\mathbb{Q}_{n})$ and $\frac{1}{n}\sum_{i=1}^{n}\Psi(x,\theta_{i})=0$.
In principle, this expansion could be developed further as in the Hermite case by writing
$\frac{\varphi(x-\theta)}{\varphi_{0}(x)}=1+\sum_{k\geq 1}\psi_{k}(x)g_{k}(\theta)$ (30)
for some functions $\psi_{k}$ which are orthogonal in $L^{2}(\varphi_{0})$ and some suitable coefficient functions $g_{k}$. Such an expansion would still possess the doubly centered property described above: indeed, averaging both sides of (30) over the mixing measure shows that in such an expansion, $\frac{1}{n}\sum_{i=1}^{n}g_{k}(\theta_{i})=0$ always holds. However, this development seems to offer no benefits in the setting of this work. The intuitive explanation of this fact is that $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ differ starting from the second moment, so the specific choices of the higher-order basis functions become unimportant (where every choice works).
Nevertheless, we conjecture that (30) may be useful in more general scenarios. For example, if $\mathbb{P}_{n}=\nu_{\mathbb{P}}\star{\mathcal{N}}(0,I_{n})$ for a model in which $\mathbb{E}_{\vartheta\sim\mathbb{P}}\left[\prod_{i=1}^{n}g_{\alpha_{i}}(\vartheta_{i})\right]=0$ for all multi-indices $\alpha$ satisfying $|\alpha|\leq K$, then we anticipate that a full basis expansion as in (30) can be used to take advantage of additional cancellations if it is possible to give good bounds on $g_{k}(\theta)$ for large $k$.
Acknowledgments.
We are grateful to Terence Tao for suggesting the use of the saddle point method in the proof of Theorem 3.3, to Cun-Hui Zhang for introducing us the question of the Bayes risk under a permutation prior in Lemma 5.3 and the reference [Zha12], to Yury Polyanskiy for pointing us to the reference [TP23], and to Yunzi Ding and Cheng Mao for discussions about an earlier version of this project. JNW is partially supported by a Sloan Research Fellowship and NSF grant DMS-2210583.
Appendix A Useful identities
We discuss several identities involving permutations and matrix permanents which will be useful in the proofs of several results. Fix an integer $\ell\in[n]$ and $n$ distributions $P_{1},\cdots,P_{n}\in{\mathcal{P}}$. The first quantity appears in (23) of the doubly centered expansion in Section 3, and is formally defined as
$R_{\ell}=\mathbb{E}_{X_{1},\ldots,X_{\ell}\sim\overline{P}}\left[\left(\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{\ell}\frac{\mathrm{d}(P_{\pi(i)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{i})\right]\right)^{2}\right],$ (31)
with $\overline{P}=\frac{1}{n}\sum_{i=1}^{n}P_{i}$ as usual. The second quantity is the $-\ell$ homogeneous polynomial $S_{\ell}$ in (27) of the matrix permanent approach in Section 4: for the matrix $A\in\mathbb{R}^{n\times n}$ defined in (26) and its eigen-decomposition $A=UDU^{\top}$ with $D=\mathrm{diag}(\lambda_{1},\cdots,\lambda_{n})$, the quantity $S_{\ell}$ is defined as
$S_{\ell}=\frac{n^{n}}{n!}\sum_{\begin{subarray}{c}(\ell_{2},\ldots,\ell_{n})\in\mathbb{N}^{n-1}\\ \sum_{i=2}^{n}\ell_{i}=\ell\end{subarray}}\frac{\mathrm{Perm}(U_{n-\ell,\ell_{2},\ldots,\ell_{n}})^{2}}{(n-\ell)!\ell_{2}!\cdots\ell_{n}!}\lambda_{2}^{\ell_{2}}\cdots\lambda_{n}^{\ell_{n}}.$ (32)
Here we recall that $U_{\ell_{1},\ldots,\ell_{n}}$ denotes the $n\times n$ matrix with the first column of $U$ appearing $\ell_{1}$ times, second column appearing $\ell_{2}$ times, and so on. The last quantity $T_{\ell}$ is a useful intermediate quantity in our proof: let $\overline{A}=A-\frac{1}{n}{\bf 1}{\bf 1}^{\top}$ be the centered version of $A$, and define
$T_{\ell}=\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=\ell\end{subarray}}\mathrm{Perm}(\overline{A}_{S,S^{\prime}}),$ (33)
where $\overline{A}_{S,S^{\prime}}$ denotes the submatrix of $\overline{A}$ with row indices in $S$ and column indices in $S^{\prime}$.
The following lemma shows that, within constant multiples of each other, the above quantities are essentially the same.
Lemma A.1.
For the quantities defined in (31)–(33), the following identity holds:
$S_{\ell}=\binom{n}{\ell}\cdot R_{\ell}=\frac{(n-\ell)!}{n!}n^{\ell}\cdot T_{\ell}.$
Proof.
We first prove the relation between $R_{\ell}$ and $T_{\ell}$. By introducing an independent copy $\pi^{\prime}\sim\text{\rm Unif}(S_{n})$ of $\pi$, for (31) it holds that
$R_{\ell}$ $=\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left\{\mathbb{E}_{X_{1},\ldots,X_{\ell}\sim\overline{P}}\left[\prod_{i=1}^{\ell}\frac{\mathrm{d}(P_{\pi(i)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{i})\prod_{i=1}^{\ell}\frac{\mathrm{d}(P_{\pi^{\prime}(i)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{i})\right]\right\}$
$=\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{\ell}\left(\int\frac{\mathrm{d}(P_{\pi(i)}-\overline{P})\times\mathrm{d}(P_{\pi^{\prime}(i)}-\overline{P})}{\mathrm{d}\overline{P}}\right)\right]$
$=\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{\ell}\left(\int\frac{\mathrm{d}P_{\pi(i)}\mathrm{d}P_{\pi^{\prime}(i)}}{\mathrm{d}\overline{P}}-1\right)\right]$
=(26)${=}\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{\ell}(nA_{\pi(i),\pi^{\prime}(i)}-1)\right]$
$=n^{\ell}\cdot\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{\ell}\overline{A}_{\pi(i),\pi^{\prime}(i)}\right]$
=$\overset{\rm(a)}{=}\frac{n^{\ell}}{\binom{n}{\ell}^{2}\ell!}\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=\ell\end{subarray}}\mathrm{Perm}(\overline{A}_{S,S^{\prime}})=\frac{n^{\ell}}{\binom{n}{\ell}^{2}\ell!}\cdot T_{\ell},$
where (a) uses $\mathbb{P}(\pi([\ell])=S)=\mathbb{P}(\pi^{\prime}([\ell])=S^{\prime})=\binom{n}{\ell}^{-1}$ for all $S,S^{\prime}$ of size $\ell$ and averages over $\ell!$ bijections between $S$ and $S^{\prime}$.
Next we prove the relation between $S_{\ell}$ and $T_{\ell}$. Fix any $t\in\mathbb{R}$, let $D_{t}:=\mathrm{diag}(\lambda_{1},t\lambda_{2},\ldots,t\lambda_{n})$. By the Cauchy–Binet formula for permanents [MM65, Theorem 2.1], it holds that
$\mathrm{Perm}(UD_{t}U^{\top})$ $=\sum_{\begin{subarray}{c}(\ell_{1},\ldots,\ell_{n})\in\mathbb{N}^{n}\\ \sum_{i=1}^{n}\ell_{i}=n\end{subarray}}\frac{\mathrm{Perm}(U_{\ell_{1},\ldots,\ell_{n}})^{2}}{\ell_{1}!\cdots\ell_{n}!}\lambda_{1}^{\ell_{1}}(t\lambda_{2})^{\ell_{2}}\cdots(t\lambda_{n})^{\ell_{n}}$
$=\sum_{\ell=0}^{n}\sum_{\begin{subarray}{c}(\ell_{2},\ldots,\ell_{n})\in\mathbb{N}^{n-1}\\ \sum_{i=2}^{n}\ell_{i}=\ell\end{subarray}}\frac{\mathrm{Perm}(U_{n-\ell,\ell_{2},\ldots,\ell_{n}})^{2}}{(n-\ell)!\ell_{2}!\cdots\ell_{n}!}(t\lambda_{2})^{\ell_{2}}\cdots(t\lambda_{n})^{\ell_{n}}=\frac{n!}{n^{n}}\sum_{\ell=0}^{n}S_{\ell}t^{\ell}$
is a polynomial in $t$. An alternative expression of $\mathrm{Perm}(UD_{t}U^{\top})$ is also possible: by the structure of leading eigenvalue and eigenvector in Lemma 4.2, $UD_{t}U^{\top}=U(tD-(t-1)\mathrm{diag}(1,0,\ldots,0))U^{\top}=tA-\frac{t-1}{n}J=t\overline{A}+\frac{J}{n}$ for $J:={\bf 1}{\bf 1}^{\top}$ being the all-one matrix. Consequently,
$\mathrm{Perm}(UD_{t}U^{\top})$ $=\mathrm{Perm}\left(t\overline{A}+\frac{J}{n}\right)$
$=\sum_{S,S^{\prime}\subseteq[n]:|S|=|S^{\prime}|}\mathrm{Perm}\left((t\overline{A})_{S,S^{\prime}}\right)\mathrm{Perm}\left((\frac{J}{n})_{[n]\backslash S,[n]\backslash S^{\prime}}\right)$
$=\sum_{\ell=0}^{n}t^{\ell}\frac{(n-\ell)!}{n^{n-\ell}}\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=\ell\end{subarray}}\mathrm{Perm}(\overline{A}_{S,S^{\prime}})=\sum_{\ell=0}^{n}\frac{(n-\ell)!}{n^{n-\ell}}T_{\ell}t^{\ell}.$
The above two expressions must give the same polynomial in $t$, so $S_{\ell}=\frac{(n-\ell)!}{n!}n^{\ell}\cdot T_{\ell}$. ∎
Appendix B Proof of Theorem 3.3
Since the complex case (24) has been established in Section 3.3, in this section we are devoted to the proof of (25) in the real case. We split the analysis into several steps.
B.1 The case of binary support
We first consider the special case where the coordinates of $x\in\mathbb{R}^{n}$ can only take two values. Thanks to the assumptions $\sum_{i=1}^{n}x_{i}=0$ and $\sum_{i=1}^{n}x_{i}^{2}=n$, the only cases are
$x^{(k)}=\Big{(}\underbrace{\sqrt{\frac{k}{n-k}},\ldots,\sqrt{\frac{k}{n-k}}}_{n-k\text{ copies}},\underbrace{-\sqrt{\frac{n-k}{k}},\ldots,-\sqrt{\frac{n-k}{k}}}_{k\text{ copies}}\Big{)},\quad\text{for some }k=1,\cdots,n-1.$
The target of this section is to show that
$|e_{\ell}(x^{(k)})|\leq\sqrt{10\binom{n}{\ell}}\qquad\text{for all }k,\ell\in[n-1].$ (34)
Fix the choice of $r=\sqrt{\ell/(n-\ell)}$. By Cauchy’s formula,
$|e_{\ell}(x^{(k)})|$ $=\left|\frac{1}{2\pi\mathrm{i}}\oint_{|z|=r}\left(1+z\sqrt{\frac{k}{n-k}}\right)^{n-k}\left(1-z\sqrt{\frac{n-k}{k}}\right)^{k}\frac{\mathrm{d}z}{z^{\ell+1}}\right|$
$\overset{\rm(a)}{\leq}\frac{1}{2\pi r^{\ell}}\int_{0}^{2\pi}\left(1+\frac{k}{n-k}r^{2}+2r\sqrt{\frac{k}{n-k}}\cos\theta\right)^{\frac{n-k}{2}}\left(1+\frac{n-k}{k}r^{2}-2r\sqrt{\frac{n-k}{k}}\cos\theta\right)^{\frac{k}{2}}\mathrm{d}\theta$
=$\overset{\rm(b)}{=}\frac{1}{\pi r^{\ell}}\int_{-1}^{1}\left(1+\frac{k}{n-k}r^{2}+2r\sqrt{\frac{k}{n-k}}t\right)^{\frac{n-k}{2}}\left(1+\frac{n-k}{k}r^{2}-2r\sqrt{\frac{n-k}{k}}t\right)^{\frac{k}{2}}\frac{\mathrm{d}t}{\sqrt{1-t^{2}}},$
where (a) uses the triangle inequality and a change of variable $z=re^{\mathrm{i}\theta}$ with $\theta\in[0,2\pi)$, and (b) applies another change of variable $t=\cos\theta\in[-1,1]$. We break the integral into two parts:
  • Fixing some $\delta\in\left[0,\frac{1}{2}\right]$ to be specified later, define
    $A_{1}(\delta):=\int_{1-\delta\leq|t|\leq 1}\left(1+\frac{k}{n-k}r^{2}+2r\sqrt{\frac{k}{n-k}}t\right)^{\frac{n-k}{2}}\left(1+\frac{n-k}{k}r^{2}-2r\sqrt{\frac{n-k}{k}}t\right)^{\frac{k}{2}}\frac{\mathrm{d}t}{\sqrt{1-t^{2}}}.$
    Note that by the AM-GM inequality, for all $t\in[-1,1]$ it holds that
    $\left(1+\frac{k}{n-k}r^{2}+2r\sqrt{\frac{k}{n-k}}t\right)^{\frac{n-k}{2}}\left(1+\frac{n-k}{k}r^{2}-2r\sqrt{\frac{n-k}{k}}t\right)^{\frac{k}{2}}$
    $\leq\left[\frac{n-k}{n}\left(1+\frac{k}{n-k}r^{2}+2r\sqrt{\frac{k}{n-k}}t\right)+\frac{k}{n}\left(1+\frac{n-k}{k}r^{2}-2r\sqrt{\frac{n-k}{k}}t\right)\right]^{\frac{n}{2}}$
    $=(1+r^{2})^{\frac{n}{2}}.$
    Therefore,
    $A_{1}(\delta)\leq(1+r^{2})^{\frac{n}{2}}\int_{1-\delta\leq|t|\leq 1}\frac{\mathrm{d}t}{\sqrt{1-t^{2}}}=2(1+r^{2})^{\frac{n}{2}}\arccos(1-\delta)\leq\frac{2\pi}{3}(1+r^{2})^{\frac{n}{2}}\sqrt{2\delta},$ (35)
    where the last step follows from $1-\cos t\geq\frac{1}{2}\left(\frac{3t}{\pi}\right)^{2}$ for all $t\in\left[0,\frac{\pi}{3}\right]$.
  • For the remaining integral
    $A_{2}(\delta):=\int_{-1+\delta}^{1-\delta}\left(1+\frac{k}{n-k}r^{2}+2r\sqrt{\frac{k}{n-k}}t\right)^{\frac{n-k}{2}}\left(1+\frac{n-k}{k}r^{2}-2r\sqrt{\frac{n-k}{k}}t\right)^{\frac{k}{2}}\frac{\mathrm{d}t}{\sqrt{1-t^{2}}},$
    we note that $\sqrt{1-t^{2}}\geq\sqrt{1-(1-\delta)^{2}}\geq\sqrt{3\delta/2}$ for $\delta\in\left[0,\frac{1}{2}\right]$. In addition,
    $t_{-}:=-\frac{1+\frac{k}{n-k}r^{2}}{2r\sqrt{\frac{k}{n-k}}}\leq-1,\qquad t_{+}:=\frac{1+\frac{n-k}{k}r^{2}}{2r\sqrt{\frac{n-k}{k}}}\geq 1$
    by AM-GM, with
    $t_{+}-t_{-}=\frac{n}{\sqrt{k(n-k)}}\cdot\frac{1+r^{2}}{2r}.$
    Using a change of variable $t=t_{-}+(t_{+}-t_{-})u$ for $u\in[0,1]$, we upper bound $A_{2}(\delta)$ as
    $A_{2}(\delta)$ $\leq\sqrt{\frac{2}{3\delta}}\int_{-1+\delta}^{1-\delta}\left(1+\frac{k}{n-k}r^{2}+2r\sqrt{\frac{k}{n-k}}t\right)^{\frac{n-k}{2}}\left(1+\frac{n-k}{k}r^{2}-2r\sqrt{\frac{n-k}{k}}t\right)^{\frac{k}{2}}\mathrm{d}t$
    $\leq\sqrt{\frac{2}{3\delta}}\int_{t_{-}}^{t_{+}}\left(1+\frac{k}{n-k}r^{2}+2r\sqrt{\frac{k}{n-k}}t\right)^{\frac{n-k}{2}}\left(1+\frac{n-k}{k}r^{2}-2r\sqrt{\frac{n-k}{k}}t\right)^{\frac{k}{2}}\mathrm{d}t$
    $=\sqrt{\frac{2}{3\delta}}(t_{+}-t_{-})^{\frac{n}{2}+1}\int_{0}^{1}\left(2r\sqrt{\frac{k}{n-k}}u\right)^{\frac{n-k}{2}}\left(2r\sqrt{\frac{n-k}{k}}(1-u)\right)^{\frac{k}{2}}\mathrm{d}u$
    $=\sqrt{\frac{2}{3\delta}}\frac{(1+r^{2})^{\frac{n}{2}+1}}{2r}\cdot\frac{n^{\frac{n}{2}+1}}{k^{\frac{k+1}{2}}(n-k)^{\frac{n-k+1}{2}}}\frac{\Gamma\left(\frac{k}{2}+1\right)\Gamma\left(\frac{n-k}{2}+1\right)}{\Gamma\left(\frac{n}{2}+2\right)},$
    where $\Gamma(x)=\int_{0}^{\infty}t^{x-1}e^{-t}\mathrm{d}t$ is the Gamma function. Using Stirling’s approximation
    $\sqrt{2\pi x}\left(\frac{x}{e}\right)^{x}\leq\Gamma(x+1)\leq\sqrt{2\pi x}\left(\frac{x}{e}\right)^{x}\cdot e^{\frac{1}{12x}},\qquad x>0,$
    we get
    $\frac{n^{\frac{n}{2}+1}}{k^{\frac{k+1}{2}}(n-k)^{\frac{n-k+1}{2}}}\frac{\Gamma\left(\frac{k}{2}+1\right)\Gamma\left(\frac{n-k}{2}+1\right)}{\Gamma\left(\frac{n}{2}+2\right)}$
    $\leq\frac{n^{\frac{n}{2}+1}}{k^{\frac{k+1}{2}}(n-k)^{\frac{n-k+1}{2}}}\cdot\frac{\sqrt{\pi k}\left(\frac{k}{2e}\right)^{\frac{k}{2}}\sqrt{\pi(n-k)}\left(\frac{n-k}{2e}\right)^{\frac{n-k}{2}}}{\left(\frac{n}{2}+1\right)\sqrt{\pi n}\left(\frac{n}{2e}\right)^{\frac{n}{2}}}\exp\left(\frac{1}{6k}+\frac{1}{6(n-k)}\right)$
    $\leq\frac{e^{1/3}\sqrt{\pi n}}{\frac{n}{2}+1}\leq\frac{2e^{1/3}\sqrt{\pi}}{\sqrt{n}},$
    which is an upper bound independent of $k$. Plugging it into the upper bound of $A_{2}(\delta)$ gives
    $A_{2}(\delta)\leq e^{1/3}\sqrt{\frac{2\pi}{3\delta}}\frac{(1+r^{2})^{\frac{n}{2}+1}}{r\sqrt{n}}<\frac{2\pi}{3}\frac{(1+r^{2})^{\frac{n}{2}+1}}{r\sqrt{n\delta}}.$ (36)
Combining (35) and (36), we get
$|e_{\ell}(x^{(k)})|\leq\frac{2(1+r^{2})^{\frac{n}{2}}}{3r^{\ell}}\left(\sqrt{2\delta}+\frac{1+r^{2}}{r\sqrt{n\delta}}\right)=\frac{2(1+r^{2})^{\frac{n}{2}}}{3r^{\ell}}\left(\sqrt{2\delta}+\frac{1}{\sqrt{\delta}}\cdot\sqrt{\frac{n}{\ell(n-\ell)}}\right),$
where the final step plugs in the choice of $r=\sqrt{\ell/(n-\ell)}$. Choosing
$\delta=\frac{1}{2}\sqrt{\frac{n}{2\ell(n-\ell)}}\leq\frac{1}{2},$
together with Stirling’s approximation, leads to the upper bound
$|e_{\ell}(x^{(k)})|$ $\leq\frac{2(1+r^{2})^{\frac{n}{2}}}{r^{\ell}}\left(\frac{n}{2\ell(n-\ell)}\right)^{\frac{1}{4}}=\frac{2n^{\frac{n}{2}}}{\ell^{\frac{\ell}{2}}(n-\ell)^{\frac{n-\ell}{2}}}\left(\frac{n}{2\ell(n-\ell)}\right)^{\frac{1}{4}}$
$\leq 2^{\frac{3}{4}}\left(\frac{n^{n+\frac{1}{2}}}{\ell^{\ell+\frac{1}{2}}(n-\ell)^{n-\ell+\frac{1}{2}}}\right)^{\frac{1}{2}}\leq 2^{\frac{3}{4}}\left(\sqrt{2\pi}e^{1/3}\binom{n}{\ell}\right)^{\frac{1}{2}}<\sqrt{10\binom{n}{\ell}},$
which proves the desired result (34).
Remark B.1.
The following insights of the saddle point method are used in the above proof. Note that $e_{\ell}(x)$ is the coefficient of $z^{\ell}$ in $f(z)=\prod_{i=1}^{n}(1+x_{i}z)$, with corresponding saddle point equation
$\frac{\ell}{z}=\frac{\mathrm{d}}{\mathrm{d}z}\log f(z)=\sum_{i=1}^{n}\frac{x_{i}}{1+x_{i}z}.$
Specializing to the vector $x=x^{(k)}$ gives the saddle points
$z_{\pm}=\frac{1}{2(n-\ell)}\left(-\frac{\ell(n-2k)}{\sqrt{k(n-k)}}\pm\sqrt{\frac{(n-2k)^{2}}{k(n-k)}\ell^{2}-4\ell(n-\ell)}\right).$
We observe that whenever the discriminant (the term under the squared root) is not positive, both saddle points $z_{\pm}$ have magnitude $\sqrt{\ell/(n-\ell)}$, motivating our choice of $r=\sqrt{\ell/(n-\ell)}$ in the proof. In addition, the Laplace approximation around $z_{\pm}$ suggests that $|e_{\ell}(x^{(k)})|$ attains the maximum when $z_{+}$ and $z_{-}$ coincide on the real axis, motivating us to decompose the integral into two quantities $A_{1}(\delta)$ and $A_{2}(\delta)$ as in (35) and (36).
B.2 Reduction to binary support
In this section, we show that the upper bound of $|e_{\ell}(x)|$ for general $x\in\mathbb{R}^{n}$ could be reduced to the case of binary support in (34). Specifically, for $\ell\geq 4$, we will prove by induction on $n\geq\ell$ that
$|e_{\ell}(x)|\leq\sqrt{10\binom{n}{\ell}},\qquad\text{for all }x\in\mathbb{R}^{n}\text{ with }\sum_{i=1}^{n}x_{i}=0,\sum_{i=1}^{n}x_{i}^{2}=n.$ (37)
The remaining cases $\ell\leq 3$ will be dealt with in the next section.
The base case $n=\ell$ of (37) is clear: by the AM-GM inequality,
$|e_{\ell}(x)|^{2}=\prod_{i=1}^{\ell}x_{i}^{2}\leq\left(\frac{1}{\ell}\sum_{i=1}^{\ell}x_{i}^{2}\right)^{\ell}=1.$
For the inductive step, assume that $n\geq\ell+1$ and the induction hypothesis (37) holds for all values smaller than $n$. As the constraint set of (37) is compact, the optimization program (37) must admit a maximizer $x^{\star}$. By the method of Lagrangian multipliers (for maximizing $e_{\ell}(x)$ or $-e_{\ell}(x)$), for any maximizer $x^{\star}$, there must exist $\lambda,\mu\in\mathbb{R}$ such that
$e_{\ell-1}\left(x^{\star}\backslash\left\{x_{i}^{\star}\right\}\right)=\lambda x_{i}^{\star}+\mu,\qquad\text{for all }i\in[n].$ (38)
Based on (38), we split the analysis into three steps.
Step I: if $x_{i}^{\star}=0$ for some $i\in[n]$.
In this case, the rescaled vector $\sqrt{1-n^{-1}}(x^{\star}\backslash\left\{x_{i}^{\star}\right\})$ satisfies the constraint in (37) with $n$ replaced by $n-1$. By the induction hypothesis,
$|e_{\ell}(x^{\star})|=\left(\frac{n}{n-1}\right)^{\frac{\ell}{2}}\left|e_{\ell}\left(\sqrt{1-n^{-1}}(x^{\star}\backslash\left\{x_{i}^{\star}\right\})\right)\right|\leq\left(\frac{n}{n-1}\right)^{\frac{\ell}{2}}\sqrt{10\binom{n-1}{\ell}}\leq\sqrt{10\binom{n}{\ell}},$
where the last step follows from simple algebra and Bernoulli’s inequality $\left(1-\frac{1}{n}\right)^{\ell}\geq 1-\frac{\ell}{n}$. Therefore, the inductive hypothesis (37) holds for $n$ in this case, and we are done. Subsequently, we may assume that every maximizer $x^{\star}$ of (37) must have $x_{i}^{\star}\neq 0$ for all $i\in[n]$.
Step II: show that $|\mathrm{supp}(x^{\star})|\leq 3$.
Assume by contradiction that $|\mathrm{supp}(x^{\star})|\geq 4$, and that the values of $(x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star})$ are distinct. Choosing $i=1,2$ in (38), a subtraction gives
$(x_{2}^{\star}-x_{1}^{\star})e_{\ell-2}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star}\right\})=\lambda(x_{1}^{\star}-x_{2}^{\star})\Longrightarrow e_{\ell-2}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star}\right\})=-\lambda.$
Similarly we also have $e_{\ell-2}(x^{\star}\backslash\left\{x_{1}^{\star},x_{3}^{\star}\right\})=-\lambda$, and a further subtraction gives
$(x_{3}^{\star}-x_{2}^{\star})e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star}\right\})=0\Longrightarrow e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star}\right\})=0.$ (39)
Again, subtracting (39) with $e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{4}^{\star}\right\})=0$ leads to
$(x_{4}^{\star}-x_{3}^{\star})e_{\ell-4}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star}\right\})=0\Longrightarrow e_{\ell-4}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star}\right\})=0.$ (40)
Moreover, since
$e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star}\right\})=x_{4}^{\star}e_{\ell-4}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star}\right\})+e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star}\right\}),$
by (39) and (40) we also have
$e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star}\right\})=0.$ (41)
Based on (40) and (41), we invoke the following result in [GY14, Fact B], which is a property for real-rooted polynomials.
Lemma B.2.
For a real vector $x\in\mathbb{R}^{n}$ and $10\leq k\leq n-$ 1, if $e_{k}(x)=e_{k+1}(x)=0$, then $e_{\ell}(x)=0$ for all $\ell\geq k$.
As $\ell\geq 4$ and $n\geq\ell+1$, Lemma B.2 applied to (40) and (41) gives $\prod_{i=5}^{n}x_{i}^{\star}=e_{n-4}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star}\right\})=0$, i.e. one of $(x_{5}^{\star},\ldots,x_{n}^{\star})$ must be zero. However, it is assumed at the end of Step I that $x_{i}^{\star}\neq 0$ for all $i\in[n]$, a contradiction! So $|\mathrm{supp}(x^{\star})|\leq 3$, as desired.
Step III: show that $|\mathrm{supp}(x^{\star})|\leq 2$.
We proceed to show that $|\mathrm{supp}(x^{\star})|=3$ is also impossible. Assume by contradiction that $|\mathrm{supp}(x^{\star})|=3$, and the values of $(x_{1}^{\star},x_{2}^{\star},x_{3}^{\star})$ are distinct. We propose to find a triple $(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$ of distinct elements such that
$x_{1}+x_{2}+x_{3}$ $=x_{1}^{\star}+x_{2}^{\star}+x_{3}^{\star},$ (42)
$x_{1}x_{2}+x_{2}x_{3}+x_{1}x_{3}$ $=x_{1}^{\star}x_{2}^{\star}+x_{2}^{\star}x_{3}^{\star}+x_{1}^{\star}x_{3}^{\star},$ (43)
$\{x_{1},x_{2},x_{3}\}\cap\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},0\}$ ==$\varnothing$.= ∅ . (44)
We show that such a triple $(x_{1},x_{2},x_{3})$ exists. Note that (42) and (43) define an intersection of a hyperplane and a sphere in $\mathbb{R}^{3}$, which is a circle. This circle is nondegenerate (and thus has infinitely many points on it), for $(x_{1}^{\star},x_{2}^{\star},x_{3}^{\star})$ and $(x_{2}^{\star},x_{1}^{\star},x_{3}^{\star})$ are two distinct points on this circle. In addition, this circle belongs to none of the hyperplanes $\{x_{i}=x_{j}\}$, $\{x_{i}=x_{j}^{\star}\}$, or $\{x_{i}=0\}$, so each hyperplane only intersects the circle at finitely many points. Therefore, we can choose any point $(x_{1},x_{2},x_{3})$ on the circle other than the above intersections, and this triple satisfies distinctness and (42)–(44).
Given such a triple, we define a new vector $x^{\prime}=(x_{1},x_{2},x_{3},x_{4}^{\star},\ldots,x_{n}^{\star})\in\mathbb{R}^{n}$. By (42) and (43), it is clear that $x^{\prime}$ satisfies the constraints in (37). In addition, viewing both $e_{\ell}(x^{\prime})$ and $e_{\ell}(x^{\star})$ as polynomials of the first three elements, we get
$e_{\ell}(x^{\prime})-e_{\ell}(x^{\star})$ $=e_{\ell-1}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star}\right\})(x_{1}+x_{2}+x_{3}-x_{1}^{\star}-x_{2}^{\star}-x_{3}^{\star})$
$\quad+e_{\ell-2}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star}\right\})(x_{1}x_{2}+x_{2}x_{3}+x_{1}x_{3}-x_{1}^{\star}x_{2}^{\star}-x_{2}^{\star}x_{3}^{\star}-x_{1}^{\star}x_{3}^{\star})$
$\quad+e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star}\right\})(x_{1}x_{2}x_{3}-x_{1}^{\star}x_{2}^{\star}x_{3}^{\star})$
$=0,$
where the last step is due to (42), (43), and (39). In other words, $x^{\prime}$ is also a maximizer of $|e_{\ell}(\cdot)|$, while $|\mathrm{supp}(x^{\prime})|\geq 4$ and $x^{\prime}$ does not contain zero by distinctness of $(x_{1},x_{2},x_{3})$ and (44). Proceeding as Step II would lead to the conclusion $|\mathrm{supp}(x^{\prime})|\leq 3$, which is a contradiction. Therefore, we have established that $|\mathrm{supp}(x^{\star})|\leq 2$, and the desired result (37) follows from the inequality (34) in the binary case.
B.3 Remaining corner cases
The only cases not covered in (37) are the scenarios $\ell\leq 3$, which we handle separately.
  • For $\ell\in\{0,1\}$, the results $e_{0}(x)=1$ and $e_{1}(x)=0$ are trivial.
  • For $\ell=2$, Newton’s identity gives that
    $|e_{2}(x)|=\frac{1}{2}\left|e_{1}(x)^{2}-\sum_{i=1}^{n}x_{i}^{2}\right|=\frac{n}{2}\leq\sqrt{\binom{n}{2}}.$
  • For $\ell=3$, Newton’s identity gives that
    $e_{3}(x)$ $=\frac{1}{3}\left(e_{1}(x)\left(e_{2}(x)-\sum_{i=1}^{n}x_{i}^{2}\right)+\sum_{i=1}^{n}x_{i}^{3}\right)=\frac{1}{3}\sum_{i=1}^{n}x_{i}^{3}.$
    As $\sum_{i=1}^{n}x_{i}^{2}=n$, we have $|x_{i}|\leq\sqrt{n}$ for all $i\in[n]$, and consequently
    $|e_{3}(x)|\leq\frac{\sqrt{n}}{3}\sum_{i=1}^{n}x_{i}^{2}=\frac{n^{\frac{3}{2}}}{3}\leq\sqrt{3\binom{n}{3}}.$
In all these scenarios, the inequality (25) of Theorem 3.3 holds. The proof of Theorem 3.3 is therefore complete.
Appendix C Deferred proofs
C.1 Proof of Corollary 1.3
By Theorem 1.2, it suffices to upper bound ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ and $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ for the given families ${\mathcal{P}}$. While the computation of $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ is typically straightforward, the evaluation of ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ may require some effort. We present a useful lemma to upper bound ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$.
Lemma C.1.
Let ${\mathcal{P}}_{1},\ldots,{\mathcal{P}}_{m}$ be families of probability distributions over the same space. Then
${\mathsf{C}}_{\chi^{2}}\left(\bigcup_{i=1}^{m}{\mathcal{P}}_{i}\right)\leq\sum_{i=1}^{m}{\mathsf{C}}_{\chi^{2}}\left({\mathcal{P}}_{i}\right)+m-1.$
Proof.
By induction it suffices to prove the lemma for $m=2$. In addition, since ${\mathsf{C}}_{\chi^{2}}\left({\mathcal{P}}\right)\leq{\mathsf{C}}_{\chi^{2}}\left({\mathcal{P}}^{\prime}\right)$ for ${\mathcal{P}}\subseteq{\mathcal{P}}^{\prime}$, without loss of generality we may assume that ${\mathcal{P}}_{1}$ and ${\mathcal{P}}_{2}$ are disjoint.
Let $\rho$ be a probability distribution over ${\mathcal{P}}_{1}\cup{\mathcal{P}}_{2}$, with $\rho({\mathcal{P}}_{1})=a$ and $\rho({\mathcal{P}}_{2})=1-a$. Let $\rho_{1},\rho_{2}$ be the restriction (conditional distribution) of $\rho$ to ${\mathcal{P}}_{1}$ and ${\mathcal{P}}_{2}$, respectively, it holds that
$\rho=a\rho_{1}+(1-a)\rho_{2}.$
We upper bound the $\chi^{2}$ mutual information $I_{\chi^{2}}(P;X)$ as follows:
$I_{\chi^{2}}(P;X)$ $=\mathbb{E}_{P\sim\rho}\left[\chi^{2}(P\|\mathbb{E}_{P^{\prime}\sim\rho}[P^{\prime}])\right]$
$=\mathbb{E}_{P\sim\rho}\left[\int\frac{(\mathrm{d}P)^{2}}{\mathbb{E}_{P^{\prime}\sim\rho}[\mathrm{d}P^{\prime}]}\right]-1$
$=a\mathbb{E}_{P\sim\rho_{1}}\left[\int\frac{(\mathrm{d}P)^{2}}{\mathbb{E}_{P^{\prime}\sim\rho}[\mathrm{d}P^{\prime}]}\right]+(1-a)\mathbb{E}_{P\sim\rho_{2}}\left[\int\frac{(\mathrm{d}P)^{2}}{\mathbb{E}_{P^{\prime}\sim\rho}[\mathrm{d}P^{\prime}]}\right]-1$
$\leq a\mathbb{E}_{P\sim\rho_{1}}\left[\int\frac{(\mathrm{d}P)^{2}}{a\mathbb{E}_{P^{\prime}\sim\rho_{1}}[\mathrm{d}P^{\prime}]}\right]+(1-a)\mathbb{E}_{P\sim\rho_{2}}\left[\int\frac{(\mathrm{d}P)^{2}}{(1-a)\mathbb{E}_{P^{\prime}\sim\rho_{2}}[\mathrm{d}P^{\prime}]}\right]-1$
$=\mathbb{E}_{P\sim\rho_{1}}\left[\chi^{2}(P\|\mathbb{E}_{P^{\prime}\sim\rho_{1}}[P^{\prime}])\right]+\mathbb{E}_{P\sim\rho_{2}}\left[\chi^{2}(P\|\mathbb{E}_{P^{\prime}\sim\rho_{2}}[P^{\prime}])\right]+1$
$\leq{\mathsf{C}}_{\chi^{2}}\left({\mathcal{P}}_{1}\right)+{\mathsf{C}}_{\chi^{2}}\left({\mathcal{P}}_{2}\right)+1.$
Taking the supremum over $\rho\in\Delta({\mathcal{P}}_{1}\cup{\mathcal{P}}_{2})$ completes the proof. ∎
Next we prove Corollary 1.3.
Gaussian family ${\mathcal{P}}=\left\{{\mathcal{N}}(\theta,1):|\theta|\leq\mu\right\}$.
Since
$\chi^{2}\left({\mathcal{N}}(\theta,1)\|{\mathcal{N}}(\theta^{\prime},1)\right)=\exp\left((\theta-\theta^{\prime})^{2}\right)-1,$
it is clear that ${\mathsf{D}}_{\chi^{2}}({\mathcal{P}})=\exp(4\mu^{2})-1$. We show that ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})=O(\mu\wedge\mu^{2})$. For $\mu\leq 1$, we simply use ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})=O(\mu^{2})$. For $\mu>1$, we split the class ${\mathcal{P}}$ into ${\mathcal{P}}\subseteq\cup_{m\in\mathbb{Z},|m|\leq\mu+1}{\mathcal{P}}_{m}$, with ${\mathcal{P}}_{m}:=\left\{{\mathcal{N}}(\theta,1):m\leq\theta<m+1\right\}$. Clearly ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}}_{m})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}}_{m})=e-1=O(1)$, so Lemma C.1 yields
${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq\sum_{m\in\mathbb{Z},|m|\leq\mu+1}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}}_{m})+\left|\left\{m\in\mathbb{Z}:|m|\leq\mu+1\right\}\right|-1=O(\mu).$
Then the claimed result follows from (4) and (6).
Gaussian family with finite support ${\mathcal{P}}=\left\{{\mathcal{N}}(\theta,1):|\theta|\leq\mu,\theta\in\Theta\right\}$, $|\Theta|<\infty$.
We only need to prove an additional upper bound ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq|\Theta|-1$. In fact, let $\Theta=\{\theta_{1},\ldots,\theta_{m}\}$, Lemma C.1 leads to
${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq\sum_{i=1}^{m}{\mathsf{C}}_{\chi^{2}}\left(\left\{{\mathcal{N}}(\theta_{i},1)\right\}\right)+m-1=m-1.$
Bernoulli family ${\mathcal{P}}=\{\text{\rm Bern}(p):p\in[\varepsilon,1-\varepsilon]\}$.
For $\varepsilon\geq\frac{1}{4}$,
${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})=\chi^{2}\left(\text{\rm Bern}(\varepsilon)\|\text{\rm Bern}(1-\varepsilon)\right)=\frac{(1-2\varepsilon)^{2}}{\varepsilon(1-\varepsilon)}=O\left((1-2\varepsilon)^{2}\right),$
so that Equation 4 gives the first upper bound. For the second upper bound, we show that ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq 1-2\varepsilon$ for general $\varepsilon>0$, so that (4) gives
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})\leq 10\sum_{\ell=2}^{n}(1-2\varepsilon)^{\ell}\leq 10\sum_{\ell=2}^{\infty}(1-2\varepsilon)^{\ell}=O\left(\frac{1}{\varepsilon}\right).$
To see this, for any prior $\rho\in\Delta([\varepsilon,1-\varepsilon])$, we upper bound the $\chi^{2}$ mutual information as
$I_{\chi^{2}}(p;X)$ $=\mathbb{E}_{p\sim\rho}\left[\chi^{2}\left(\text{\rm Bern}(p)\|\text{\rm Bern}(\mathbb{E}_{p^{\prime}\sim\rho}[p^{\prime}]\right)\right]$
$=\frac{\mathbb{E}_{p\sim\rho}[p^{2}]}{\mathbb{E}_{p\sim\rho}[p]}+\frac{\mathbb{E}_{p\sim\rho}[(1-p)^{2}]}{\mathbb{E}_{p\sim\rho}[1-p]}-1$
$\leq 1-\varepsilon+1-\varepsilon-1=1-2\varepsilon.$
Taking the supremum over the prior $\rho$ leads to ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq 1-2\varepsilon$.
Poisson family ${\mathcal{P}}=\left\{\text{\rm Poi}(\lambda):\lambda\in[0,M]\right\}$.
First, we have $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})=e^{M}$, for
$\sup_{P_{1},P_{2}\in{\mathcal{P}}}\frac{1}{\int\sqrt{\mathrm{d}P_{1}\mathrm{d}P_{2}}}=\frac{1}{\int\sqrt{\mathrm{d}\text{\rm Poi}(0)\mathrm{d}\text{\rm Poi}(M)}}=e^{M/2}.$
Next we prove ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})=O(\sqrt{M}\wedge M)$, by considering the cases $M\leq 1$ and $M>1$. For $M\leq 1$, pick any prior $\rho$ on $[0,M]$. We upper bound the $\chi^{2}$ mutual information as
$I_{\chi^{2}}(\lambda;X)$ $=\mathbb{E}_{\lambda\sim\rho}\left[\chi^{2}\left(\text{\rm Poi}(\lambda)\|\mathbb{E}_{\lambda^{\prime}\sim\rho}[\text{\rm Poi}(\lambda^{\prime})]\right)\right]$
$=\frac{\mathbb{E}_{\lambda\sim\rho}\left[\mathbb{P}(\text{\rm Poi}(\lambda)=0)^{2}\right]}{\mathbb{E}_{\lambda\sim\rho}\left[\mathbb{P}(\text{\rm Poi}(\lambda)=0)\right]}+\sum_{k=1}^{\infty}\frac{\mathbb{E}_{\lambda\sim\rho}\left[\mathbb{P}(\text{\rm Poi}(\lambda)=k)^{2}\right]}{\mathbb{E}_{\lambda\sim\rho}\left[\mathbb{P}(\text{\rm Poi}(\lambda)=k)\right]}-1$
$\overset{\rm(c)}{\leq}1+\sum_{k=1}^{\infty}\mathbb{P}(\text{\rm Poi}(M)=k)-1=1-e^{-M}\leq M,$
where (c) uses the monotonicity $\mathbb{P}(\text{\rm Poi}(\lambda)=k)\leq\mathbb{P}(\text{\rm Poi}(M)=k)$ for all $\lambda\leq M\leq 1$ and $k\geq 1$. Taking the supremum over the prior $\rho$ gives that ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq M$. For $M>1$, we write ${\mathcal{P}}=\cup_{i=1}^{m}{\mathcal{P}}_{i}$, with
${\mathcal{P}}_{i}:=\left\{\text{\rm Poi}(\lambda):(i-1)^{2}\leq\lambda\leq i^{2}\wedge M\right\}.$
Clearly we can choose $m=O(\sqrt{M})$ to ensure that ${\mathcal{P}}=\cup_{i=1}^{m}{\mathcal{P}}_{i}$. In addition, our previous argument shows that ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}}_{1})\leq 1$, and for $i\geq 2$,
${\mathsf{C}}_{\chi^{2}}({\mathcal{P}}_{i})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}}_{i})\leq\max_{(i-1)^{2}\leq\lambda_{1},\lambda_{2}\leq i^{2}}\exp\left(\frac{(\lambda_{1}-\lambda_{2})^{2}}{\lambda_{2}}\right)-1\leq\exp\left(\frac{(i^{2}-(i-1)^{2})^{2}}{(i-1)^{2}}\right)-1=O(1).$
Therefore, by Lemma C.1, we have
${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq\sum_{i=1}^{m}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}}_{i})+m-1=O(\sqrt{M}).$
Then the claimed result follows from (4) and (5).
C.2 Proof of Corollary 1.4
For the first inequality (1), it is clear that
$nI(\vartheta_{1};X_{1})-I(\vartheta;X)$ $=nI(\vartheta_{1};X_{1})-\mathbb{E}_{(\vartheta,X)}\left[\log\frac{\prod_{i=1}^{n}\mathrm{d}P_{\vartheta_{i}}}{\mathrm{d}P_{X}}\right]$
$=nI(\vartheta_{1};X_{1})-\mathbb{E}_{(\vartheta,X)}\left[\log\frac{\prod_{i=1}^{n}\mathrm{d}P_{\vartheta_{i}}}{\mathrm{d}Q_{X}}\right]+\mathbb{E}_{P_{X}}\left[\log\frac{\mathrm{d}P_{X}}{\mathrm{d}Q_{X}}\right]$
$=nI(\vartheta_{1};X_{1})-n\mathbb{E}_{(\vartheta_{1},X_{1})}\left[\log\frac{\mathrm{d}P_{\vartheta_{1}}}{\mathrm{d}Q_{X_{1}}}\right]+{\sf KL}(P_{X}\|Q_{X})$
=$\overset{\rm(a)}{=}{\sf KL}(P_{X}\|Q_{X}),$
where $P_{X},Q_{X}$ are the marginal distributions of the random vector $X$ under the permutation prior and the i.i.d. prior of $\vartheta$, respectively, and (a) uses $P_{X_{1}}=Q_{X_{1}}$. So $nI(\vartheta_{1};X_{1})-I(\vartheta;X)\geq 0$, and
$nI(\vartheta_{1};X_{1})-I(\vartheta;X)$ $\leq\log\left(1+\chi^{2}(P_{X}\|Q_{X})\right)$
$\leq\min\left\{10\sum_{\ell=2}^{n}I_{\chi^{2}}(\vartheta_{1};X_{1})^{\ell},I_{\chi^{2}}(\vartheta_{1};X_{1})(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))\right\},$
where we use Theorem 1.2 and replace the channel capacity ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ by a tighter quantity $I_{\chi^{2}}(\vartheta_{1};X_{1})$, which is an intermediate step in the proof of Theorem 1.2.
For the second inequality (8), let $Q_{1},\dots,Q_{m}$ be an $\varepsilon$-cover of $\mathrm{conv}({\mathcal{P}})$ under the KL divergence, with a fixed $\varepsilon>0$ and $m=N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)$. By the standard entropic upper bound of the mutual information (cf. [YB99] or [PW24, Theorem 32.4]), one has
$I(\theta;X)$ $\leq\mathbb{E}_{\theta}\left[\min_{i\in[m]}{\sf KL}(P_{X|\theta}\|Q_{i}^{\otimes n})\right]+\log m$
=$\overset{\rm(b)}{=}\mathbb{E}_{\theta}\left[{\sf KL}(P_{X|\theta}\|Q_{X|\theta})+\min_{i\in[m]}\mathbb{E}_{P_{X|\theta}}\left[\log\frac{\mathrm{d}Q_{X|\theta}}{\mathrm{d}Q_{i}^{\otimes n}}\right]\right]+\log m$
=$\overset{\rm(c)}{=}\mathbb{E}_{\theta}\left[{\sf KL}(P_{X|\theta}\|Q_{X|\theta})+\min_{i\in[m]}{\sf KL}(Q_{X|\theta}\|Q_{i}^{\otimes n})\right]+\log m$
$\overset{\rm(d)}{\leq}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))+n\varepsilon^{2}+\log m,$
where in (b) $Q_{X|\theta}$ denotes the i.i.d. distribution $\left(\frac{1}{n}\sum_{i=1}^{n}P_{\theta_{i}}\right)^{\otimes n}$, (c) uses the i.i.d. structures for both $Q_{X|\theta}$ and $Q_{i}^{\otimes n}$, and notes that the marginal distributions $P_{X_{1}|\theta}$ and $Q_{X_{1}|\theta}$ are the same, and (d) follows from Theorem 1.2 and the definition of the KL covering applied to $\frac{1}{n}\sum_{i=1}^{n}P_{\theta_{i}}\in\mathrm{conv}({\mathcal{P}})$. Taking the infimum over $\varepsilon>0$ then completes the proof of (8).
C.3 Proof of Theorem 1.5
Since $P_{X^{n}}$ is exchangeable, it can be written as a mixture $P_{X^{n}}=\mathbb{E}_{\lambda}\left[\nu_{\mathbb{P}}(\theta_{1,\lambda},\ldots,\theta_{n,\lambda})\right]$, where $\nu_{\mathbb{P}}(\theta_{1},\ldots,\theta_{n})$ denotes the discrete probability measure on $\mathbb{R}^{n}$ induced by sampling without replacement:
$\nu_{\mathbb{P}}(\theta_{1},\ldots,\theta_{n})(x)=\frac{1}{n!}\sum_{\pi\in S_{n}}\prod_{i=1}^{n}{\mathbf{1}_{\left\{{x_{i}=\theta_{\pi(i)}}\right\}}}.$
By joint convexity of the $\chi^{2}$-divergence, it suffices to deal with the scenario $P_{X^{n}}=\nu_{\mathbb{P}}(\theta_{1},\ldots,\theta_{n})$, in which case $P_{Y^{n}}$ is a permutation mixture with components $(P_{1},\ldots,P_{n}):=({\mathsf{K}}(\cdot|\theta_{1}),\ldots,{\mathsf{K}}(\cdot|\theta_{n}))$. The marginal distribution $P_{Y^{k}}$ is then
$\mathrm{d}P_{Y^{k}}(y^{k})=\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{k}\mathrm{d}P_{\pi(i)}(y_{i})\right].$
We will show that the product distribution $Q_{Y^{k}}=\overline{P}^{\otimes k}$ achieves the theorem statement, where as usual $\overline{P}:=\frac{1}{n}\sum_{i=1}^{n}P_{i}$. Specifically, the $\chi^{2}$-divergence is
$\chi^{2}(P_{Y^{k}}\|Q_{Y^{k}})+1$ $=\mathbb{E}_{Q_{Y^{k}}}\left[\left(\frac{\mathrm{d}P_{Y^{k}}}{\mathrm{d}Q_{Y^{k}}}\right)^{2}\right]$
$=\mathbb{E}_{Y_{1},\ldots,Y_{k}\sim\overline{P}}\left[\left(\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{k}\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(Y_{i})\right]\right)^{2}\right]$
=$\overset{\rm(a)}{=}\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{k}\mathbb{E}_{Y_{i}\sim\overline{P}}\left[\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(Y_{i})\frac{\mathrm{d}P_{\pi^{\prime}(i)}}{\mathrm{d}\overline{P}}(Y_{i})\right]\right]$
=$\overset{\rm(b)}{=}\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[n^{k}\prod_{i=1}^{k}A_{\pi(i),\pi^{\prime}(i)}\right]$
=$\overset{\rm(c)}{=}\frac{n^{k}}{\binom{n}{k}^{2}k!}\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=k\end{subarray}}\mathrm{Perm}(A_{S,S^{\prime}}),$
where in (a) we introduce an independent copy $\pi^{\prime}\sim\text{\rm Unif}(S_{n})$ of $\pi$, (b) recalls the definition of $A$ in (26), and (c) uses $\mathbb{P}(\pi([k])=S)=\mathbb{P}(\pi^{\prime}([k])=S^{\prime})=\binom{n}{k}^{-1}$ for all subsets $S,S^{\prime}$ of size $k$ and averages over $k!$ bijections between $S$ and $S^{\prime}$. Next we relate $A$ to its centered version $\overline{A}=A-\frac{1}{n}{\bf 1}{\bf 1}^{\top}$. Let $J={\bf 1}{\bf 1}^{\top}$ denote the all-one matrix, then
$\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=k\end{subarray}}\mathrm{Perm}(A_{S,S^{\prime}})$ $=\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=k\end{subarray}}\mathrm{Perm}\left(\overline{A}_{S,S^{\prime}}+\frac{J_{S,S^{\prime}}}{n}\right)$
$=\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=k\end{subarray}}\sum_{\begin{subarray}{c}T\subseteq S,T^{\prime}\subseteq S^{\prime}\\ |T|=|T^{\prime}|\end{subarray}}\mathrm{Perm}(\overline{A}_{T,T^{\prime}})\mathrm{Perm}\left(\frac{J_{S\backslash T,S^{\prime}\backslash T^{\prime}}}{n}\right)$
$=\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=k\end{subarray}}\sum_{\ell=0}^{k}\sum_{\begin{subarray}{c}T\subseteq S,T^{\prime}\subseteq S^{\prime}\\ |T|=|T^{\prime}|=\ell\end{subarray}}\mathrm{Perm}(\overline{A}_{T,T^{\prime}})\frac{(k-\ell)!}{n^{k-\ell}}$
=$\overset{\rm(d)}{=}\sum_{\ell=0}^{k}\sum_{\begin{subarray}{c}T,T^{\prime}\subseteq[n]\\ |T|=|T^{\prime}|=\ell\end{subarray}}\binom{n-\ell}{k-\ell}^{2}\mathrm{Perm}(\overline{A}_{T,T^{\prime}})\frac{(k-\ell)!}{n^{k-\ell}},$
where (d) swaps the sum and counts the number of $S\supseteq T,S^{\prime}\supseteq T^{\prime}$ of size $k$. After some algebra, the above two identities give that
$\chi^{2}(P_{Y^{k}}\|Q_{Y^{k}})+1$ $=\frac{n^{\ell}k!}{[n!]^{2}}\sum_{\ell=0}^{k}\frac{[(n-\ell)!]^{2}}{(k-\ell)!}\sum_{\begin{subarray}{c}T,T^{\prime}\subseteq[n]\\ |T|=|T^{\prime}|=\ell\end{subarray}}\mathrm{Perm}(\overline{A}_{T,T^{\prime}})$
=$\overset{\rm(e)}{=}\frac{n^{\ell}k!}{[n!]^{2}}\sum_{\ell=0}^{k}\frac{[(n-\ell)!]^{2}}{(k-\ell)!}\cdot\frac{n!}{(n-\ell)!n^{\ell}}S_{\ell}=\sum_{\ell=0}^{k}\frac{\binom{k}{\ell}}{\binom{n}{\ell}}S_{\ell},$
where (e) follows from the identity between $T_{\ell}$ and $S_{\ell}$ in Lemma A.1. Finally, as $S_{0}=1$, $S_{1}=0$, and $S_{\ell}\geq 0$,
$\chi^{2}(P_{Y^{k}}\|Q_{Y^{k}})=\sum_{\ell=2}^{k}\frac{\binom{k}{\ell}}{\binom{n}{\ell}}S_{\ell}\leq\frac{k(k-1)}{n(n-1)}\sum_{\ell=2}^{n}S_{\ell}\overset{\rm(f)}{=}\frac{k(k-1)}{n(n-1)}\chi^{2}(P_{Y^{n}}\|Q_{Y^{n}}),$
where (f) is due to the identity (27). Now the claimed result follows from the upper bounds of $\chi^{2}(P_{Y^{n}}\|Q_{Y^{n}})$ in Theorem 1.2.
C.4 Proof of Corollary 1.7
We first show that
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)\leq M\sum_{i=1}^{n}\mathbb{E}\left[{\sf TV}(P_{\vartheta_{i}|X_{i}},P_{\vartheta_{i}|X})\right].$ (45)
In fact,
$r^{\mathrm{S}}(\theta)$ $=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{S}}}\mathbb{E}\left[L(\vartheta,\widehat{\vartheta})\right]=\inf_{\Delta(\cdot)}\sum_{i=1}^{n}\mathbb{E}\left[\ell(\vartheta_{i},\Delta(X_{i}))\right]=\mathbb{E}\left[\sum_{i=1}^{n}\inf_{u}\mathbb{E}_{\vartheta_{i}|X_{i}}\left[\ell(\vartheta_{i},u)\right]\right],$
$r^{\mathrm{PI}}(\theta)$ $=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{PI}}}\mathbb{E}\left[L(\vartheta,\widehat{\vartheta})\right]=\inf_{\widehat{\vartheta}\in{\mathcal{D}}^{\mathrm{PI}}}\sum_{i=1}^{n}\mathbb{E}\left[\ell(\vartheta_{i},\widehat{\vartheta}_{i}(X))\right]=\mathbb{E}\left[\sum_{i=1}^{n}\inf_{u}\mathbb{E}_{\vartheta_{i}|X}\left[\ell(\vartheta_{i},u)\right]\right],$
and the optimal decision rules are the Bayes optimal decision rules under the posterior distributions $P_{\vartheta_{i}|X_{i}}$ and $P_{\vartheta_{i}|X}$, respectively, for the loss $\ell$. Consequently, as $0\leq\ell(\cdot,\cdot)\leq M$,
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)$ $\leq\sum_{i=1}^{n}\mathbb{E}\left[\sup_{u}\left|\mathbb{E}_{\vartheta_{i}|X_{i}}\left[\ell(\vartheta_{i},u)\right]-\mathbb{E}_{\vartheta_{i}|X}\left[\ell(\vartheta_{i},u)\right]\right|\right]$
$\leq M\sum_{i=1}^{n}\mathbb{E}\left[{\sf TV}(P_{\vartheta_{i}|X_{i}},P_{\vartheta_{i}|X})\right],$
establishing (45). To proceed, note that by the Bayes rule, we have
$\frac{\mathrm{d}P_{\vartheta_{i}|X_{i}}}{\mathrm{d}P_{\vartheta_{i}|X}}=\frac{\mathrm{d}P_{X_{\backslash i}|X_{i}}}{\mathrm{d}P_{X_{\backslash i}|\vartheta_{i}}},$
and therefore
$\mathbb{E}\left[{\sf TV}(P_{\vartheta_{i}|X_{i}},P_{\vartheta_{i}|X})\right]$ $=\mathbb{E}_{\vartheta_{i},X}\left[\left|\frac{\mathrm{d}P_{X_{\backslash i}|X_{i}}}{\mathrm{d}P_{X_{\backslash i}|\vartheta_{i}}}-1\right|\right]=\mathbb{E}_{\vartheta_{i},X_{i}}\left[{\sf TV}(P_{X_{\backslash i}|X_{i}},P_{X_{\backslash i}|\vartheta_{i}})\right]$
$\overset{\rm(a)}{\leq}\mathbb{E}_{X_{i}}\mathbb{E}_{\vartheta_{i},\vartheta_{i}^{\prime}|X_{i}}\left[{\sf TV}(P_{X_{\backslash i}|\vartheta_{i}^{\prime}},P_{X_{\backslash i}|\vartheta_{i}})\right]\leq\max_{\vartheta_{i},\vartheta_{i}^{\prime}}{\sf TV}(P_{X_{\backslash i}|\vartheta_{i}},P_{X_{\backslash i}|\vartheta_{i}^{\prime}})$
$\overset{\rm(b)}{\leq}\sqrt{\frac{3{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}}{n-1}}.$
Here (a) follows from convexity and $P_{X_{\backslash i}|X_{i}}=\mathbb{E}_{\vartheta_{i}^{\prime}|X_{i}}\left[P_{X_{\backslash i}|\vartheta_{i}^{\prime}}\right]$, where $\vartheta_{i}^{\prime}$ is an independent copy of $\vartheta_{i}$ given $X_{i}$. For (b), note that $P_{X_{\backslash i}|\vartheta_{i}}$ is the permutation mixture based on $\{P_{\theta_{1}},\dots,P_{\theta_{n}}\}\backslash\{P_{\vartheta_{i}}\}$, so that Theorem 1.6 gives the claimed upper bound. Plugging this upper bound into (45) and using $n\geq 2$ leads to the claimed corollary.
C.5 Proof of Lemma 4.1
By the definition of $\chi^{2}$-divergence,
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})+1=\mathbb{E}_{\mathbb{Q}_{n}}\left[\left(\frac{\mathrm{d}\mathbb{P}_{n}}{\mathrm{d}\mathbb{Q}_{n}}\right)^{2}\right]=\mathbb{E}_{X_{1},\ldots,X_{n}\sim\overline{P}}\left[\left(\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{n}\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\right]\right)^{2}\right].$
By introducing an independent copy $\pi^{\prime}\sim\text{\rm Unif}(S_{n})$ of $\pi$, it holds that
$\chi^{2}(\mathbb{P}_{n}\|\mathbb{Q}_{n})+1$ $=\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left\{\mathbb{E}_{X_{1},\ldots,X_{n}\sim\overline{P}}\left[\prod_{i=1}^{n}\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\prod_{i=1}^{\ell}\frac{\mathrm{d}P_{\pi^{\prime}(i)}}{\mathrm{d}\overline{P}}(X_{i})\right]\right\}$
$=\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{n}\left(\int\frac{\mathrm{d}P_{\pi(i)}\mathrm{d}P_{\pi^{\prime}(i)}}{\mathrm{d}\overline{P}}\right)\right]$
=(26)${=}\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{n}(nA_{\pi(i),\pi^{\prime}(i)})\right]=\frac{n^{n}}{n!}\mathrm{Perm}(A),$
as claimed.
C.6 Proof of Lemma 4.2
The definition of $\overline{P}=\frac{1}{n}\sum_{i=1}^{n}P_{i}$ trivially implies that $A$ is doubly stochastic. To see why $A$ is PSD, simply note that
$A_{ij}=\frac{1}{n}\mathbb{E}_{X\sim\overline{P}}\left[\frac{\mathrm{d}P_{i}}{\mathrm{d}\overline{P}}(X)\frac{\mathrm{d}P_{j}}{\mathrm{d}\overline{P}}(X)\right]$
could be written as an inner product of $\frac{\mathrm{d}P_{i}}{\mathrm{d}\overline{P}}$ and $\frac{\mathrm{d}P_{j}}{\mathrm{d}\overline{P}}$. The trace upper bound is also straightforward:
$\mathrm{Tr}(A)=\frac{1}{n}\sum_{i=1}^{n}\int\frac{(\mathrm{d}P_{i})^{2}}{\mathrm{d}\overline{P}}=\frac{1}{n}\sum_{i=1}^{n}\chi^{2}(P_{i}\|\overline{P})+1\leq{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})+1.$
As for the eigenstructure of $A$, the same leading eigenvalue/eigenvector holds for any doubly stochastic matrix. The only non-trivial property is the spectral gap. To this end, note that the Laplacian $L=I-A$ satisfies that
$x^{\top}Lx=\sum_{1\leq i<j\leq n}A_{ij}(x_{i}-x_{j})^{2}.$
For each entry $A_{ij}$, we can lower bound it as
$A_{ij}=\frac{1}{n}\int\frac{\mathrm{d}P_{i}\mathrm{d}P_{j}}{\mathrm{d}\overline{P}}\overset{\rm(a)}{\geq}\frac{1}{n}\left(\int\sqrt{\mathrm{d}P_{i}\mathrm{d}P_{j}}\right)^{2}=\frac{1}{n}\left(1-\frac{H^{2}(P_{i},P_{j})}{2}\right)^{2}\geq\frac{1}{n\mathsf{\Delta}_{H^{2}}({\mathcal{P}})},$
where (a) is due to the Cauchy–Schwarz inequality. Consequently, for any unit vector $x\in\mathbb{R}^{n}$ with ${\bf 1}^{\top}x=0$, we have
$x^{\top}Lx\geq\frac{1}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}\sum_{1\leq i<j\leq n}(x_{i}-x_{j})^{2}=\frac{1}{2\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}\sum_{i,j=1}^{n}(x_{i}-x_{j})^{2}=\frac{1}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})},$
which proves the spectral gap lower bound.
C.7 Proof of Lemma 4.4
Let $P=(P_{ij})_{i\in[m],j\in[n]}$ and $z=(z_{1},\ldots,z_{n})$. Then
$\mathbb{E}\left[\prod_{i=1}^{m}|(Pz)_{i}|^{2}\right]$ $=\mathbb{E}\left[\prod_{i=1}^{m}\left|\sum_{j=1}^{n}P_{ij}z_{j}\right|^{2}\right]$
$=\mathbb{E}\left[\prod_{i=1}^{m}\left(\sum_{j=1}^{n}P_{ij}z_{j}\right)\left(\sum_{j=1}^{n}P_{ij}\bar{z}_{j}\right)\right]$
=$\overset{\rm(a)}{=}\sum_{\pi\in S_{m}}\prod_{i=1}^{m}\mathbb{E}\left[\left(\sum_{j=1}^{n}P_{ij}z_{j}\right)\left(\sum_{j=1}^{n}P_{\pi(i)j}\bar{z}_{j}\right)\right]$
$=\sum_{\pi\in S_{m}}\prod_{i=1}^{m}\left(\sum_{j=1}^{n}P_{ij}P_{\pi(i)j}\right)$
$=\sum_{\pi\in S_{m}}\prod_{i=1}^{m}(PP^{\top})_{i\pi(i)}=\mathrm{Perm}(PP^{\top}),$
where (a) is due to Isserlis’ theorem [Iss18].
C.8 Proof of Lemma 4.6
Let $v=(v_{1},\cdots,v_{n}):=\widetilde{P}z\sim{\mathcal{C}}{\mathcal{N}}(0,\widetilde{P}\widetilde{P}^{\top})$, where
$\widetilde{P}\widetilde{P}^{\top}=\widetilde{U}\widetilde{D}\widetilde{U}^{\top}=A-\lambda_{1}u_{1}u_{1}^{\top}=A-\frac{1}{n}{\bf 1}{\bf 1}^{\top}=\overline{A},$
where $\lambda_{1}=1$ and $u_{1}={\bf 1}/\sqrt{n}$ are the leading eigenvalue and eigenvector of $A$, respectively (cf. Lemma 4.2), and $\overline{A}$ is the centered version of $A$ defined above (33). Consequently,
$\mathbb{E}\left[|e_{\ell}(v)|^{2}\right]$ $=\mathbb{E}\left[\left|\sum_{S\subseteq[n]:|S|=\ell}\prod_{i\in S}v_{i}\right|^{2}\right]=\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=\ell\end{subarray}}\mathbb{E}\left[\prod_{i\in S}v_{i}\prod_{j\in S^{\prime}}\bar{v}_{j}\right]$
=$\overset{\rm(a)}{=}\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=\ell\end{subarray}}\sum_{\pi:S\leftrightarrow S^{\prime}}\prod_{i\in S}\mathbb{E}\left[v_{i}\bar{v}_{\pi(i)}\right]$
[…]
where (a) is due to Isserlis’ theorem [Iss18] and the observation $\mathbb{E}[v_{i}v_{j}]=\mathbb{E}[\bar{v}_{i}\bar{v}_{j}]=0$ for all $i,j\in[n]$ (here $\pi:S\leftrightarrow S^{\prime}$ denotes that $\pi$ is a bijection between $S$ and $S^{\prime}$), and (b) follows from $v\sim{\mathcal{C}}{\mathcal{N}}(0,\overline{A})$. Then the claimed result is a direct consequence of the identity between $S_{\ell}$ and $T_{\ell}$ in Lemma A.1.
C.9 Proof of Lemma 4.7
By the Cauchy–Schwarz inequality, for any probability measure $\mathbb{R}_{n}$ it holds that
${\sf TV}(\mathbb{P}_{n},\mathbb{Q}_{n})^{2}=\left(\frac{1}{2}\int|\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{Q}_{n}|\right)^{2}\leq\frac{1}{4}\int\frac{(\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{Q}_{n})^{2}}{\mathrm{d}\mathbb{R}_{n}}.$
Choosing $\mathbb{R}_{n}=\overline{P}^{\otimes n}$ completes the proof of the first claim. The likelihood ratio between the signed measure $\mathbb{P}_{n}-\mathbb{Q}_{n}$ and the product measure $\overline{P}^{\otimes n}$ is computed as
$\frac{\mathrm{d}(\mathbb{P}_{n}-\mathbb{Q}_{n})}{\mathrm{d}\overline{P}^{\otimes n}}(X)$ =$\overset{\rm(a)}{=}\mathbb{E}_{I\sim\text{\rm Unif}([n]),\pi:[n]\backslash\{I\}\leftrightarrow\{2,\ldots,n\}}\left[\frac{\mathrm{d}(P_{1}-Q_{1})}{\mathrm{d}\overline{P}}(X_{I})\cdot\prod_{i\neq I}\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\right]$
=$\overset{\rm(b)}{=}\mathbb{E}_{I\sim\text{\rm Unif}([n]),\pi:[n]\backslash\{I\}\leftrightarrow\{2,\ldots,n\}}\left[\frac{\mathrm{d}(P_{1}-Q_{1})}{\mathrm{d}\overline{P}}(X_{I})\cdot\sum_{T\subseteq[n]\backslash\{I\}}\prod_{i\in T}\frac{\mathrm{d}(P_{\pi(i)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{i})\right]$
=$\overset{\rm(c)}{=}\sum_{S\subseteq[n]:|S|\geq 1}\left(\frac{1}{n}\sum_{i\in S}\frac{\mathrm{d}(P_{1}-Q_{1})}{\mathrm{d}\overline{P}}(X_{i})\cdot\mathbb{E}_{\pi:[n]\backslash\{i\}\leftrightarrow\{2,\ldots,n\}}\prod_{j\in S\backslash\{i\}}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})\right)$
$=:\sum_{S\subseteq[n]:|S|\geq 1}f_{S}(X),$
where (a) decomposes the random permutation over $S_{n}$ into a random choice of $I=\pi^{-1}(1)$ and a random bijection between the rest (i.e. $[n]\backslash\{i\}$ and $\{2,\cdots,n\}$), and (b) is the identity $\prod_{i=1}^{n}(1+x_{i})=\sum_{T\subseteq[n]}\prod_{i\in T}x_{i}$. In (c), we expand the expectation over $I\sim\text{\rm Unif}([n])$ and swap the sum as $\sum_{i=1}^{n}\sum_{T\subseteq[n]\backslash\{i\}}g_{i}(T)=\sum_{S\subseteq[n]:|S|\geq 1}\sum_{i\in S}g_{i}(S\backslash\{i\})$.
Next, we argue that the functions $\{f_{S}(X):S\subseteq[n],|S|\geq 1\}$ are orthogonal under $\overline{P}^{\otimes n}$. Since $f_{S}(X)$ could be expressed as $\frac{1}{n}\sum_{i\in S}\mathbb{E}_{\pi\sim\mu_{i}}\left[g_{S,i,\pi}(X)\right]$ for some probability measure $\mu_{i}$ and function $g_{S,i,\pi}(X)$, it suffices to show the orthogonality between $g_{S,i,\pi}(X)$ and $g_{T,i^{\prime},\pi^{\prime}}(X)$ for all $S\neq T$ and $(i,i^{\prime},\pi,\pi^{\prime})$. To see so, simply note that
$g_{S,i,\pi}(X)=\frac{\mathrm{d}(P_{1}-Q_{1})}{\mathrm{d}\overline{P}}(X_{i})\prod_{j\in S\backslash\{i\}}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})$
is a product of zero-mean functions of $\{X_{j}:j\in S\}$ under $\overline{P}^{\otimes n}$, so the orthogonality follows. This orthogonality yields
$\int\frac{(\mathrm{d}\mathbb{P}_{n}-\mathrm{d}\mathbb{Q}_{n})^{2}}{\mathrm{d}\overline{P}^{\otimes n}}=\mathbb{E}_{\overline{P}^{\otimes n}}\left[\left(\sum_{S\subseteq[n]:|S|\geq 1}f_{S}(X)\right)^{2}\right]=\sum_{S\subseteq[n]:|S|\geq 1}\mathbb{E}_{\overline{P}^{\otimes n}}\left[f_{S}(X)^{2}\right]$
=$\overset{\rm(d)}{=}\sum_{\ell=1}^{n}\binom{n}{\ell}\mathbb{E}_{\overline{P}^{\otimes n}}\left[f_{[\ell]}(X)^{2}\right]$
$\overset{\rm(e)}{\leq}\sum_{\ell=1}^{n}\binom{n}{\ell}\frac{\ell}{n^{2}}\sum_{i=1}^{\ell}\mathbb{E}_{X_{1},\ldots,X_{\ell}\sim\overline{P}}\left[\left(\frac{\mathrm{d}(P_{1}-Q_{1})}{\mathrm{d}\overline{P}}(X_{i})\cdot\mathbb{E}_{\pi:[n]\backslash\{i\}\leftrightarrow\{2,\ldots,n\}}\prod_{j\in[\ell]\backslash\{i\}}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})\right)^{2}\right]$
=$\overset{\rm(f)}{=}\sum_{\ell=1}^{n}\binom{n}{\ell}\frac{\ell^{2}}{n^{2}}\int\frac{(\mathrm{d}P_{1}-\mathrm{d}Q_{1})^{2}}{\mathrm{d}\overline{P}}\cdot\mathbb{E}_{X_{2},\ldots,X_{\ell}\sim\overline{P}}\left[\left(\mathbb{E}_{\pi:\{2,\ldots,n\}\leftrightarrow\{2,\ldots,n\}}\prod_{j=2}^{\ell}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})\right)^{2}\right],$
where (d) notes that the second moment of $f_{S}$ only depends on the size of $S$, (e) uses the Cauchy–Schwarz inequality $(\sum_{i=1}^{\ell}x_{i})^{2}\leq\ell\sum_{i=1}^{\ell}x_{i}^{2}$, and (f) notes that the inner expectation does not depend on $i$. To proceed, note that
$\int\frac{(\mathrm{d}P_{1}-\mathrm{d}Q_{1})^{2}}{\mathrm{d}\overline{P}}$ $\leq 2\int\frac{(\mathrm{d}P_{1}-\mathrm{d}\overline{P})^{2}}{\mathrm{d}\overline{P}}+2\int\frac{(\mathrm{d}Q_{1}-\mathrm{d}\overline{P})^{2}}{\mathrm{d}\overline{P}}$
$=2(\chi^{2}(P_{1}\|\overline{P})+\chi^{2}(Q_{1}\|\overline{P}))\leq 4{\mathsf{D}}_{\chi^{2}}({\mathcal{P}}),$
and
$\mathbb{E}_{X_{2},\ldots,X_{\ell}\sim\overline{P}}\left[\left(\mathbb{E}_{\pi:\{2,\ldots,n\}\leftrightarrow\{2,\ldots,n\}}\prod_{j=2}^{\ell}\frac{\mathrm{d}(P_{\pi(j)}-\overline{P})}{\mathrm{d}\overline{P}}(X_{j})\right)^{2}\right]=R_{\ell-1}=\frac{S_{\ell-1}}{\binom{n-1}{\ell-1}}$
by the definition of $R_{\ell}$ in (31) and the identity between $R_{\ell}$ and $S_{\ell}$ in Lemma A.1, where the matrix $A\in\mathbb{R}^{(n-1)\times(n-1)}$ is now constructed from $(P_{2},\ldots,P_{n})$. A combination of all the above then gives
${\sf TV}(\mathbb{P}_{n},\mathbb{Q}_{n})^{2}\leq\frac{1}{4}\sum_{\ell=1}^{n}\binom{n}{\ell}\frac{\ell^{2}}{n^{2}}\cdot 4{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})\frac{S_{\ell-1}}{\binom{n-1}{\ell-1}}=\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}\sum_{\ell=1}^{n}\ell S_{\ell-1},$
as claimed.
C.10 Proof of Lemma 5.1
Fix any $n$ and the choices $P_{1},\dots,P_{n}\in{\mathcal{P}}$. For a large integer $m$, consider the probability measures $\mathbb{P}_{mn}$ and $\mathbb{Q}_{mn}$ based on $mn$ distributions $\left\{P_{1},\dots,P_{1},P_{2},\dots,P_{2},\dots,P_{n}\right\}$, where each $P_{i}$ appears $m$ times. Let $\overline{P}=\frac{1}{n}\sum_{i=1}^{n}P_{i}$, and $f\in L^{2}(\overline{P})$ be any function with $\mathbb{E}_{\overline{P}}[f]=0$. We use $f$ as the test function, and study the distributions of the test statistic $\frac{1}{\sqrt{mn}}\sum_{i=1}^{mn}f(X_{i})$ under both $\mathbb{P}_{mn}$ and $\mathbb{Q}_{mn}$, denoted by $\mathbb{P}_{f,m}$ and $\mathbb{Q}_{f,m}$, respectively.
Under $\mathbb{P}_{mn}$, it is clear that
$\frac{1}{\sqrt{mn}}\sum_{i=1}^{mn}f(X_{i})\overset{\mathrm{d}}{=}\frac{1}{\sqrt{mn}}\sum_{i=1}^{n}\sum_{j=1}^{m}f(Z_{i,j}),$
where $Z_{i,1},\dots,Z_{i,m}\overset{\mathrm{i.i.d.}}{\sim}P_{i}$ for all $i\in[n]$, and are mutually independent for different $i\in[n]$. By CLT, it is then clear that
$\mathbb{P}_{f,m}\rightsquigarrow{\mathcal{N}}\left(0,\frac{1}{n}\sum_{i=1}^{n}\mathrm{Var}_{P_{i}}(f)\right)\qquad\text{as }m\to\infty,$
where $\rightsquigarrow$ denotes the weak convergence of probability measures. Under $\mathbb{Q}_{mn}$, it is clear that $X_{1},\dots,X_{mn}$ follow an i.i.d. distribution $\overline{P}$, so CLT gives
$\mathbb{Q}_{f,m}\rightsquigarrow{\mathcal{N}}\left(0,\mathrm{Var}_{\overline{P}}(f)\right)\qquad\text{as }m\to\infty.$
Based on the limiting distributions, the $\chi^{2}$-divergence can be lower bounded as
$\liminf_{m\to\infty}\chi^{2}(\mathbb{P}_{mn}\|\mathbb{Q}_{mn})$ $\overset{\rm(a)}{\geq}\liminf_{m\to\infty}\chi^{2}(\mathbb{P}_{f,m}\|\mathbb{Q}_{f,m})$
$\overset{\rm(b)}{\geq}\chi^{2}\left({\mathcal{N}}\left(0,\frac{1}{n}\sum_{i=1}^{n}\mathrm{Var}_{P_{i}}(f)\right)\|{\mathcal{N}}\left(0,\mathrm{Var}_{\overline{P}}(f)\right)\right)$
=$\overset{\rm(c)}{=}\left[1-\left(1-\frac{\frac{1}{n}\sum_{i=1}^{n}\mathrm{Var}_{P_{i}}(f)}{\mathrm{Var}_{\overline{P}}(f)}\right)^{2}\right]^{-\frac{1}{2}}-1,$
where (a) is due to the data-processing inequality, (b) is the lower-semicontinuity of the $\chi^{2}$ divergence (using arguments similar to [PW24, Theorem 4.9]), and (c) uses
$\chi^{2}({\mathcal{N}}(0,\sigma_{1}^{2})\|{\mathcal{N}}(0,\sigma_{2}^{2}))=\sqrt{\frac{\sigma_{2}^{4}}{\sigma_{1}^{2}(2\sigma_{2}^{2}-\sigma_{1}^{2})}}-1=\left(1-\left(1-\frac{\sigma_{1}^{2}}{\sigma_{2}^{2}}\right)^{2}\right)^{-\frac{1}{2}}-1$
as long as $\sigma_{2}\geq\sigma_{1}$, where we also note that $\mathrm{Var}_{\overline{P}}(f)\geq\frac{1}{n}\sum_{i=1}^{n}\mathrm{Var}_{P_{i}}(f)$ by the concavity of variance. Consequently, it remains to evaluate
$S:=\sup_{f\in L^{2}(\overline{P}):\mathbb{E}_{\overline{P}}[f]=0}\left(1-\frac{\frac{1}{n}\sum_{i=1}^{n}\mathrm{Var}_{P_{i}}(f)}{\mathrm{Var}_{\overline{P}}(f)}\right)=\sup_{f\in L^{2}(\overline{P}):\mathbb{E}_{\overline{P}}[f]=0}\frac{\frac{1}{n}\sum_{i=1}^{n}\left(\mathbb{E}_{P_{i}}[f]\right)^{2}}{\mathbb{E}_{\overline{P}}[f^{2}]}.$
We show that $S\geq\lambda_{2}(A)$, the second largest eigenvalue of the matrix $A$ constructed from (26). To this end, we choose
$f=\sum_{j=1}^{n}u_{j}\frac{\mathrm{d}P_{j}}{\mathrm{d}\overline{P}},$
with $u=(u_{1},\dots,u_{n})$ being the unit-length eigenvector of $A$ associated with $\lambda_{2}(A)$. Clearly $\mathbb{E}_{\overline{P}}[f]=\sum_{j=1}^{n}u_{j}=0$ by the orthogonality of $u$ and $\bf 1$ in Lemma 4.2, and
$\frac{1}{n}\sum_{i=1}^{n}\left(\mathbb{E}_{P_{i}}[f]\right)^{2}$ $=\frac{1}{n}\sum_{i=1}^{n}\left(n\sum_{j=1}^{n}A_{ij}u_{j}\right)^{2}=\frac{1}{n}\sum_{i=1}^{n}(n\lambda_{2}(A)u_{i})^{2}=n\lambda_{2}(A)^{2}$
$\mathbb{E}_{\overline{P}}[f^{2}]$ $=n\sum_{i=1}^{n}\sum_{j=1}^{n}A_{ij}u_{i}u_{j}=n\sum_{i=1}^{n}\lambda_{2}(A)u_{i}^{2}=n\lambda_{2}(A).$
This shows that $S\geq\lambda_{2}(A)$ (in fact $S=\lambda_{2}(A)$ holds, but we do not need this upper bound).
A combination of the above steps then gives
$\liminf_{m\to\infty}\chi^{2}(\mathbb{P}_{mn}\|\mathbb{Q}_{mn})\geq\frac{1}{\sqrt{1-\lambda_{2}(A)^{2}}}-1,$
and taking the supremum over the choice of $n$ and $P_{1},\dots,P_{n}\in{\mathcal{P}}$ leads to the first result. For the second result, pick $n=2$, so that
$\lambda_{2}(A)=\mathrm{Tr}(A)-1=\frac{1}{2}\int\frac{(\mathrm{d}P_{1}-\mathrm{d}P_{2})^{2}}{\mathrm{d}P_{1}+\mathrm{d}P_{2}}={\sf LC}(P_{1},P_{2})\geq\frac{H^{2}(P_{1},P_{2})}{2}.$
Here the last step follows from [PW24, Eqn. (7.35)]. Consequently, the second statement follows from
$\frac{1}{\sqrt{1-\lambda_{2}(A)^{2}}}-1\geq\frac{\lambda_{2}(A)^{2}}{2},\quad\text{and}\quad\frac{1}{\sqrt{1-\lambda_{2}(A)^{2}}}-1\geq\frac{1}{\sqrt{2(1-\lambda_{2}(A))}}-1,$
and taking the supremum over $P_{1},P_{2}\in{\mathcal{P}}$.
C.11 Proof of Lemma 5.2
The constants ${\mathsf{C}}_{0}\geq 3$ and $\Delta_{0}\leq\frac{1}{4}$ will be specified later in the proof. Given ${\mathsf{C}}\geq{\mathsf{C}}_{0}$ and $\Delta\leq\Delta_{0}$, construct a matrix $A\in\mathbb{R}^{mn\times mn}$ with
$A_{ij}=\begin{cases}\frac{\Delta}{mn}+\frac{1-\Delta}{n}&\text{if }i,j\in[(k-1)n+1,kn]\text{ for some }k\in[m],\\ \frac{\Delta}{mn}&\text{otherwise,}\end{cases}$
with $m=\lceil{\mathsf{C}}\rceil$, and $n\in\mathbb{N}$ to be specified later. In other words, $A$ is the Kronecker product
$A=\left(\frac{\Delta}{m}J_{m}+(1-\Delta)I_{m}\right)\otimes\frac{J_{n}}{n},$
where $I$ and $J$ are the identity and all-one matrices, respectively. By simple algebra, $A$ is doubly stochastic, and has eigenvalues $\lambda_{1}(A)=1,\lambda_{2}(A)=\cdots=\lambda_{m}(A)=1-\Delta$, and $\lambda_{m+1}(A)=\cdots=\lambda_{mn}(A)=0$. In particular, $\mathrm{Tr}(A)\leq m<1+{\mathsf{C}}$, and the spectral gap of $A$ is $\Delta$.
It remains to lower bound the permanent $\mathrm{Perm}(A)$. Restricting only to the diagonal blocks, one has
$\mathrm{Perm}(A)\geq(n!)^{m}\left(\frac{\Delta}{mn}+\frac{1-\Delta}{n}\right)^{mn}\geq(n!)^{m}\left(\frac{1-\Delta}{n}\right)^{mn}.$
By Stirling’s approximation $\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}\leq n!\leq\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}\exp\left(\frac{1}{12n}\right)$, we have
$\mathrm{Perm}(A)\geq(2\pi n)^{\frac{m}{2}}\left(\frac{1-\Delta}{e}\right)^{mn}\geq\frac{(mn)!}{(mn)^{mn}}\cdot\frac{(2\pi n)^{\frac{m}{2}}}{\sqrt{2\pi mn}}(1-\Delta)^{mn}\exp\left(-\frac{1}{12mn}\right).$
Now we choose $n=\lceil\frac{1}{2\log\left(\frac{1}{1-\Delta}\right)}\rceil\geq 2$, so that
$\mathrm{Perm}(A)\geq\frac{(mn)!}{(mn)^{mn}}\cdot(2\pi n)^{\frac{m-1}{2}}\frac{(1-\Delta)^{m}e^{-\frac{m}{2}}}{e^{1/12}\sqrt{m}}\overset{\rm(a)}{\geq}\frac{(mn)!}{(mn)^{mn}}\cdot\left(\frac{3}{\Delta}\right)^{\frac{m-1}{2}}\frac{(1-\Delta)^{m}e^{-\frac{m}{2}}}{e^{1/12}\sqrt{m}},$
where (a) uses that $n\sim\frac{1}{2\Delta}$ as $\Delta\to 0$, and thus $2\pi n\geq\frac{3}{\Delta}$ as long as $\Delta\leq\Delta_{0}$ is small enough. Moreover, since the leading exponential term of
$3^{\frac{m-1}{2}}\frac{(1-\Delta)^{m}e^{-\frac{m}{2}}}{e^{1/12}\sqrt{m}}$
is $\left(\frac{3}{e}(1-\Delta)^{2}\right)^{m/2}$, the above quantity is no smaller than $1$ for ${\mathsf{C}}_{0}$ large enough and $\Delta_{0}$ small enough. Consequently, under the above choice of $({\mathsf{C}}_{0},\Delta_{0})$, it holds that
$\mathrm{Perm}(A)\geq\frac{(mn)!}{(mn)^{mn}}\left(\frac{1}{\Delta}\right)^{\frac{{\mathsf{C}}}{3}},$
i.e. the first statement of the lemma holds with $r=\frac{1}{3}$.
For the second statement, let ${\mathcal{P}}=\{P_{1},\dots,P_{m}\}$ with $m=\lceil{\mathsf{C}}\rceil$, and
$P_{i}=\sqrt{\Delta}\delta_{0}+(1-\sqrt{\Delta})\delta_{i}$
Here $\delta_{x}$ is the Dirac delta measure on the singleton $\{x\}$. It is clear that ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq|{\mathcal{P}}|-1<{\mathsf{C}}$ (cf. Lemma C.1), and $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})=\frac{1}{\Delta}$. Now consider $\chi^{2}(\mathbb{P}_{mn}\|\mathbb{Q}_{mn})$, with each $P_{i}$ appearing $n$ times. By Lemma 4.1,
$\chi^{2}(\mathbb{P}_{mn}\|\mathbb{Q}_{mn})=\frac{(mn)^{mn}}{(mn)!}\mathrm{Perm}(A)-1,$
where the matrix $A\in\mathbb{R}^{mn\times mn}$ is given by
$A=\left(\frac{\sqrt{\Delta}}{m}+(1-\sqrt{\Delta})I_{m}\right)\otimes\frac{J_{n}}{n}.$
By the proof of the first part, as long as ${\mathsf{C}}\geq{\mathsf{C}}_{0}$ is large enough and $\sqrt{\Delta}\leq\sqrt{\Delta_{0}}$ is small enough, there is a suitable choice of $n\in\mathbb{N}$ such that
$\chi^{2}(\mathbb{P}_{mn}\|\mathbb{Q}_{mn})\geq\left(\frac{1}{\sqrt{\Delta}}\right)^{\frac{{\mathsf{C}}}{3}}-1=\left(\frac{1}{\Delta}\right)^{\frac{{\mathsf{C}}}{6}}-1,$
establishing the second part with $r^{\prime}=\frac{1}{6}$.
C.12 Proof of Lemma 5.3
Let $\delta=\delta(\varepsilon)>0$ be any fixed constant such that
$(1+\delta)(1-\varepsilon)<1-\frac{\varepsilon}{2}.$ (46)
Our proof is based on the following I-MMSE formula [GSV05]: for an SNR parameter $\lambda\geq 0$ and $X^{\lambda}:=\sqrt{\lambda}\theta+N$ with $N\sim{\mathcal{N}}(0,I_{n})$ independent of $\theta$, for any prior distribution of $\theta$ it holds that
$I(\theta;X^{\lambda})=\frac{1}{2}\int_{0}^{\lambda}\mathbb{E}\left[\|\theta-\mathbb{E}\left[\theta|X^{t}\right]\!\|^{2}\right]\mathrm{d}t=:\frac{1}{2}\int_{0}^{\lambda}\mathrm{mmse}(t)\mathrm{d}t.$
We denote by $I_{\mathbb{P}}(\theta;X^{\lambda})$ and $I_{\mathbb{Q}}(\theta;X^{\lambda})$ the mutual information under the permutation prior and the i.i.d. prior, respectively. Similarly we use the notations $(\mathrm{mmse}_{\mathbb{P}}(t),\mathrm{mmse}_{\mathbb{Q}}(t))$ for the minimum mean squared errors (MMSEs) under these priors. By the I-MMSE formula,
$\frac{1}{2}\int_{1}^{1+\delta}\left(\mathrm{mmse}_{\mathbb{P}}(t)-\mathrm{mmse}_{\mathbb{Q}}(t)\right)\mathrm{d}t=I_{\mathbb{P}}(\theta;X^{1+\delta})-I_{\mathbb{Q}}(\theta;X^{1+\delta})-(I_{\mathbb{P}}(\theta;X^{1})-I_{\mathbb{Q}}(\theta;X^{1})).$
For the difference in the mutual information, the inequality (1) in Corollary 1.4 states that
$0\leq I_{\mathbb{Q}}(\theta;X^{\lambda})-I_{\mathbb{P}}(\theta;X^{\lambda})\leq 10\sum_{\ell=2}^{n}I_{\chi^{2}}(\theta_{1};X_{1}^{\lambda})^{\ell}.$
Consequently, using the decreasing property of $t\mapsto\mathrm{mmse}(t)$ under both priors, we have
$\frac{\delta}{2}\left(\mathrm{mmse}_{\mathbb{P}}(1)-\mathrm{mmse}_{\mathbb{Q}}(1+\delta)\right)\geq\frac{1}{2}\int_{1}^{1+\delta}\left(\mathrm{mmse}_{\mathbb{P}}(t)-\mathrm{mmse}_{\mathbb{Q}}(t)\right)\mathrm{d}t\geq-10\sum_{\ell=2}^{n}I_{\chi^{2}}(\theta_{1};X_{1}^{1+\delta})^{\ell},$
which rearranges to
$\mathrm{mmse}_{\mathbb{P}}(1)\geq\mathrm{mmse}_{\mathbb{Q}}(1+\delta)-\frac{20}{\delta}\sum_{\ell=2}^{n}I_{\chi^{2}}(\theta_{1};X_{1}^{1+\delta})^{\ell}.$ (47)
In the above inequality, note that $\mathrm{mmse}_{\mathbb{P}}(1)$ is precisely the target Bayes risk under the permutation prior. As for $\mathrm{mmse}_{\mathbb{Q}}(1+\delta)$, this quantity corresponds to the Bayes risk under a rescaled i.i.d. prior $\left((1-\alpha)\delta_{0}+\alpha\delta_{\mu_{0}}\right)^{\otimes n}$ for the Gaussian mean vector, with
$\mu_{0}:=\sqrt{(1+\delta)(1-\varepsilon)}\mu=\sqrt{2(1+\delta)(1-\varepsilon)\log(1/R^{p})},\qquad\alpha:=\frac{\lfloor s\rfloor}{n}.$
As $\mu_{0}<\mu$ by (46) and $s=\omega(1)$, this Bayes risk has been characterized in [DJ94, Section 6] that $\mathrm{mmse}_{\mathbb{Q}}(1+\delta)=(1-o(1))s\mu_{0}^{2}>(1-o(1))(1-\varepsilon)s\mu^{2}$. Finally, since $\delta=\delta(\varepsilon)>0$ is a fixed constant and $s\mu^{2}=\Omega(1)$, the inequality (47) proves the corollary as long as we show that
$I_{\chi^{2}}(\theta_{1};X_{1}^{1+\delta})=o_{\varepsilon}(1).$ (48)
To prove (48), note that as $R=o(1)$, for $n$ large enough we have $\mu\geq 1$, so that
[…]
Consequently,
$I_{\chi^{2}}(\theta_{1};X_{1}^{1+\delta})$ =$\overset{\rm(a)}{=}\alpha(1-\alpha)\int\frac{(\mathrm{d}{\mathcal{N}}(\mu_{0},1)-\mathrm{d}{\mathcal{N}}(0,1))^{2}}{\mathrm{d}[\alpha{\mathcal{N}}(\mu_{0},1)+(1-\alpha){\mathcal{N}}(0,1)]}$
=$\overset{\rm(b)}{=}\alpha(1-\alpha)\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\frac{\left(\exp(\mu_{0}Z-\mu_{0}^{2}/2)-1\right)^{2}}{\alpha\exp(\mu_{0}Z-\mu_{0}^{2}/2)+1-\alpha}\right],$
where (a) uses the definition of the $\chi^{2}$ mutual information, and (b) follows from a simple change of measure. To proceed, note that
$\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\frac{\left(\exp(\mu_{0}Z-\mu_{0}^{2}/2)-1\right)^{2}}{\alpha\exp(\mu_{0}Z-\mu_{0}^{2}/2)+1-\alpha}{\mathbf{1}_{\left\{{Z\leq(1+\frac{\varepsilon}{4})\mu_{0}}\right\}}}\right]$
$\leq\frac{1}{1-\alpha}\left(1+\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\exp(2\mu_{0}Z-\mu_{0}^{2}){\mathbf{1}_{\left\{{Z\leq(1+\frac{\varepsilon}{4})\mu_{0}}\right\}}}\right]\right)$
=$\overset{\rm(c)}{=}\frac{1}{1-\alpha}\left(1+e^{\mu_{0}^{2}}\cdot\mathbb{P}\left({\mathcal{N}}(2\mu_{0},1)\leq\left(1+\frac{\varepsilon}{4}\right)\mu_{0}\right)\right)$
$\overset{\rm(d)}{\leq}\frac{1}{1-\alpha}\left(1+\exp\left(\mu_{0}^{2}-\frac{1}{2}\left(1-\frac{\varepsilon}{4}\right)^{2}\mu_{0}^{2}\right)\right)=\frac{1}{1-\alpha}\left(1+\exp\left(\left(\frac{1}{2}+\frac{\varepsilon}{4}-\frac{\varepsilon^{2}}{32}\right)\mu_{0}^{2}\right)\right),$
where (c) is due to a change of measure, and (d) uses $\mathbb{P}({\mathcal{N}}(0,1)\geq t)\leq\exp(-t^{2}/2)$ for every $t\geq 0$. For the remaining quantity, we similarly have
$\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\frac{\left(\exp(\mu_{0}Z-\mu_{0}^{2}/2)-1\right)^{2}}{\alpha\exp(\mu_{0}Z-\mu_{0}^{2}/2)+1-\alpha}{\mathbf{1}_{\left\{{Z>(1+\frac{\varepsilon}{4})\mu_{0}}\right\}}}\right]$
$\leq\frac{1}{\alpha}\mathbb{E}_{Z\sim{\mathcal{N}}(0,1)}\left[\exp(\mu_{0}Z-\mu_{0}^{2}/2){\mathbf{1}_{\left\{{Z>(1+\frac{\varepsilon}{4})\mu_{0}}\right\}}}\right]$
$=\frac{1}{\alpha}\mathbb{P}\left({\mathcal{N}}(\mu_{0},1)>\left(1+\frac{\varepsilon}{4}\right)\mu_{0}\right)\leq\frac{1}{\alpha}\exp\left(-\frac{\varepsilon^{2}\mu_{0}^{2}}{32}\right).$
Since $\alpha\leq\exp\left(-\frac{(1+\varepsilon/2)\mu_{0}^{2}}{2}\right)$, we have
$I_{\chi^{2}}(\theta_{1};X_{1}^{1+\delta})\leq\alpha+2\exp\left(-\frac{\varepsilon^{2}\mu_{0}^{2}}{32}\right)=o_{\varepsilon}(1),$
as claimed. The proof of (48) is then complete.
C.13 Proof of Lemma 5.4
Without loss of generality assume that $i=0$. The proof is via a careful coupling in [GR09]: for $\pi\in S_{n}$ and $j\in[n]$, define joint distributions of $(X_{1},\dots,X_{n})$ as
$P_{\pi}$ $:=P_{\pi(1)}\otimes\cdots\otimes P_{\pi(n)}$
$P_{\pi,j}$ $:=P_{\pi(1)}\otimes\cdots\otimes P_{\pi(\pi^{-1}(j)-1)}\otimes P_{0}\otimes P_{\pi(\pi^{-1}(j)+1)}\otimes\cdots\otimes P_{\pi(n)}.$
For $j=0$ we also define $P_{\pi,0}=P_{\pi}$. Since $\mathbb{P}_{-j}=\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}[P_{\pi,j}]$ for all $j\in\{0,\dots,n\}$, the joint convexity of the $\chi^{2}$ divergence yields
$\chi^{2}(\overline{\mathbb{P}}\|\mathbb{P}_{-0})$ $=\chi^{2}\left(\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\frac{1}{n+1}\sum_{j=0}^{n}P_{\pi,j}\right]\bigg{\|}\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[P_{\pi}\right]\right)$
$\leq\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\chi^{2}\left(\frac{1}{n+1}\sum_{j=0}^{n}P_{\pi,j}\bigg{\|}P_{\pi}\right)\right]$
$\overset{\rm(a)}{\leq}\frac{n}{n+1}\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\chi^{2}\left(\frac{1}{n}\sum_{j=1}^{n}P_{\pi,j}\bigg{\|}P_{\pi}\right)\right].$
Here (a) uses $\frac{1}{n+1}\sum_{j=0}^{n}P_{\pi,j}=\frac{P_{\pi}}{n+1}+\frac{n}{n+1}\cdot\frac{1}{n}\sum_{j=1}^{n}P_{\pi,j}$ and the convexity again. To proceed we perform the second moment computation. By simple algebra, the likelihood ratio between $P_{\pi,j}$ and $P_{\pi}$ is
$\frac{\mathrm{d}P_{\pi,j}}{\mathrm{d}P_{\pi}}(X_{1},\dots,X_{n})=\frac{\mathrm{d}P_{0}}{\mathrm{d}P_{j}}(X_{\pi^{-1}(j)}).$
Consequently, for $j\neq j^{\prime}$,
$\int\frac{\mathrm{d}P_{\pi,j}\mathrm{d}P_{\pi,j^{\prime}}}{\mathrm{d}P_{\pi}}$ $=\mathbb{E}_{(X_{1},\dots,X_{n})\sim P_{\pi}}\left[\frac{\mathrm{d}P_{0}}{\mathrm{d}P_{j}}(X_{\pi^{-1}(j)})\frac{\mathrm{d}P_{0}}{\mathrm{d}P_{j^{\prime}}}(X_{\pi^{-1}(j^{\prime})})\right]$
$=\mathbb{E}_{X_{\pi^{-1}(j)}\sim P_{j}}\left[\frac{\mathrm{d}P_{0}}{\mathrm{d}P_{j}}(X_{\pi^{-1}(j)})\right]\mathbb{E}_{X_{\pi^{-1}(j^{\prime})}\sim P_{j^{\prime}}}\left[\frac{\mathrm{d}P_{0}}{\mathrm{d}P_{j^{\prime}}}(X_{\pi^{-1}(j^{\prime})})\right]=1,$
and
$\int\frac{(\mathrm{d}P_{\pi,j})^{2}}{\mathrm{d}P_{\pi}}=\mathbb{E}_{(X_{1},\dots,X_{n})\sim P_{\pi}}\left[\left(\frac{\mathrm{d}P_{0}}{\mathrm{d}P_{j}}(X_{\pi^{-1}(j)})\right)^{2}\right]=\chi^{2}(P_{0}\|P_{j})+1\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1.$
Therefore, the second moment method gives
$\chi^{2}\left(\frac{1}{n}\sum_{j=1}^{n}P_{\pi,j}\bigg{\|}P_{\pi}\right)=\frac{1}{n^{2}}\sum_{j,j^{\prime}=1}^{n}\int\frac{\mathrm{d}P_{\pi,j}\mathrm{d}P_{\pi,j^{\prime}}}{\mathrm{d}P_{\pi}}-1\leq\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n},$
and we obtain the first claim. The second claim follows from
$H^{2}(\mathbb{P}_{-i},\mathbb{P}_{-j})\overset{\rm(b)}{\leq}2H^{2}(\overline{\mathbb{P}},\mathbb{P}_{-i})+2H^{2}(\overline{\mathbb{P}},\mathbb{P}_{-j})\overset{\rm(c)}{\leq}2\chi^{2}(\overline{\mathbb{P}}\|\mathbb{P}_{-i})+2\chi^{2}(\overline{\mathbb{P}}\|\mathbb{P}_{-j})\leq\frac{4{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n+1},$
where (b) is the triangle inequality of the Hellinger distance, and (c) uses $H^{2}(P,Q)\leq\chi^{2}(P\|Q)$.
C.14 Proof of Lemma 5.5
Since
$\mathbb{E}\left[\vartheta_{1}|X_{1}\right]-\mathbb{E}\left[\vartheta_{1}|X\right]=\sum_{i=1}^{n}\theta_{i}\left(P_{\vartheta_{1}=\theta_{i}|X_{1}}-P_{\vartheta_{1}=\theta_{i}|X}\right)=\sum_{i=1}^{n}\theta_{i}P_{\vartheta_{1}=\theta_{i}|X_{1}}\left(1-\frac{P_{X_{2}^{n}|\vartheta_{1}=\theta_{i}}}{P_{X_{2}^{n}|X_{1}}}\right)$
according to the Bayes rule, we have
$\mathbb{E}\left[\left(\mathbb{E}\left[\vartheta_{1}|X_{1}\right]-\mathbb{E}\left[\vartheta_{1}|X\right]\right)^{2}\right]$
$=\mathbb{E}_{X_{1}}\sum_{i=1}^{n}\sum_{j=1}^{n}\theta_{i}\theta_{j}P_{\vartheta_{1}=\theta_{i}|X_{1}}P_{\vartheta_{1}=\theta_{j}|X_{1}}\mathbb{E}_{X_{2}^{n}|X_{1}}\left[\left(1-\frac{P_{X_{2}^{n}|\vartheta_{1}=\theta_{i}}}{P_{X_{2}^{n}|X_{1}}}\right)\left(1-\frac{P_{X_{2}^{n}|\vartheta_{1}=\theta_{j}}}{P_{X_{2}^{n}|X_{1}}}\right)\right].$
Next we fix any realization of $X_{1}$. Let $w_{i}:=P_{\vartheta_{1}=\theta_{i}|X_{1}}$ and $P_{i}:=P_{X_{2}^{n}|\vartheta_{1}=\theta_{i}}$, then $\sum_{i=1}^{n}w_{i}=1$, and
$\overline{P}:=P_{X_{2}^{n}|X_{1}}=\sum_{i=1}^{n}w_{i}P_{i}.$
Therefore, the entire term inside $\mathbb{E}_{X_{1}}$ can be written as $\theta^{\top}M\theta$, where
$M_{ij}=w_{i}w_{j}\mathbb{E}_{\overline{P}}\left[\left(1-\frac{\mathrm{d}P_{i}}{\mathrm{d}\overline{P}}\right)\left(1-\frac{\mathrm{d}P_{j}}{\mathrm{d}\overline{P}}\right)\right]=w_{i}w_{j}\left(\int\frac{\mathrm{d}P_{i}\mathrm{d}P_{j}}{\mathrm{d}\overline{P}}-1\right).$
Comparing with the matrix $A$ defined in (26) and the centered version $\overline{A}$ before (33), it is clear that
$M=n\mathrm{diag}(w_{1},\dots,w_{n})\overline{A}\mathrm{diag}(w_{1},\dots,w_{n}),$
and therefore
$\theta^{\top}M\theta\leq n\lambda_{\max}(\overline{A})\sum_{i=1}^{n}w_{i}^{2}\theta_{i}^{2}\leq nM^{2}\lambda_{\max}(\overline{A})\|w\|_{2}^{2}.$
To upper bound the largest eigenvalue of $\overline{A}$, note that Lemma 5.4 implies that $H^{2}(P_{i},P_{j})\leq\frac{4{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}$ for all $i,j\in[n]$. Consequently, following the spectral gap argument in Lemma 4.2, it holds that
$\lambda_{\max}(\overline{A})=\lambda_{2}(A)\leq 1-\left(1-\max_{i,j}\frac{H^{2}(P_{i},P_{j})}{2}\right)^{2}\leq\max_{i,j}{H^{2}(P_{i},P_{j})}\leq\frac{4{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}{n}.$
In addition,
$\mathbb{E}_{X_{1}}\left[\|w\|_{2}^{2}\right]$ $=\mathbb{E}_{X_{1}}\left[\sum_{i=1}^{n}P_{\vartheta_{1}=\theta_{i}|X_{1}}^{2}\right]=\sum_{i=1}^{n}\mathbb{E}_{X_{1}}\left[\left(\frac{\mathrm{d}P_{\theta_{i}}}{\sum_{j=1}^{n}\mathrm{d}P_{\theta_{j}}}\right)^{2}\right]$
$=\frac{1}{n^{2}}\sum_{i=1}^{n}\left(1+\chi^{2}\bigg{(}P_{\theta_{i}}\bigg{\|}\frac{1}{n}\sum_{j=1}^{n}P_{\theta_{j}}\bigg{)}\right)\leq\frac{1+{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}{n}.$
A combination of the above inequalities shows that
$\mathbb{E}\left[\left(\mathbb{E}\left[\vartheta_{1}|X_{1}\right]-\mathbb{E}\left[\vartheta_{1}|X\right]\right)^{2}\right]=\mathbb{E}_{X_{1}}[\theta^{\top}M\theta]$ $\leq 4M^{2}{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})\cdot\mathbb{E}_{X_{1}}\left[\|w\|_{2}^{2}\right]$
$\leq\frac{4M^{2}{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(1+{\mathsf{C}}_{\chi^{2}}({\mathcal{P}}))}{n},$
as claimed.